### v2

# MLC at Feigenbaum points

Abstract.
We prove a priori bounds for Feigenbaum quadratic polynomials, i.e., infinitely renormalizable polynomials $f_{c}:z\mapsto z^{2}+c$ of bounded type. It implies local connectivity of the corresponding Julia sets $J(f_{c})$ and MLC (local connectivity of the Mandelbrot set ${\mathcal{M}}$) at the corresponding parameters $c$. It also yields the scaling Universality, dynamical and parameter, for the corresponding combinatorics. The MLC Conjecture was open for the most classical period-doubling Feigenbaum parameter as well as for the complex tripling renormalizations. Universality for the latter was conjectured by Goldberg-Khanin-Sinai in the early 1980s.
1. Introduction
1.1. Brief history
The MLC Conjecture on the local connectivity of the Mandelbrot set ${\mathcal{M}}$, put forward by Douady and Hubbard in the mid 1980s [DH1], is the central problem of contemporary Holomorphic Dynamics. It would imply a precise topological model for ${\mathcal{M}}$ and the Fatou Conjecture on the density of hyperbolic maps in ${\mathcal{M}}$, and it is intimately related to the Mostow-Thurston Rigidity phenomenon in 3D Hyperbolic Geometry.
Around 1990 Yoccoz proved MLC at any parameter $c\in{\mathcal{M}}$ that is not infinitely renormalizable (in the quadratic-like sense) thus linking the problem tightly to the Quadratic-like Renormalization Theory (see [H, M1]). First advances in this direction appeared in [L1], where MLC was established for infinitely renormalizable parameters of high type and the general problem was reduced, under some circumstances, to the problem of a priori bounds.
Quadratic-like renormalization appears in two flavors, primitive and satellite. The above results were concerned with the primitive case. Next breakthrough in this direction appeared in the work of Jeremy Kahn [K] who established a priori bounds, and hence MLC, for all infinitely renormalizable parameters of bounded primitive type (using the Covering Lemma [KL1]). It followed up with the work [KL2, KL3] handling the definitely primitive case.
It took more than 10 years to prove MLC at some infinitely renormalizable parameters of bounded satellite type [DL1] (based upon the Pacmen Renormalization Theory developed in [DLS]). However, the most prominent parameter, the period-doubling Feigenbaum point corresponding to the cascade of doubling renormalizations (see Figure 1) was not covered by this result. In this paper we are filling in this gap by proving MLC at the Feigenbaum point, and in fact, in all infinitely renormalizable maps of bounded type (for which we will still preserve the same name).
Remark 1.1.
The universality of period-doubling cascade was discovered in the mid 70s by Feigenbaum [F1, F2] (the parameter part, see Figure 1) and independently by Coullet and Tresser [TC, CT] (the dynamical part). It is intimately related to the renormalization phenomenon in the Quantum Field Theory and Statistical Mechanics; the discovery opened up a new universality paradigm in Dynamical Systems.
Refer to captionRefer to captionRefer to caption
Figure 1. Self-similarity of the Mandelbrot set at the period-doubling Feigenbaum parameter: ${\mathcal{M}}$ scales almost linearly under $w\mapsto 4.6692\dots w+o(|w|^{1+\varepsilon})$.
1.2. Statement of the main result and its consequences
A Feigenbaum map is an infinitely renormalizable quadratic-like map $f:U\rightarrow V$ with bounded combinatorics, i.e., all renormalization periods of $f$ are bounded by some $\bar{p}$ (see §2.2.5 for precise definitions). One says that such an $f$ has a priori bounds if its quadratic-like renormalizations ${\mathcal{R}}^{n}f:U^{n}\rightarrow V^{n}$ can be selected so that
$\operatorname{mod}(V^{n}\setminus U^{n})\geq\mu>0.$
The bounds are called beau if $\mu$ depends only on the combinatorial bound $\bar{p}$ as long as $n$ is big enough (depending on $\operatorname{mod}(V\setminus U)$).
Refer to captionRefer to captionRefer to caption
Figure 2. Illustration to the Goldberg-Khanin-Sinai Conjecture: the Mandelbrot set scales almost linearly at the parameter of the period-tripling bifurcation. (The first zoom magnifies a deep renormalization level.)
Theorem A.
Any Feigenbaum quadratic-like map has a priori beau bounds.
Together with the Rigidity Theorem of [L1], it implies MLC at the corresponding parameters:
Theorem B.
The Mandelbrot set is locally connected at any Feigenbaum parameter.
The Renormalization Conjecture asserts that for any combinatorial bound $\bar{p}$ the renormalization transformation ${\mathcal{R}}$ has a renormalization horseshoe ${\mathcal{A}}_{\bar{p}}$ of Feigenbaum maps of type $\bar{p}$ on which it acts hyperbolically with one-dimensional unstable foliation. Together with [L2] (see also [S2, McM2, AL]) we obtain:
Theorem C.
For any combinatorial bound $\bar{p}$, the Renormalization Conjecture is valid in the space of quadratic-like maps.
Let us consider the set ${\mathcal{I}}_{\bar{p}}$ of Feigenbaum parameters $c\in{\mathcal{M}}$ of type $\bar{p}$. Theorem A implie that each $c\in{\mathcal{I}}_{\bar{p}}$ is the intersection of the nest $M^{n}(c)$ of little $M$-copies corresponding to the $n$-fold renormalizations of $f_{c}$. Let us say that Mandelbrot set is self-similar over ${\mathcal{I}}_{\bar{p}}$ if the straightening map $\chi:{\mathcal{M}}^{n+1}(c)\rightarrow{\mathcal{M}}^{n}(\chi(c))$ (see §2.2.3) is $C^{1+{\delta}}$-conformal at any $c\in{\mathcal{I}}_{\bar{p}}$ (with uniform constants).
The Renormalization Conjecture (Theorem C) implies:
Theorem D.
For any $\bar{p}$, the Mandelbrot set is self-similar over ${\mathcal{I}}_{\bar{p}}$.
For the satellite triplings, the last two theorems confirm the Goldberg-Sinai-Khanin Conjecture [GKS] from the 1980s, see Figure 2.
1.3. Rough outline of the proof
The historical developments of ideas behind the proof are summarized in [L4]. For a ql map $f\colon U\to V$, its width is
${\mathcal{W}}(f)={\mathcal{W}}(V\setminus{\mathfrak{K}}_{f})=\frac{1}{\operatorname{mod}(V\setminus{\mathfrak{K}}_{f})},$
where ${\mathfrak{K}}_{f}$ is the filled Julia set of $f$. Our goal is to bound ${\mathcal{W}}(f)$ from above. We will first review the main ideas for the bounded primitive case, then we will discuss how to complete the argument in the bounded satellite case.
${\mathcal{G}}_{1}$${\mathcal{G}}_{2}$${\mathcal{G}}^{\prime}_{1}$${\mathcal{G}}^{\prime\prime}_{1}$${\mathcal{G}}^{\prime}_{2}$
Figure 3. Primitive pull-off in the airplane combinatorics. If ${\mathcal{W}}_{\mathrm{loc}}^{\mathrm{ver}}({\mathfrak{K}}_{j})\ll{\mathcal{W}}_{\mathrm{loc}}({\mathfrak{K}}_{j})$, then most curves of the horizontal rectangles ${\mathcal{G}}_{1},{\mathcal{G}}_{2}$ overflow their preimages ${\mathcal{G}}^{\prime}_{1},{\mathcal{G}}^{\prime}_{2},{\mathcal{G}}^{\prime\prime}_{1}$ as illustrated. Such a configuration is impossible by the Grötzsch inequality; i.e, (1.4) holds. Peripheral preimages are omitted.
1.3.1. Pull-off in the primitive case [K, KL1]
Consider the $n$th renormalization $f_{n}={\mathcal{R}}^{n}f$ of $f=f_{0}$, and let ${\mathfrak{K}}_{0},{\mathfrak{K}}_{1},\dots,{\mathfrak{K}}_{p-1}$ be the periodic cycle of little filled Julia sets of $f_{n}$ in the dynamical plane of $f_{0}\colon U\to V$. Assuming that the renormalization is primitive, all ${\mathfrak{K}}_{i}$ are pairwise disjoint.
Consider the thin-thick decomposition of $V^{\prime}\coloneqq V\setminus\bigcup_{i}{\mathfrak{K}}_{i}\,$: there are finitely many pairwise disjoint rectangles ${\mathcal{G}}_{i}$ between components of $\partial V^{\prime}$ such that, up to $O_{p}(1)$ of width, the ${\mathcal{G}}_{i}$ contain all non-peripheral curves in $V^{\prime}$ with endpoints in $\partial V^{\prime}$, see Figure 9 for an illustration of a thin-thick decomposition. A rectangle ${\mathcal{G}}_{i}$ is called
  • vertical if it connects the outer boundary $\partial^{\mathrm{out}}V^{\prime}\coloneqq\partial V$ of $V^{\prime}$ to one of its inner boundary components;
  • horizontal if it connects two inner (i.e., non-outer) boundary components of $V^{\prime}$.
Below we assume that ${\mathcal{W}}(f_{n})\gg_{p}1$. We have:
(1.1) ${\mathcal{W}}(f)\geq\sum_{\text{ vertical }{\mathcal{G}}_{i}}{\mathcal{W}}({\mathcal{G}}_{i})$
and, for every ${\mathfrak{K}}_{j}$:
(1.2) ${\mathcal{W}}(f_{n})={\mathcal{W}}_{\mathrm{loc}}({\mathfrak{K}}_{j})+O_{p}(1)\coloneqq\sum_{\begin{subarray}{c}{\mathcal{G}}_{i}\text{ adjacent}\\ \text{to }{\mathfrak{K}}_{j}\end{subarray}}{\mathcal{W}}({\mathcal{G}}_{i})+O_{p}(1),$
i.e. the sum in (1.2) is taken over all rectangles ${\mathcal{G}}_{i}$ adjacent to ${\mathfrak{K}}_{j}$. (A “$\psi$-ql renormalization” is used to achieve (1.2), see §3.43.5.2, and (3.23).) Set also
(1.3) ${\mathcal{W}}_{\mathrm{loc}}^{\mathrm{ver}}({\mathfrak{K}}_{j})\coloneqq\sum_{\begin{subarray}{c}\text{vertical }{\mathcal{G}}_{i}\\ \text{adjacent to }{\mathfrak{K}}_{j}\end{subarray}}{\mathcal{W}}({\mathcal{G}}_{i}).$
A fundamental fact is that the horizontal ${\mathcal{G}}_{i}$ are eventually (after several restrictions of the domain) aligned with the Hubbard tree $T_{f}$ of $f$ as it shown on Figure 3. Applying the Grötzsch inequality to the horizontal rectangles and their pullbacks (see the caption of Figure 3), we experience a definite loss of the horizontal weight in favor of the vertical one. The created definite vertical weight can then be pushed forward by means of the Covering Lemma [KL1]. Since all the local weights are comparable,  [^1] This observation goes back to McMullen [McM1, Theorem 9.3]. we obtain:
(1.4) ${\mathcal{W}}_{\mathrm{loc}}^{\mathrm{ver}}({\mathfrak{K}}_{j})\asymp{\mathcal{W}}_{\mathrm{loc}}({\mathfrak{K}}_{j}){\ \ }{\ \ }{\ \ }\text{ for every }{\mathfrak{K}}_{j},$
where $\asymp$” is independent of $p$.
We stress that a key combinatorial ingredient used for (1.4) is the absence of periodic horizontal rectangles; i.e. a rectangle that has an iterated lift homotopic to itself rel small Julia sets, see §2.4, §3.5.8.
Combining  (1.1), (1.4), (1.2), and assuming $p\gg 1$, we obtain:
(1.5) ${\begin{array}[]{c}{\mathcal{W}}(f)\geq\sum_{\text{ vertical }{\mathcal{G}}_{i}}{\mathcal{W}}({\mathcal{G}}_{i})\ =\ \sum_{j=1}^{p}{\mathcal{W}}_{\mathrm{loc}}^{\mathrm{ver}}({\mathfrak{K}}_{j})\ \asymp\\ \\ \sum_{j=1}^{p}{\mathcal{W}}_{\mathrm{loc}}({\mathfrak{K}}_{j})\ \asymp p\ {\mathcal{W}}(f_{n})\ \gg\ 2{\mathcal{W}}(f_{n});\end{array}}$
i.e:
(1.6) ${\mathcal{W}}(f)\geq 2{\mathcal{W}}(f_{n})$
implying a priori bounds in the primitive case. (For instance, by the Record Argument: by selecting “record levels”, on which the degeneration exceeds all the preceding ones, we immediately arrive at a contradiction.)
${\mathcal{R}}_{0}$${\mathcal{G}}$${\mathcal{R}}_{1}$${\mathfrak{B}}_{0}$${\mathfrak{B}}_{1}$${\mathfrak{B}}_{2}$${\mathfrak{B}}_{3}$${\mathcal{R}}^{\prime}_{0}$$f^{2}$${\mathcal{R}}^{\prime\prime}_{0}$${\mathcal{G}}^{\prime}$$f^{2}$${\mathcal{R}}^{\prime}_{1}$$f^{2}$${\mathcal{R}}^{\prime\prime}_{1}$
Figure 4. Pull-off in the Feigenbaum combinatorics for non-periodic rectangles. Either most of the horizontal degeneration is within periodic rectangles ${\mathcal{R}}_{0},{\mathcal{R}}_{1}$ such that a $(1-\delta)$-portion of each ${\mathcal{R}}_{i}$ overflows its periodic lifts ${\mathcal{R}}^{\prime}_{i}$. Or we have ${\mathcal{W}}^{\mathrm{ver}}_{\mathrm{loc}}({\mathfrak{K}}_{j})\succeq_{\delta}{\mathcal{W}}_{\mathrm{loc}}({\mathfrak{K}}_{j}).$
1.3.2. Satellite case
Let us now assume that $f$ is an infinitely renormalizable map with the periodic-doubling combinatorics; i.e., $f$ is hybrid equivalent to a map in the period-doubling Feigenbaum combinatorial class. The general bounded case is similar.
Let ${\mathfrak{B}}\coloneqq{\mathfrak{K}}^{1}_{0}\cup{\mathfrak{K}}^{1}_{1}$ be the bouquet consisting of two level 11 11 little Julia sets of $f$. Instead of ${\mathcal{W}}(f)$, we will measure
${\mathcal{W}}_{\bullet}(f)\coloneqq{\mathcal{W}}(V\setminus{\mathfrak{B}}).$
In the dynamical plane of $f=f_{0}$, consider the periodic cycle of little bouquets ${\mathfrak{B}}_{0},{\mathfrak{B}}_{1},\dots,{\mathfrak{B}}_{p-1}$ (where $p=2^{n}$) associated with $f_{n}={\mathcal{R}}^{n}f$. We enumerate these bouquets linearly from left-to-right. Consider the thick-thin decomposition of $V^{\prime}\coloneqq V\setminus\bigcup_{i}{\mathfrak{B}}_{i}$. The following properties are established similarly as in the primitive case considered above:
  • Estimates (1.1) and (1.2), where little Julia sets ${\mathfrak{K}}_{i}$ are replaced with bouquets ${\mathfrak{B}}_{i}$;
  • alignment of the horizontal rectangles with the Hubbard tree of $f$ (after several restrictions of the domain).
However, unlike in the primitive case §1.3.1, $f$ may have periodic horizontal rectangles ${\mathcal{R}}_{k}$ aligned with the Hubbard tree between ${\mathfrak{B}}_{2k}$ and ${\mathfrak{B}}_{2k+1}$ as shown on Figure 4. This leads us to the following alternative: either most of the total degeneration is within the ${\mathcal{R}}_{k}$ or a substantial part of the total degeneration is vertical (i.e. (1.4) holds). More precisely, choose a small $\delta>0$. We have:
  1. (I)
    either there is a periodic rectangle ${\mathcal{R}}_{k}$ connecting ${\mathfrak{B}}_{2k},{\mathfrak{B}}_{2k+1}$ such that a $(1-\delta)$ part of ${\mathcal{R}}_{k}$ overflows its iterated lift ${\mathcal{R}}^{\prime}_{k}$ under $f^{2^{n-1}}$;
  2. (II)
    or else we have:
    ${\mathcal{W}}_{\mathrm{loc}}^{\mathrm{ver}}({\mathfrak{K}}_{j})\asymp_{\delta}{\mathcal{W}}_{\mathrm{loc}}({\mathfrak{K}}_{j}){\ \ }{\ \ }{\ \ }\text{ for every }{\mathfrak{K}}_{j}$
If (II) holds, then assuming $p\gg_{\delta}1$ and ${\mathcal{W}}_{\bullet}(f_{n})\gg_{p,\delta}1$ and repeating the argument of (1.5), we obtain
  1. (${}^{\prime}$)
    ${\mathcal{W}}_{\bullet}(f)>2{\mathcal{W}}_{\bullet}(f_{n})$;
the Record Argument leads to a contradiction in case (${}^{\prime}$).
The dichotomy “Case (I) vs Case (${}^{\prime}$)” is stated as Theorem 4.1. This is a refinement of [K].
Consider now Case (I). As Figure 5 illustrates, bouquets ${\mathfrak{B}}_{0},{\mathfrak{B}}_{1}$ grow under lifting thus the “combinatorial distance” between $\widetilde{\mathfrak{B}}_{0},\widetilde{\mathfrak{B}}_{1}$ shrinks. This allows us to detect laminations ${\mathcal{S}}_{i}$ (the “differences” between ${\mathcal{R}}={\mathcal{R}}_{k}$ and its lift ${\mathcal{R}}^{\prime}$) that are much wider than ${\mathcal{W}}({\mathcal{R}})\succeq{\mathcal{W}}_{\bullet}(f_{n})$; i.e. ${\mathcal{W}}({\mathcal{S}}_{i})\geq C_{\delta}{\mathcal{W}}({\mathcal{R}})$ with $C_{\delta}\to\infty$ as $\delta\to 0$. There are two possibilities: either a substantial part of ${\mathcal{S}}_{i}$ hits preperiodic bouquets associated with $f_{n+1}$ or a substantial part of ${\mathcal{S}}_{i}$ creates a “wave” (see §5) above such a preperiodic bouquet. In the latter case, the Wave Lemma implies that ${\mathcal{W}}_{\bullet}(f_{n-1})\gg{\mathcal{W}}_{\bullet}(f_{n})$ and the Record Argument is applicable.
In the former case, we obtain ${\mathcal{W}}_{\bullet}(f_{n+1})\succeq{\mathcal{W}}({\mathcal{S}}_{i})\succeq C_{\delta}{\mathcal{W}}(f_{n})$; i.e.:
${\mathcal{W}}_{\bullet}(f_{n+1})\geq C^{\prime}_{\delta}{\mathcal{W}}(f_{n}){\ \ }{\ \ }\text{ with }{\ \ }C^{\prime}_{\delta}\to\infty\text{{\ \ }}{as}{\ \ }\delta\to 0.$
For a sufficiently small $\delta$, this eventually contradicts the Teichmüller contraction within hyperbolic classes [S1] (stated as Proposition 2.7):
${\mathcal{W}}_{\bullet}(f_{n})=O\left(\Delta^{n}\right){\ \ }{\ \ }{\ \ }\text{ for some }\Delta>1.$
Finally, we will convert in Theorem 7.2 bounds for ${\mathcal{W}}_{\bullet}(f_{m})$ into bounds for ${\mathcal{W}}(f_{m})$.
Remark 1.2.
In the paper, we will be measuring the degeneration around “bushes” §2.3, which are Hubbard continua enhanced with the periodic cycle of level one little Julia sets. An alternative would be to keep the Julia sets for primitive renormalization and use the satellite flowers in the satellite case. We believe that it is also possible to work with Hubbard continua; however, the Wave Lemma §5 and the transition toward classical QL-bounds will be more subtle.
Acknowledgments
The first author was partially supported by the NSF grant DMS 20555322055532 20555322055532 and the ERC grant “HOLOGRAM”. The second author has been partly supported by the NSF, the Hagler and Clay Fellowships, the Institute for Theoretical Studies at ETH (Zurich), and MSRI (Berkeley).
We would like to thank Jeremy Kahn for many stimulating discussions.
Results of this paper were first announced in 2021 at the Quasiworld Seminar and at the CIRM conference “Advancing Bridges in Complex Dynamics”, see the talk by the second author [L4].
A number of our pictures are made with W. Jung’s program Mandel.
${\mathcal{R}}$${\mathfrak{B}}_{0}$${\mathfrak{B}}_{1}$${\mathcal{R}}^{\prime}$${\mathcal{S}}_{1}$${\mathcal{S}}_{2}$$\widetilde{\mathfrak{B}}_{0}$$\widetilde{\mathfrak{B}}_{1}$
Figure 5. Pull-off in the Feigenbaum combinatorics for periodic rectangles. Assume that a $(1-\delta)$ part of a periodic rectangle ${\mathcal{R}}$ overflows its iterated lift ${\mathcal{R}}^{\prime}$. Since ${\mathcal{R}}$ also overflows ${\mathcal{S}}_{1},{\mathcal{S}}_{2}$ (that are disjoint from ${\mathcal{R}}^{\prime}$), we obtain from the Grötzsch inequality that ${\mathcal{W}}({\mathcal{S}}_{i})>C_{\delta}{\mathcal{W}}({\mathcal{R}})$ with $C_{\delta}\to\infty$ as $\delta\to 0$.
2. QL renormalization, bushes, and invariant arc diagrams
The reader can find most of the needed background material in the book [L3]. Let us briefly outline the context of this section with emphasis on the notation.
2.1. Outline
Basics aspects of quadratic-like maps are summarized in §2.2. We will usually denote by
  • $f\colon X\to Y$ a ql map,
  • ${\mathfrak{K}}_{f}$ its (filled) Julia set,
  • ${\mathfrak{K}}^{[n]}_{i}\subset{\mathfrak{K}}_{f}$ its level $n$ little Julia sets,
  • ${\boldsymbol{\mathfrak{K}}}^{[n]}\coloneqq\bigcup_{i}{\mathfrak{K}}^{[n]}_{i}$ the union of periodic level $n$ little Julia sets, see (2.6),
  • ${\mathfrak{T}}_{f}\subset{\mathfrak{K}}$ the Hubbard continuum of $f$, see §2.2.2,
  • ${\mathfrak{B}}_{f}\coloneqq{\mathfrak{T}}_{f}\cup{\boldsymbol{\mathfrak{K}}}^{[1]}$ the bush of $f$, see (2.8),
  • ${\mathfrak{B}}_{f}^{(m)}\coloneqq f^{-m}({\mathfrak{B}}_{f})$ the little bush of height $m$,
  • ${\mathfrak{B}}^{[n]}_{i}\equiv{\mathfrak{B}}^{[n],(0)}\subset{\mathfrak{B}}^{[n],(m)}\subset{\mathfrak{K}}^{[n]}_{i}$ the associated little bushes,
  • ${\boldsymbol{\mathfrak{B}}}^{[n]}\coloneqq\bigcup_{i}{\mathfrak{B}}^{[n]}_{i}\subset{\boldsymbol{\mathfrak{K}}}^{[n]}$ the union of periodic level $n$ little bushes,
  • ${\boldsymbol{\mathfrak{B}}}^{[n],(m)}\coloneqq f^{-1}\left({\boldsymbol{\mathfrak{B}}}^{[n]}\right)$ the union of level $n$ and height $m$ little periodic and preperiodic bushes,
  • ${\boldsymbol{\mathfrak{B}}}^{[n],(m)}_{\operatorname{per}}\coloneqq{\boldsymbol{\mathfrak{B}}}^{[n],(m)}\cap{\boldsymbol{\mathfrak{K}}}^{[n]}$ the union of periodic level $n$ and height $m$ little bushes.
If ${\mathfrak{B}}_{f}={\boldsymbol{\mathfrak{K}}}^{[1]}$, i.e., ${\mathfrak{T}}_{f}\subset{\boldsymbol{\mathfrak{K}}}^{[1]}$, then we will also refer to ${\mathfrak{B}}_{f}$ as the satellite flower of $f$. Preperiodic little Julia sets and bushes are defined accordingly, 2.3.1. We will usually assume that
  1. (O)
    ${\mathfrak{K}}^{[n]}_{0}$ contains the critical point of $f$.
The Teichmüller contraction of ql renormalization will be established in §2.5.
In Lemma 2.2, we will show that invariant horizontal arc diagrams are aligned with the Hubbard dendrite. Figures 7 and 8 demonstrate new subtleties arising in the satellite case. In particular, there are genuine periodic arcs, described in Lemma 2.3; they encode almost periodic rectangles (see §3.5.8 and Figure 4) that will have a separate treatment in Section 6.
2.2. Quadratic-like maps
A quadratic-like map $f:X\rightarrow Y$, which will also be abbreviated as a ql map, is a holomorphic double branched covering between two Jordan disks $X\Subset Y\subset{\mathbb{C}}$. It has a single critical point that we usually put at the origin 00 0. The annulus $A=Y\setminus\overline{X}$ is called the fundamental annulus of $f\colon X\to Y$. We let $\operatorname{mod}f:=\operatorname{mod}A$. The filled Julia set ${\mathfrak{K}}_{f}$ is the set of non-escaping points:
${\mathfrak{K}}\equiv{\mathfrak{K}}_{f}\equiv{\mathfrak{K}}(f)\coloneqq\{z:\,f^{n}z\in X,\ n=0,1,2,\dots\}.$
Its boundary is called the Julia set ${\mathfrak{J}}\equiv{\mathfrak{J}}(f)$. The (filled) Julia set is either connected or Cantor, depending on whether the critical point is non-escaping (i.e., $0\in{\mathfrak{K}}(f)$) or otherwise. In this paper, filled Julia sets will play a bigger role than the Julia sets, and we will often skip the adjective “filled” when it is clear (e.g., from the notation) what we mean.
We say that two quadratic-like maps with connected Julia set represent the same germ if they have the same filled Julia set and coincide in some neighborhood of it. For a ql germ $f$, we define
(2.1) $\operatorname{mod}f\coloneqq\sup\operatorname{mod}(Y^{\prime}\setminus X^{\prime}),$
where the supremum is taken over all $f\colon X^{\prime}\to Y^{\prime}$ representing the ql germ of $f$.
Two quadratic-like maps $f:X\rightarrow Y$ and $\tilde{f}:\tilde{X}\rightarrow\tilde{Y}$ are called hybrid conjugate if they are conjugate by a quasiconformal map $h:(Y,X)\rightarrow(\tilde{Y},\tilde{X})$ such that ¯$\bar{\partial}h=0$ a.e. on ${\mathfrak{K}}$.
A simplest example of a quadratic-like map is provided by a quadratic polynomial $f_{c}:z\mapsto z^{2}+c$ restricted to a disk $Y=\mathbb{D}_{R}$ of sufficiently big radius. The Douady and Hubbard Straightening Theorem asserts that any quadratic-like map $f$ is hybrid conjugate to some restricted quadratic polynomial $f_{c}$. Moreover, if $J$ is connected then the parameter $c\in{\mathcal{M}}$ is unique.
As for quadratic polynomials, two fixed points of a quadratic-like map with connected Julia set have a different dynamical meaning. One of them, called $\beta$, is the landing point of a proper arc $\gamma\subset X\setminus K(f)$ such that $f(\gamma)\supset\gamma$. It is either repelling or parabolic with multiplier one. The other fixed point, called $\alpha$, is either non-repelling or a cut-point of the Julia set.
2.2.1. Combinatorial models
A quadratic polynomial $f$ is called periodically hyperbolic (resp., repelling) if it does not have neutral (resp., non-repelling) cycles. Note that in the periodically repelling case, $\operatorname{int}{\mathfrak{K}}=\emptyset$, so ${\mathfrak{J}}={\mathfrak{K}}$. The filled Julia set of a periodically hyperbolic map admits a locally connected combinatorial model ${\mathfrak{K}}^{\mathrm{com}}={\mathfrak{K}}^{\mathrm{com}}(f)$ obtained by taking the quotient of the unit disk by the associated rational lamination (see [L3, §32.1.3]). The combinatorial model is endowed with the induced dynamics $f_{\mathrm{com}}:{\mathfrak{K}}^{\mathrm{com}}\rightarrow{\mathfrak{K}}^{\mathrm{com}}$ which is semi-conjugate to $f$ by a natural projection $\pi:{\mathfrak{K}}\rightarrow{\mathfrak{K}}^{\mathrm{com}}$ whose fibers are “combinatorial classes” – sets of points inseparable by preperiodic points.
2.2.2. Hubbard continua
In the superattracting case (when the crittical point is periodic) the Hubbard tree ${\mathfrak{T}}_{f}$ is defined as the smallest tree in ${\mathfrak{K}}$ containing the postcritical set such that ${\mathfrak{T}}_{f}$ intersects the components of $\operatorname{int}{\mathfrak{K}}$ along internal rays. It is invariant under $f$ and is marked with the orbit of 00 0. If ${\mathfrak{K}}$ is periodically repelling and locally connected, we can define the Hubbard dendrite ${\mathfrak{T}}_{f}$ as the smallest connected closed forward-invariant subset of ${\mathfrak{K}}$ containing the postcritical set.
For a general periodically repelling map, the (combinatorial) Hubbard dendrite ${\mathfrak{T}}_{f}^{\mathrm{com}}\subset{\mathfrak{K}}^{\mathrm{com}}$ is defined for its combinatorial model $f\colon{\mathfrak{K}}^{\mathrm{com}}\selfmap$. Lifting ${\mathfrak{T}}_{f}^{\mathrm{com}}$ via $\pi\colon{\mathfrak{K}}\to{\mathfrak{K}}^{\mathrm{com}}$, we obtain the Hubbard continuum ${\mathfrak{T}}_{f}$.
For a periodically repelling $f$ and $x,y\in{\mathfrak{J}}_{f}={\mathfrak{K}}_{f}$, we define a geodesic continuum
(2.2) $[x,y]\coloneqq\pi^{-1}[\pi(x),\pi(y)],$
where $[\pi(x),\pi(y)]$ is a unique arc in ${\mathfrak{J}}^{\mathrm{com}}$ connecting $\pi(x),\pi(y)$.
2.2.3. QL-renormalization
A quadratic-like map $f:X\rightarrow Y$ is called DH renormalizable (after Douady and Hubbard) if there is a quadratic-like restriction
(2.3) $f_{1}=f_{1,0}={\mathcal{R}}f=f^{p_{1}}:X_{1,0}\rightarrow Y_{1,0}$
with connected Julia set ${\mathfrak{K}}_{0}^{[1]}$ such that the little Julia sets
(2.4) ${\mathfrak{K}}^{[1]}_{i}\coloneqq f^{i}\left({\mathfrak{K}}_{0}^{[1]}\right),\hskip 28.45274ptk=0,\dots,p_{1}-1,$
are either pairwise disjoint or else touch at their $\beta$-fixed points. In the former case the renormalization is called primitive, while in the latter it is called satellite.
Note that there are many ql maps $f_{1}\colon X_{1,0}^{\prime}\rightarrow Y^{\prime}_{1,0}$ that satisfy the above requirements (if $f$ is renormalizable). However, all of them represent the same ql germ.
The map $f_{1}$ in (2.3) is called a pre-renormalization of $f$. If it is considered up to a linear rescaling, it is called the renormalization of $f$. (In what follows, we will often skip the prefix “-pre” as long as it does not lead to a confusion.)
For every ${\mathfrak{K}}^{[1]}_{i}$, there are $Y_{1,i}\Supset X_{1,i}\Supset{\mathfrak{K}}^{[1]}_{i}$ such that
(2.5) $f_{1}=f_{1,i}={\mathcal{R}}f=f^{p_{1}}:X_{1,i}\rightarrow Y_{1,i}$
is a ql map with non-escaping set ${\mathfrak{K}}^{[1]}_{i}={\mathfrak{K}}_{f_{1,i}}$. All maps (2.5) represent conformally conjugate germs. We assume that ${\mathfrak{K}}^{[1]}_{0}$ contains the critical point of $f$; see (O).
2.2.4. Little copies of ${\mathcal{M}}$
The sets ${\mathfrak{K}}^{[1]}_{i}$ in (2.4) are referred to as the little (filled) Julia sets. Their positions in the big Julia set ${\mathfrak{K}}_{f}$ determine the renormalization combinatorics. By the Douady-Hubbard Straightening Theorem [DH2], the set of parameters $c$ for which the quadratic polynomial $f_{c}\colon z\mapsto z^{2}+c$ is renormalizable with a given combinatorics forms a little Mandelbrot copy ${\mathcal{M}}_{1}\subset{\mathcal{M}}$ (see [L3, Theorem 43.1]). The renormalization combinatorics can be formally encoded by the Hubbard tree ${\mathfrak{T}}_{\circ}$ of the superattracting center $c_{\circ}$ of ${\mathcal{M}}_{1}$.
Each little copy ${\mathcal{M}}_{1}$ can be canonically mapped onto the whole Mandelbrot set ${\mathcal{M}}$ by the straightening homeomorphism $\chi_{1}:{\mathcal{M}}_{1}\rightarrow{\mathcal{M}}$.
A little Mandelbrot copy ${\mathcal{M}}_{1}$ is called primitive or satellite depending on the type of the corresponding renormalization. They can be easily distinguished as any satellite copy is attached to some hyperbolic component of $\operatorname{int}{\mathcal{M}}$ and does not have the cusp at its root point.
2.2.5. Infinitely renormalizable maps
The notions of an infinitely DH renormalizable map $f$ with renormalization periods $p_{n}$, and its renormalizations $f_{n}={\mathcal{R}}^{n}f$, are defined naturally.
By default, we assume that $p_{n}$ is the smallest renormalization period after $p_{n-1}$. We will denote by ${\mathfrak{K}}^{[n]}_{i},i\in\{0,1,\dots,p_{n}-1\}$ the level $n$ little Julia sets of $f$ enumerated dynamically so that ${\mathfrak{K}}^{[n]}_{0}$ contains the critical point of $f$. We will write
(2.6) ${\boldsymbol{\mathfrak{K}}}^{[n]}\coloneqq\bigcup_{i}{\mathfrak{K}}^{[n]}_{i}$
The ratios $q_{n}:=p_{n}/p_{n-1}$ are called relative periods. One says that such a map has bounded combinatorics of type $\bar{p}$ if the relative periods are bounded by $\bar{p}$. In this case, the map $f$ is called Feigenbaum of type $\bar{p}$. We say that a Feigenbaum map is primitive/satellite if all its renormalizations are such. A Feigenabaum map has a priori bounds if
(2.7) $\operatorname{mod}{\mathcal{R}}^{n}f\geq{\epsilon}>0$
We say that the family ${\mathcal{F}}_{\bar{p}}$ of Feigenbaum maps of type $\bar{p}$ have beau bounds if there exists $\mu>0$ depending only on $\bar{p}$ such that for any $\nu>0$ there exists $n_{0}=n_{0}(\bar{p},\nu)$ such that for any $f\in{\mathcal{F}}_{\bar{p}}$ with $\operatorname{mod}f\geq\nu$ we have
$\operatorname{mod}{\mathcal{R}}^{n}f\geq\mu\quad\text{for all }\ n\geq n_{0}.$
2.3. Bushes
Consider a DH renormalizable quadratic-like map $f\colon X\to Y$. The bush of $f$ is
(2.8) ${\mathfrak{B}}_{f}\equiv{\mathfrak{B}}(f)\coloneqq{\mathfrak{T}}_{f}\cup{\boldsymbol{\mathfrak{K}}}^{[1]},$
where ${\mathfrak{T}}_{f}$ is the Hubbard continuum and ${\boldsymbol{\mathfrak{K}}}^{[1]}$ is the periodic cycle of level one little Julia sets (2.6). For $m\geq 0$, we define the bush of height $m$ to be
(2.9) ${\mathfrak{B}}^{(m)}={\mathfrak{B}}^{(m)}_{f}\coloneqq f^{-m}({\mathfrak{B}}_{f}).$
2.3.1. Little bushes
Suppose that $f$ is at least $n+1$ times DH renormalizable and let $f_{n}={\mathcal{R}}^{n}f$ be its $n$th renormalization of $f$. Then $f_{n}$ has a well defined bush ${\mathfrak{B}}(f_{n})$. Consider the periodic cycle of little level $n$ filled Julia sets ${\mathfrak{K}}_{i}^{[n]}$ associated with $f_{n}$ in the dynamical plane of $f$. Let $f_{n,i}$ be the $n$th prerenormalization around ${\mathfrak{K}}_{i}^{[n]}$, compare (2.5). Then ${\mathfrak{K}}_{i}^{[n]}$ contains the little bush ${\mathfrak{B}}_{i}^{[n]}\equiv{\mathfrak{B}}(f_{n,i})\simeq{\mathfrak{B}}(f_{n})$ as well as ${\mathfrak{B}}_{i}^{[n],(m)}\equiv{\mathfrak{B}}^{(m)}(f_{n,i})\simeq{\mathfrak{B}}^{(m)}(f_{n})$. Note that ${\mathfrak{B}}^{[n],(m)}_{i}$ is the unique periodic lift of ${\mathfrak{B}}_{i+m}^{[n]}$ under $f^{m}$.
We write
${\boldsymbol{\mathfrak{B}}}^{[n]}\coloneqq\bigcup_{i}{\mathfrak{B}}_{i}^{[n]},{\ \ }{\boldsymbol{\mathfrak{B}}}^{[n],(m)}\coloneqq f^{-1}\left({\boldsymbol{\mathfrak{B}}}^{[n]}\right),{\ \ }{\boldsymbol{\mathfrak{B}}}^{[n],(m)}_{\operatorname{per}}\coloneqq{\boldsymbol{\mathfrak{B}}}^{[n],(m)}\cap{\boldsymbol{\mathfrak{K}}}^{[n]}.$
Observe that ${\boldsymbol{\mathfrak{B}}}^{[n],(m)}_{\operatorname{per}}$ is the union of periodic little bushes of height $m$. A preperiodic bush ${\mathfrak{B}}^{[n],(m)}_{a}$ is a non-periodic connected component of ${\boldsymbol{\mathfrak{B}}}^{[n],(m)}$. The preperiod of ${\mathfrak{B}}^{[n],(m)}_{a}$ is the smallest $s\leq m$ such that $f^{s}\left({\mathfrak{B}}^{[n],(m)}_{a}\right)$ is periodic.
Every periodic or preperiodic little bush ${\mathfrak{B}}_{a}^{[n],(m)}$ is within a unique little periodic or preperiodic filled Julia set ${\mathfrak{K}}^{[n]}_{a}$ associated with $f_{n}$. The ${\mathfrak{B}}_{t}^{[n],(m)}$ exhaust ${\mathfrak{K}}^{[n]}_{a}$:
${\mathfrak{K}}^{[n]}_{a}=\overline{\bigcup_{m}{\mathfrak{B}}^{[n],(m)}_{a}}\ .$
By construction, the ${\mathfrak{B}}^{[n],(m)}_{a}$ are pairwise disjoint but the ${\mathfrak{K}}^{[n]}_{a}$ may touch each other in the satellite case.
2.3.2. Superattracting model
Consider a ql map $f\colon X\to Y$ and assume it is $n+1$ DH renormalizable. Then $f$ is hybrid equivalent to $z^{2}+c$, where $c$ is in a level $n+1$ little copy ${\mathcal{M}}^{[n+1]}_{i}\subset{\mathcal{M}}$. A superattracting model for $f$ of level $n+1$ is any ql map $f_{\circ}\colon X_{\circ}\to Y_{\circ}$ hybrid equivalent to the center of ${\mathcal{M}}^{[n+1]}_{i}$.
It is well-known (follows, for instance, from the lamination theory) that the Hubbard continua ${\mathfrak{T}}_{f}$ and ${\mathfrak{T}}_{\circ}\equiv{\mathfrak{T}}_{f_{\circ}}$ are combinatorially equivalent up to ${\boldsymbol{\mathfrak{B}}}^{[n]}$: there is a bijection between
  • components of ${\mathfrak{T}}_{f}\setminus f^{-m}\left({\boldsymbol{\mathfrak{B}}}^{[n]}_{f}\right)$ and components of ${\mathfrak{T}}_{\circ}\setminus f^{-m}_{\circ}\left({\boldsymbol{\mathfrak{B}}}^{[n]}_{\circ}\right)$ for all $m\geq 0$; and
  • components of $f^{-m}\left({\boldsymbol{\mathfrak{B}}}^{[n]}_{f}\right)$ and components of $f^{-m}_{\circ}\left({\boldsymbol{\mathfrak{B}}}^{[n]}_{\circ}\right)$ for all $m\geq 0$
that respects the adjacency, natural embedding, and dynamical relations between respective components. In other words, the above components define equivalent Markov partitions for $f,f_{\circ}$.
${\mathfrak{B}}^{[1]}_{a}$${\mathfrak{B}}^{[1],(6p_{2})}_{a}$${\mathfrak{B}}^{\prime\prime}_{1}$$\ell_{1}$${\mathfrak{B}}^{\prime}_{1}$${\mathfrak{B}}_{i}^{[2]}$${\mathfrak{B}}_{i+1}^{[2]}$${\mathfrak{B}}^{\prime\prime}_{2}$${\mathfrak{B}}^{\prime}_{2}$$\ell_{2}$${\mathfrak{B}}^{[1]}_{b}$$\alpha$
Figure 6. Illustration to Lemma 2.1.
2.3.3. Satellite combinatorics
Suppose that $f$ is at least 33 33 times renormalizable and the first renormalization of $f$ is satellite. Then all ${\mathfrak{K}}_{i}^{[1]}$ are organized in the satellite flower around the $\alpha$ fixed point of $f$:
[…]
where ${\mathfrak{T}}_{f}$ is the Hubbard continuum of $f$. For $s\geq 0$ we write ${\mathfrak{T}}^{(s)}_{f}=f^{-s}({\mathfrak{T}}_{f})$. As before, the $p_{k}$ are the periods of level $k$ Julia sets.
Lemma 2.1 (Satellite flower).
For a satellite $f$ as above, let ${\mathfrak{B}}^{[1]}_{a},{\mathfrak{B}}^{[1]}_{b}$ be its two level 11 11 bushes. Then the continuum ${\mathfrak{T}}^{(6p_{2})}_{f}$ contains geodesic continua $\ell_{1}$ and $\ell_{2}$ (as in §2.2.2) connecting preperiodic lifts ${\mathfrak{B}}^{\prime}_{1},{\mathfrak{B}}^{\prime\prime}_{1}$ and ${\mathfrak{B}}^{\prime}_{2},{\mathfrak{B}}^{\prime\prime}_{2}$ of ${\mathfrak{B}}^{[1]}_{b}$ of preperiods $\leq 6p_{2}$ such that, see Figure 6:
  • $\ell_{1}$ and $\ell_{2}$ separate ${\mathfrak{B}}^{[1]}_{a}$ from $\alpha$ (within ${\mathfrak{K}}_{f}$),
  • the geodesic continuum $\ell\subset{\mathfrak{T}}_{f}$ connecting $\ell_{1},\ell_{2}$ intersects little preperidic bushes ${\mathfrak{B}}_{i}^{[2]},{\mathfrak{B}}^{[2]}_{i+1}$of level 22 22 with preperiods less than $6p_{1}$;
  • ${\mathfrak{B}}_{i}^{[2]},{\mathfrak{B}}^{[2]}_{i+1}$ are disjoint from $\ell_{1}\cup\ell_{2}$.
We remark that $\ell,{\mathfrak{B}}_{i}^{[2]},{\mathfrak{B}}^{[2]}_{i+1}$ are within ${\mathfrak{B}}^{[1],(6p_{2})}_{a}$.
Proof.
Let ${\mathfrak{B}}_{\gamma}^{[2]}\subset{\mathfrak{B}}^{[1]}_{a}$ be the level 22 22 periodic bush closest to $\alpha$; i.e. $({\mathfrak{T}}_{f}\cap{\mathfrak{B}}^{[1]}_{a})\setminus{\mathfrak{B}}^{[2]}_{\gamma}$ and $\alpha$ are in different components of ${\mathfrak{T}}_{f}\setminus{\mathfrak{B}}_{\gamma}^{[2]}$. Let the ${\mathfrak{B}}_{t}^{[2]}\subset{\mathfrak{B}}^{[1],(tp_{1})}_{a}$ be the lifts of ${\mathfrak{B}}_{\gamma}^{[2]}$ under $f^{tp_{1}}$ towards $\alpha$; i.e. each ${\mathfrak{B}}_{t}^{[2]}$ separate $\alpha$ from ${\mathfrak{B}}^{[1],(tp_{1}-p_{1})}_{a}$.
Since $f^{6p_{2}}$ has critical points in ${\mathfrak{B}}^{[2]}_{1}$ and in ${\mathfrak{B}}_{4}^{[2]}$, we can select $\ell_{1}$ and $\ell_{2}$ passing through these critical points such that $\ell_{1},\ell_{2}$ connect preperiodic lifts ${\mathfrak{B}}^{\prime}_{1},{\mathfrak{B}}^{\prime\prime}_{1}$ and ${\mathfrak{B}}^{\prime}_{2},{\mathfrak{B}}^{\prime\prime}_{2}$ of ${\mathfrak{B}}^{[1]}_{b}$. Then $\ell_{1},\ell_{2}$ separate ${\mathfrak{B}}^{[1]}_{a}$ from $\alpha$, and $\ell_{1}\cup\ell_{2}$ separate ${\mathfrak{B}}^{[2]}_{2},{\mathfrak{B}}^{[2]}_{3}$ from $\alpha$ and ${\mathfrak{B}}^{[1]}_{a}$; i.e. we can take ${\mathfrak{B}}^{[2]}_{i},{\mathfrak{B}}^{[2]}_{i+1}$ to be ${\mathfrak{B}}^{[2]}_{2},{\mathfrak{B}}^{[2]}_{3}$. ∎
2.4. Invariant arc diagrams
In this subsection, we will discuss invariant up to homotopy arc diagrams of ql maps. Arc diagrams endowed with weights will naturally appear from the thin-thick decompositions of the dynamical planes of $\psi^{\bullet}$-ql maps, see §3.3, §3.5.
2.4.1. Arc diagrams
Consider a hyperbolic open Riemann surface $S$ of finite type without cusps. We endow $S$ with its ideal boundary $\partial^{i}S$. This naturally makes $V\coloneqq\overline{S}\equiv S\cup\partial^{i}S$ a compact surface.
A path (closed or open) $\ell$ in $S$ is an embedded (closed or open) interval $\ell\colon I\to S$.
An open path $\gamma\colon(0,1)\to S$ is proper if it extends to $\gamma\colon[0,1]\to V\equiv S\cup\partial^{i}S$ with $\gamma\{0,1\}\subset\partial^{i}S$. Two proper paths $\gamma_{0},\gamma_{1}$ in $S$ are homotopic if there is a homotopy $\gamma_{t}$ among proper paths. Similarly, proper curves and their homotopy are defined.
An arc on $S$ is a class of properly homotopic paths, $\alpha=[\gamma]$. A curve $\gamma$ is trivial if it can be represented in an arbitrary small neighborhood of $\partial^{i}S$. Two different arcs are non-crossing if they can be represented by non-crossing paths.
An arc diagram (AD) is a family of non-trivial pairwise non-crossing arcs $A=\{\alpha_{i}\}$. A weighted arc diagram (WAD) ${\mathcal{A}}=\sum_{\alpha_{i}\in A}w_{i}\alpha_{i},\ w_{i}\in{\mathbb{R}}_{>0}$ is an arc diagram endowed with positive weights.
2.4.2. Arc diagrams of ql maps
Consider a ql map $f\colon X\to Y$ and assume that it is $n+1$ times DH renormalizable.
An arc diagram $A=\{\alpha_{i}\}$ on $X\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ is horizontal if every $\alpha_{i}$ connects components of ${\boldsymbol{\mathfrak{B}}}^{[n]}$. A horizontal arc diagram $A=\{\alpha_{i}=[\gamma_{i}]\}$ is called invariant if every $\alpha_{i}$ can be represented up to a proper homotopy in $X\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ in an arbitrary small neighborhood of $f^{-1}\left({\boldsymbol{\mathfrak{B}}}^{[n]}\cup\Gamma\right)$, where $\Gamma=\bigcup_{i}\gamma_{i}$. In other words, every $[\gamma]\in A$ can be presented as a concatenation
(2.10) $\gamma=\ell_{0}\#\gamma_{1}\#\ell_{1}\#\gamma_{2}\#\dots\#\gamma_{s}\#\ell_{s},$
where
  • $[\gamma_{j}]\in f^{*}(A)$ are proper arcs in $X\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}=f^{-1}\left(Y\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}\right)$; and
  • every component of $\ell_{i}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}$ is trivial in $X\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}$ (with respect to a proper homotopy, see §2.4.1).
(In other words, the $\ell_{i}$ are contractible into ${\boldsymbol{\mathfrak{B}}}^{[n],(1)}$.) Figure 7 gives an example of an invariant arc diagram.
$f$${\mathfrak{B}}^{[1]}_{0}$$\alpha$$\beta$${\mathfrak{B}}^{[1],(1)}_{0}$$\tilde{\alpha}$$\tilde{\beta}$
Figure 7. The arc diagram $\{[\alpha],[\beta]\}$ is invariant. Here $\widetilde{\alpha},\widetilde{\beta}$ are lifts of $\alpha,\beta$, the map $f$ is the Rabbit map tuned with the Feigenbaum map, peripheral lifts of $\alpha,\beta$ and preperiodic bushes are omitted on the right side. The arc diagram is invariant because $\alpha$ can be properly homotoped into $\widetilde{\beta}$ and $P\coloneqq{\mathfrak{B}}^{[1],(1)}_{0}\setminus{\mathfrak{B}}^{[1]}_{0}$ while $\beta$ can be properly homotoped into $\widetilde{\beta},P,\widetilde{\alpha}$.
Lemma 2.2 (Alignment with ${\mathfrak{T}}_{f}$, following [K, §4]).
If $A=\{\alpha_{i}\}$ is an invariant horizontal AD on $X\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$, where $f\colon X\to Y$ is a ql map, then $A$ is aligned with the Hubbard continuum ${\mathfrak{T}}_{f}$: every $\alpha_{i}$ can be represented in an arbitrary small neighborhood of a geodesic continuum $T_{i}\subset{\mathfrak{T}}_{f}\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ connecting components of ${\boldsymbol{\mathfrak{B}}}^{[n]}$.
Proof.
Write $\alpha_{i}=[\gamma_{i}]$ and consider $Y^{\prime}\coloneqq Y\setminus\left({\boldsymbol{\mathfrak{B}}}^{[n]}\cup\Gamma\right)$, where $\Gamma=\bigcup_{i}\gamma_{i}$. Then one of the little bushes ${\mathfrak{B}}^{[n]}_{i}$ is accessible from the outermost component of $\partial Y^{\prime}$. Therefore, we can select a proper path $\ell\subset Y^{\prime}$ connecting $\partial Y$ and ${\boldsymbol{\mathfrak{B}}}^{n}$. For every such a path $\ell$, its legal pullback $\widetilde{\ell}$ is any of its lift connecting ${\boldsymbol{\mathfrak{B}}}^{[n]}$ and $\partial X$ and concatenated by a path in $Y\setminus X$ so that ~$\widetilde{\ell}\subset Y\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ is a proper path connecting $\partial Y$ and ${\boldsymbol{\mathfrak{B}}}^{n}$. Since $A$ is invariant, $\widetilde{\ell}$ can be again represented in $Y^{\prime}$ up to a proper homotopy in $Y\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$.
It is well-known that iterated pullbacks of $\ell$ can realize all periodic rays landing at ${\boldsymbol{\mathfrak{B}}}^{[n]}$ up to a proper homotopy in $Y\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$. (It is sufficient to verify the property for the superattracting model §2.3.2.) Therefore, $A$ is aligned with ${\mathfrak{T}}_{f}$. ∎
2.4.3. Genuine periodic arcs
Let $A=\{\alpha_{i}\}$ be a horizontal invariant AD on $Y\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ of $f\colon X\to Y$. For every $m\geq 1$, we can present $[\gamma]\in A$ as a concatenation
(2.11) $\gamma=\ell_{0}\#\gamma_{1}\#\ell_{1}\#\gamma_{2}\#\dots\#\gamma_{s}\#\ell_{s}$
(compare with (2.10)), where
  • $[\gamma_{j}]\in\big{(}f^{m}\big{)}^{*}(A)$ are proper arcs in $f^{-m}(Y)\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}=f^{-m}\left(Y\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}\right)$; and
  • every component of $\ell_{i}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ is trivial in $X\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ (with respect to a proper homotopy, see §2.4.1) .
We call (2.11) the decomposition of $\gamma$ rel ${\boldsymbol{\mathfrak{B}}}^{[n],(m)}$. We define the expansivity number
[…]
where the minimum is taking over all paths possible $\gamma$ as above.
We call an arc $\alpha\in A$ a genuine periodic if ${\operatorname{EN}}_{A}(f^{m},\alpha)=1$ for all $m\geq 1$.
$f$${\mathfrak{B}}^{[1]}_{0}$$\alpha_{1}$$\alpha_{2}$$\alpha_{3}$$\alpha_{4}$$\beta$${\mathfrak{B}}^{[1],(1)}_{0}$$\tilde{\beta}$$\tilde{\alpha}_{2}$$\tilde{\alpha}_{3}$$\tilde{\alpha}_{4}$$\tilde{\alpha}_{1}$
Figure 8. An invariant arc diagram $A=\{\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\beta\}$ is depicted for the $1/4$ Rabbit map tuned with the Feigenbaum map. Here, ${\operatorname{EN}}_{A}(f,\alpha_{i})=1$ while ${\operatorname{EN}}_{A}(f,\beta)=2$ because $\beta$ must travel through $\tilde{\alpha}_{2}$ and $\tilde{\alpha}_{1}$ ($\beta$ can not cross-intersect its lift $\widetilde{\beta}$). Note that if we replace in $A$ the arc $\beta$ with its orthogonal arc $\beta^{\perp}$ connecting ${\mathfrak{B}}^{[1]}_{0}$ and ${\mathfrak{B}}^{[1]}_{2}$, then the new arc diagram $A^{\operatorname{new}}$ will not be invariant: $\widetilde{\beta}^{\perp}=\beta$ will completely block $\beta^{\perp}$.
Lemma 2.3 (Expansivity of $f\mid{\big{[}{\mathfrak{T}}_{f}\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}\big{]}}$).
Consider an invariant AD $A$ and an arc $\gamma\in A$ aligned with a proper geodesic continuum $T_{\gamma}$ of ${\mathfrak{T}}_{f}\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$. Then
  • $\gamma$ is genuine periodic if and only if $T_{\gamma}$ connects two neighboring bushes ${\mathfrak{B}}^{[n]}_{a},{\mathfrak{B}}^{[n]}_{a+1}$ (with respect to the cyclic order) of a periodic satellite flower ${\mathfrak{B}}^{[n-1]}_{j}\supset{\mathfrak{B}}^{[n]}_{a},{\mathfrak{B}}^{[n]}_{a+1}$ of level $n-1$. In particular, the $n$-th renormalization of $f$ is satellite.
  • if $\gamma$ is not genuine periodic, then
    (2.12) ${\operatorname{EN}}_{A}(f^{p_{n}},\gamma)\geq 2.$
See Figure 8 for illustration.
Proof.
Clearly, if $T_{\gamma}$ is in a satellite flower ${\mathfrak{B}}^{[n-1]}_{j}$ and connects two neighboring bushes ${\mathfrak{B}}^{[n]}_{a},\ {\mathfrak{B}}^{[n]}_{a+1}$, then $\gamma$ is genuine periodic. It is also clear that (2.12) holds unless $T_{\gamma}$ is in a satellite flower ${\mathfrak{B}}^{[n-1]}_{j}$. (It can be easily proven for a superattracting model §2.3.2.)
Assume $T_{\gamma}$ is in a satellite flower ${\mathfrak{B}}^{[n-1]}_{j}$ but connects two non-neighboring bushes ${\mathfrak{B}}^{[n]}_{a},\ {\mathfrak{B}}^{[n]}_{b},\ |a-b|>1$. Then a certain lift $T^{\prime}$ of $T_{\gamma}$ under $f^{m}$ with $m=kp_{n-1}<p_{n}$ cross-intersects $T_{\gamma}$. It follows that $\big{(}f^{m}\big{)}^{*}A$ has an arc aligned with $T^{\prime}$ but has no arc aligned with $T_{\gamma}\setminus\left({\mathfrak{B}}^{[n],(m)}_{a}\cup{\mathfrak{B}}^{[n],(m)}_{b}\right)$. This implies that ${\operatorname{EN}}_{A}(f^{m},\gamma)\geq 2$.
2.4.4. Arcs inside and outside of satellite flowers
Consider an invariant AD $A$ on $X\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$. Consider an arc $\alpha\in A$ aligned with a geodesic continuum $T_{\alpha}\subset{\mathfrak{T}}_{f}\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$. We say that $\alpha$ is in ${\mathfrak{B}}^{[n-1]}_{c}$ if $T_{\alpha}\subset{\mathfrak{B}}^{[n-1]}_{c}$.
Lemma 2.4.
Consider an invariant arc diagram $A$ on $X\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ and an arc $\alpha\in A$. Assume that $\alpha$ is not in any satellite flower ${\mathfrak{B}}^{[n-1]}_{c}$ of level $n-1$. Then there are two arcs
$\alpha_{1},\alpha_{2}\in(f^{2p_{n}})^{*}(A)\hskip 17.07182pt\text{ with }{\ \ }f^{2p_{n}}(\alpha_{1})=f^{2p_{n}}(\alpha_{2})=\alpha_{\operatorname{new}}\in A$
such that $\alpha$ overflows $\alpha_{1},\alpha_{2}$ in the following sense: any decomposition 2.11 of $\alpha=[\gamma]$ rel ${\boldsymbol{\mathfrak{B}}}^{[n],(2p_{n})}$ contains two paths $\gamma_{a},\gamma_{b}$ representing $\alpha_{1}=[\gamma_{a}]$ and $\alpha_{2}=[\gamma_{b}]$.
Proof.
Since $T_{\alpha}$ is not in any satellite flower ${\mathfrak{B}}^{[n-1]}_{c}$, there is a strictly preperiodic component ${\mathfrak{B}}^{[n],(p_{n})}_{d}$ of ${\boldsymbol{\mathfrak{B}}}^{[n],(p_{n})}$ intersecting $T_{\alpha}$. There are two components $T_{1},T_{2}\subset T_{\alpha}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(2p_{n})}$ adjacent to ${\mathfrak{B}}^{[n],(2p_{n})}_{d}$. Since there is an $m\leq 2p_{n}$ such that $f^{m}\big{(}{\mathfrak{B}}^{[n],(2p_{n})}_{d}\big{)}$ contains the critical value, $T_{1},T_{2}$ have a common injective image of generation $m\leq 2p_{n}$. There must be two arcs $\alpha_{1},\alpha_{2}\in(f^{2p_{n}})^{*}(A)$ aligned with $T_{1},T_{2}$. ∎
Assume that the $n$th renormalization is satellite and consider a satellite flower ${\mathfrak{B}}^{[n-1]}_{c}$. Let $A$ be an invariant AD as above. Let $A_{c}=A({\mathfrak{B}}^{[n-1]}_{c})\subset A$ be the AD consisting of all arcs from $A$ that are in ${\mathfrak{B}}^{[n-1]}_{c}$. Similarly, $A^{(p_{n})}_{c}\subset(f^{p_{n}})^{*}(A)$ be the AD consisting of all arcs from $(f^{p_{n}})^{*}(A)$ that are in ${\mathfrak{B}}^{[n-1]}_{c}$. Then
(2.13) $(f^{p_{n}})^{*}\colon A_{c}^{p_{n}}\overset{1:1}{\longrightarrow}A_{c}$
is a bijection and every arc $\alpha\in A_{c}$ is homotopic to its unique lift $\tilde{\alpha}\in A_{c}^{p_{n}},(f^{p_{n}})^{*}(\tilde{\alpha})=\alpha$ in the following sense: $T_{\alpha}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(p_{n})}=T_{\tilde{\alpha}}$.
2.4.5. Inhomogeneous configuration
The discussion here will be only used in Section 7. Consider a ql map $f\colon X\to Y$ with a cycle of bushes ${\boldsymbol{\mathfrak{B}}}^{[1]}$. Write ${\boldsymbol{\Upsilon}}\coloneqq{\boldsymbol{\mathfrak{B}}}^{[1]}\cup{\mathfrak{K}}^{[1]}_{0}$. Note that $f$ does not permute components of ${\boldsymbol{\Upsilon}}$.
An AD $A=\{\alpha_{i}\}$ on $X\setminus{\boldsymbol{\Upsilon}}$ is horizontal if every $\alpha_{i}$ connects components of ${\boldsymbol{\Upsilon}}$. A horizontal AD $A=\{\alpha_{i}=[\gamma_{i}]\}$ is $f^{p_{1}}$-invariant if every $\alpha_{i}$ can be represented up to a proper homotopy in $X\setminus{\boldsymbol{\Upsilon}}$ in an arbitrary small neighborhood of $f^{-p_{1}}\left({\boldsymbol{\Upsilon}}\cup\Gamma\right)$, where $\Gamma=\bigcup_{i}\gamma_{i}$. By replacing $f$ with its superattracting model (§2.3.2, compare also wit Lemma 2.2), we obtain:
Lemma 2.5 (Alignment with ${\mathfrak{T}}_{f}$).
If $A=\{\alpha_{i}\}$ is an $f^{p_{1}}$-invariant AD on $Y\setminus{\boldsymbol{\Upsilon}}$ as above, then $A$ is aligned with the Hubbard continuum ${\mathfrak{T}}_{f}$: every $\alpha_{i}$ can be represented in an arbitrary small neighborhood of a geodesic continuum $T_{i}$ of ${\mathfrak{T}}_{f}\setminus{\boldsymbol{\Upsilon}}$ connecting components of ${\boldsymbol{\Upsilon}}.$
For $m\geq 1$ and $\alpha\in A$, the expansivity number ${\operatorname{EN}}_{A}(f^{mp_{1}},\alpha)$ is defined in the same way as in § 2.4.3; i.e., it is the smallest number of arcs in $f^{-p_{1}m}\left(A\right)$ overflown by $\alpha$. An arc $\alpha\in A$ is genuine periodic if ${\operatorname{EN}}_{A}(f^{p_{1}m},\alpha)=1$ for all $m\geq 1$. By the same argument as in Lemma 2.3, we have:
Lemma 2.6 (Expansivity of $f\mid[{\mathfrak{T}}_{f}\setminus{\boldsymbol{\Upsilon}}]$).
Consider an $f^{p_{1}}$-invariant AD $A$ on $Y\setminus{\boldsymbol{\Upsilon}}$ and an arc $\gamma\in A$ aligned with a proper geodesic continuum $T_{\gamma}$ of ${\mathfrak{T}}_{f}\setminus{\boldsymbol{\Upsilon}}$. If the first renormalization of $f$ is primitive, then
(2.14) ${\operatorname{EN}}_{A}(f^{p_{1}},\gamma)\geq 2.$
2.5. Teichmüller contraction
The Teichmüller contraction comes from the observation that the restriction of a qc conjugacy to deeper renormalization levels can only decrease the dilatation. Therefore, the renormalization orbits (for bounded-type Feigenbaum families) can escape to infinity with at most linear rate. This fact can be traced back to Sullivan [S1], where structures of the Teichmüller spaces (reminiscent to the Thurston’s machinery) were brought into the Renormalization Theory. The proposition below is stated for quadratic maps; the Straightening Theorem easily extends it quadratic-like maps.
Proposition 2.7.
For every combinatorial bound $\bar{p}$, there is a constant $\Delta=\Delta_{\overline{p}}>1$ such that the following holds. Let $f_{c}(z)=z^{2}+c$ be an infinitely renormalizable quadratic polynomial of bounded type $\overline{p}$, see §2.2.5. Then $g_{n}={\mathcal{R}}^{n}f_{c}$ has ql prerenormalization $g_{n}=f_{c}^{p_{n}}\colon X_{n}\to Y_{n}$, $X_{n}\Subset Y_{n}$ such that
  1. (I)
    ${\mathcal{W}}(Y_{n}\setminus X_{n})=O(\Delta^{n})$; and
  2. (II)
    $Y_{n}\setminus{\mathfrak{K}}_{g_{n}}$ is disjoint from ${\boldsymbol{\mathfrak{B}}}^{[n]}_{f_{c}}$.
Proof.
Let ${\mathcal{M}}_{i},i\in I$ be the (finite) set of all maximal ${\mathcal{M}}$-copies in ${\mathcal{M}}$ of period $\leq\bar{p}$, and let
$R\colon\bigcup_{i\in I}{\mathcal{M}}_{i}\to{\mathcal{M}}$
be the ql straightening map. Let us remove from every satellite ${\mathcal{M}}_{i}$ a small open neighborhood of its cusp such that the remaining ${\mathcal{M}}^{\circ}_{i}$ contains all the preimages of the ${\mathcal{M}}_{j}$ under $R\colon{\mathcal{M}}_{i}\to{\mathcal{M}}$. For a primitive copy ${\mathcal{M}}_{i}$, set ${\mathcal{M}}_{i}^{\circ}\coloneqq{\mathcal{M}}_{i}$. By compactness, there is a $\Delta>1$ such that every $f_{c}$ with $c\in\bigcup_{i}{\mathcal{M}}_{i}^{\circ}$ has a ql restriction $g_{1}=f_{c}^{p_{1}}\colon X_{1,c}\to Y_{1,c}$ around its maximal little Julia set containing 00 0 such that $g_{1}$ satisfies (II) and such that $g_{1}$ is hybrid conjugate via $h_{c}\colon Y_{1,1}\to Y_{R(c)}$ with dilatation $\leq\Delta$ to a ql restriction $f_{R(c)}\colon X_{R(c)}\to Y_{R(c)}$. Moreover, we can select the $Y_{1,c}$ and $Y_{c}$ so that $Y_{c}\Supset Y_{1,c}$.
Wright $c_{n}\coloneqq R^{n}(c)$. Then
$\bar{h}\coloneqq\big{(}h_{c_{n-1}}\circ\dots\circ h_{c_{1}}\circ h_{c}\big{)}^{-1}$
is a hybrid conjugacy from $f_{c_{n}}\colon X_{c_{n}}\to Y_{c_{n}}$ to a ql restriction $g_{n}=f^{p_{1}\dots p_{n}}_{c}\colon X_{n,c}\to Y_{n,c}$ satisfying (II). Since the dilatation of $\overline{h}$ is $\leq\Delta^{n}$, Property (I) follows. ∎
Remark 2.8.
Property (II) implies that the prerenormalization $g_{n}\colon X_{n}\to Y_{n}$ is unbranched: ${\mathfrak{P}}(f_{c})\cap Y_{n}\subset{\mathfrak{K}}(g_{n})$, compare with [McM2]. Moreover, by induction, $Y_{n}$ is disjoint from
$\Upsilon\coloneqq\bigcup_{m\leq n}\left({\boldsymbol{\mathfrak{B}}}^{[m]}_{f}\setminus{\mathfrak{B}}^{[m]}_{f,0}\right).$
This implies that the quadratic-like prerenormalizations $g_{n}\colon X_{n}\to Y_{n}$ can be univalently lifted to the dynamical plane of any $\psi^{\bullet}$ renormalization
${\mathcal{R}}^{n_{1}\bullet}\circ{\mathcal{R}}^{n_{2}\bullet}\dots\dots\circ{\mathcal{R}}^{n_{s}\bullet}(f_{c})$
of $f_{c}$, see §3.4.3.
3. $\psi^{\bullet}$-ql renormalization and near-degenerate regime
In this section, we will introduce $\psi^{\bullet}$-ql renormalization and discuss tools to detect its degeneration; see [A, K, KL1, L3, DL2] for details.
Given a compact subset $K\Subset S$, we denote by ${\mathcal{F}}(S,K)$ the family of non-trivial proper curves in $S\setminus K$ emerging from $\partial S$. We write
(3.1) ${\mathcal{W}}(S,K)={\mathcal{W}}\big{(}{\mathcal{F}}(S,K)\big{)}.$
3.1. Outline
Consider a ql map $f\colon X\to Y$ and let ${\boldsymbol{\mathfrak{B}}}^{[n]}$ be its level $n$ cycle of little bushes.
Around every periodic ${\mathfrak{K}}^{[n]}_{i}$ there is an associated ql prerenormalization
(3.2) $f_{n,i}=f^{p_{n}}\colon X_{n,i}\to Y_{n,i}.$
In §3.4, we will define the $\psi^{\bullet}$-ql renormalization
(3.3) $F_{n,i}=(f_{n,i},\ \iota_{n,i})\colon U_{n,i}\rightrightarrows V_{n,i}$
of $f\colon X\to Y$ by extending (3.2) along all curves in $Y\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$, see §3.4.3. This is similar to the $\psi$-ql renormalization [K], where the extension is performed along all curves in $Y\setminus{\boldsymbol{\mathfrak{K}}}^{[n]}$ under the assumption that the ${\mathfrak{K}}^{[n]}_{i}$ are pairwise disjoint (the primitive case); see §3.4.1 and Remark 3.5.
The correspondence $F_{n,i}$ in (3.3) is called a $\psi^{\bullet}$-ql map. It consists of a degree 22 22 branched covering $f_{n,i}$ and an immersion $\iota_{n,i}$, see definitions in §3.4.2. We define the Julia set and bush of $F_{n,i}$ to be that of the ql map $f_{n,i}$ (3.2):
${\mathfrak{K}}_{F_{n,i}}\coloneqq{\mathfrak{K}}_{f_{n,i}}{\ \ }{\ \ }\text{ and }{\ \ }{\ \ }{\mathfrak{B}}_{F_{n,i}}\coloneqq{\mathfrak{B}}_{f_{n,i}}.$
Similarly, ${\mathfrak{K}}_{F}$ and ${\mathfrak{B}}_{F}$ are defined for any $\psi^{\bullet}$-ql map, see (3.14).
A $\psi^{\bullet}$-ql renormalization can be naturally defined for a $\psi^{\bullet}$-ql map; i.e., $\psi^{\bullet}$-ql renormalization can be iterated. Sup-Chain Rule for renormalization domains is stated in §3.4.4.
Just like a ql map $f\colon X\to Y$ can be restricted to $f\colon X^{k+1}\to X^{k}$, where $X^{k+1}=f^{-k}(X)$, a $\psi^{\bullet}$-ql map $F=(f,\iota)\colon U\rightrightarrows V$ can be restricted to
$F=(f,\iota)\colon U^{k+1}\rightrightarrows U^{k}$
by considering the fiber product §3.4.6. We have a natural $1\displaystyle 2^{k}:$ 1 branched covering $f^{k}\colon U^{k}\to V$ and an immersion $\iota^{k}\colon U^{k}\to V$.
The width of a $\psi^{\bullet}$-ql map $F\colon U\rightrightarrows V$ is
${\mathcal{W}}_{\bullet}(F)\coloneqq{\mathcal{W}}(V\setminus{\mathfrak{B}}_{F}).$
We denote by ${\mathcal{A}}^{[n]}\equiv{\mathcal{A}}^{[n]}_{F}$ the weighted arc diagram (WAD) of $V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$: it is a formal sum of weighted arcs representing rectangles in the thick-thin decomposition of $V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$; see §3.3 and Figure 3.3. The horizontal part ${\mathcal{A}}^{[n]}_{\mathrm{hor}}$ of ${\mathcal{A}}^{[n]}$ consists of weighted arcs connecting components of ${\boldsymbol{\mathfrak{B}}}^{[n]}$. The vertical part ${\mathcal{A}}^{[n]}_{\mathrm{ver}}$ of ${\mathcal{A}}^{[n]}$ consists of weighted arcs connecting $\partial V$ and components of ${\boldsymbol{\mathfrak{B}}}^{[n]}$. The local WAD ${\mathcal{A}}^{[n]}_{i}$ consists of arcs in ${\mathcal{A}}^{[n]}$ adjacent to ${\mathfrak{B}}^{[n]}_{i}$, where the weight of self-arcsto ${\mathfrak{B}}^{[n]}_{i}$ is doubled. More generally, ${\mathcal{A}}^{[n],(m),k}$ is the WAD of $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$, and the WADs ${\mathcal{A}}^{[n],(m),k}_{\mathrm{hor}},{\mathcal{A}}^{[n],(m),k}_{i},\dots$ are defined accordingly, see §3.5.2. We have
${\mathcal{W}}(F_{n,i})={\mathcal{W}}\left({\mathcal{A}}^{[n]}_{i}\right)+O_{p_{n}}(1).$
By identifying up to homotopy $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ and $U^{k+1}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$, we obtain ${\mathcal{A}}^{[n],(m),k+1}_{\mathrm{hor}}\leq{\mathcal{A}}^{[n],(m),k}_{\mathrm{hor}}$, see §3.5.6. Since the complexity of $A^{[n],k}_{\mathrm{hor}}\equiv{\operatorname{AD}}\left[{\mathcal{A}}^{[n],k}_{{\mathrm{hor}}}\right]$ decreases, the $A^{[n],k}_{{\mathrm{hor}}}$ are essentially invariant for $k\geq 3p_{n}$. Then either most of ${\mathcal{A}}^{[n]}_{\mathrm{hor}}$ is in ${\mathcal{A}}^{[n],k}_{\mathrm{hor}}$ or a substantial part of ${\mathcal{A}}^{[n]}_{\mathrm{hor}}$ is in ${\mathcal{A}}^{[n],k}_{\mathrm{ver}}$ – this is a key dichotomy in Section 4; see also the dichotomy “(I) vs (II)” in §1.3.2.
3.2. Rectangles
A Euclidean rectangle is a rectangle $E_{x}\coloneqq[0,x]\times[0,1]\subset{\mathbb{C}}$, where:
  • $(0,0),(x,0),(x,1),(0,1)$ are four vertices of $E_{x}$,
  • $\partial^{h}E_{x}=[0,x]\times\{0,1\}$ is the horizontal boundary of $E_{x}$,
  • $\partial^{h,0}E_{x}\coloneqq[0,x]\times\{0\}$ is the base of $E_{x}$,
  • $\partial^{h,1}E_{x}\coloneqq[0,x]\times\{1\}$ is the roof of $E_{x}$,
  • $\partial^{v}E_{x}=\{0,x\}\times[0,1]$ is the vertical boundary of $E_{x}$,
  • ${\mathcal{F}}(E_{x})\coloneqq\{\{t\}\times[0,1]\mid t\in[0,x]\}$ is the vertical foliation of $E_{x}$,
  • ${\mathcal{F}}^{\operatorname{full}}(E_{x})\coloneqq\{\gamma\colon[0,1]\to E_{x}\mid\gamma(0)\in\partial^{h,0}E_{x},\ \gamma(1)\in\partial^{h,1}E_{x}\}$ is the full family of curves in $E_{x}$;
  • ${\mathcal{W}}(E_{x})={\mathcal{W}}({\mathcal{F}}(E_{x}))={\mathcal{W}}\big{(}{\mathcal{F}}^{\operatorname{full}}(E_{x})\big{)}=x$ is the width of $E_{x}$,
  • $\operatorname{mod}(E_{x})=1/{\mathcal{W}}(E_{x})=1/x$ the extremal length of $E_{x}$.
By a (topological) rectangle in a Riemann surface we mean a closed Jordan disk ${\mathcal{R}}$ together with a conformal map $h\colon{\mathcal{R}}\to E_{x}$ into the standard rectangle $E_{x}$. The vertical foliation ${\mathcal{F}}({\mathcal{R}})$, the full family ${\mathcal{F}}^{\operatorname{full}}({\mathcal{R}})$, the base $\partial^{h,0}{\mathcal{R}}$, the roof $\partial^{h,1}{\mathcal{R}}$, the vertices of ${\mathcal{R}}$, and other objects are defined by pulling back the corresponding objects of $E_{x}$. Equivalently, a rectangle ${\mathcal{R}}$ is a closed Jordan disk together with four marked vertices on $\partial{\mathcal{R}}$ and a chosen base between two vertices.
A genuine subrectangle of $E_{x}$ is any rectangle of the form $E^{\prime}=[x_{1},x_{2}]\times[0,1]$, where $0\leq x_{1}<x_{2}\leq x$; it is identified with the standard Euclidean rectangle $[0,x_{2}-x_{1}]\times[0,1]$ via $z\mapsto z-x_{1}$. A genuine subrectangle of a topological rectangle is defined accordingly.
A subrectangle of a rectangle ${\mathcal{R}}$ is a topological rectangle ${\mathcal{R}}_{2}\subset{\mathcal{R}}$ such that $\partial^{h,0}{\mathcal{R}}_{2}\subset{\mathcal{R}}$ and $\partial^{h,1}{\mathcal{R}}_{2}\subset{\mathcal{R}}$. By monotonicity: ${\mathcal{W}}({\mathcal{R}}_{2})\leq{\mathcal{W}}({\mathcal{R}})$.
Assume that ${\mathcal{W}}(E_{x})>2$. The left and right 11 11-buffers of $E_{x}$ are defined
$B^{\ell}_{1}\coloneqq[0,1]\times[0,1]{\ \ }\text{ and }{\ \ }B^{\rho}_{1}\coloneqq[x-1,x]\times[0,1]$
respectively. We say that the rectangle
$E^{\operatorname{new}}_{x}\coloneqq[1,x-1]\times[0,1]=E_{x}\setminus\big{(}B^{\ell}_{1}\cup B^{\rho}_{1}\big{)}$
is obtained from $E_{x}$ by removing 11 11-buffers. If ${\mathcal{W}}(E_{x})\leq 2$, then we set $E_{x}^{\operatorname{new}}\coloneqq\emptyset$. Similarly, buffers of any width are defined.
3.2.1. Monotonicity and Grötzsch inequality
We say a family of curves $SS\displaystyle\SS$ overflows a family ${\mathcal{G}}$ if every curve $SS\displaystyle\gamma\in\SS$ contains a subcurve $\gamma^{\prime}\in{\mathcal{G}}$. We also say that
  • a family of curves ${\mathcal{F}}$ overflows a rectangle ${\mathcal{R}}$ if ${\mathcal{F}}$ overflows ${\mathcal{F}}^{\operatorname{full}}({\mathcal{R}})$;
  • a rectangle ${\mathcal{R}}_{1}$ overflows another rectangle ${\mathcal{R}}_{2}$ if ${\mathcal{F}}({\mathcal{R}}_{1})$ overflows ${\mathcal{F}}^{\operatorname{full}}({\mathcal{R}}_{2})$.
If ${\mathcal{F}}$ overflows a family or a rectangle ${\mathcal{G}}$, then ${\mathcal{G}}$ is wider than ${\mathcal{F}}$:
(3.4) ${\mathcal{W}}({\mathcal{F}})\leq{\mathcal{W}}({\mathcal{G}}).$
If ${\mathcal{F}}$ overflows both ${\mathcal{G}}_{1},{\mathcal{G}}_{2}$, and ${\mathcal{G}}_{1},{\mathcal{G}}_{2}$ are disjointly supported, then the Grötzsch inequality states:
(3.5) ${\mathcal{W}}({\mathcal{F}})\leq{\mathcal{W}}({\mathcal{G}}_{1})\oplus{\mathcal{W}}({\mathcal{G}}_{2}),$
where $x\oplus y=(x^{-1}+y^{-1})^{-1}$ is the harmonic sum.
3.3. Weighted arc diagrams
Let us recall the notion of the Weighted Arc Diagram (WAD) describing wide rectangles in the thick-thin decomposition of a Riemann surface with boundary, see Figure 9 for illustration and brief summary. We also recall that WAD were abstractly defined in §2.4.1.
$V$$X$$Z$$Y$$X$$Z$$V$$Y$${\boldsymbol{\Pi}}^{\operatorname{can}}_{\alpha_{i}}$${\boldsymbol{\Pi}}^{\operatorname{can}}_{\alpha_{j}}$
Figure 9. Top: the thin-thick decomposition of $V\setminus(X\cup Y\cup Z)$. As we are interested in homotopy properties of wide rectangles, they will usually be depicted as narrow rectangles (bottom) and denoted by ${\boldsymbol{\Pi}}^{\operatorname{can}}_{\alpha_{i}}$, where $\alpha_{i}$ is an arc representing the homotopy class of the rectangle ${\boldsymbol{\Pi}}^{\operatorname{can}}_{\alpha_{i}}$. The Weighted Arc Diagram (WAD) is the formal sum ${\mathcal{A}}=\sum_{i}W({\boldsymbol{\Pi}}^{\operatorname{can}}_{\alpha_{i}})\alpha_{i}$, while the Arc Diagram (AD) is the associated set of curves $A=\{\alpha_{i}\}\equiv{\operatorname{AD}}[{\mathcal{A}}]$. Thin annuli of the thick-thin decomposition are disregarded in the WAD.
3.3.1. Proper rectangles
A proper rectangle $R$ in an open hyperbolic surface $S$ is a rectangle in its ideal completion $\overline{S}=S\cup\partial^{i}S$ such that $\partial^{h}R\subset\partial^{i}S$. We naturally view vertical curves of $R$ as proper (and open) paths in $S$. We say that $R$ connects boundary components $J_{0},J_{1}\subset\partial^{i}R$ if $\partial^{h,0}R\subset J_{0},\ \partial^{h,1}R\subset J_{1}$.
If $S=U\setminus K$, where $U$ is a topological disk and $K\Subset U$ is a compact subset with finitely many connected components, then $\partial^{i}S$ has one outer component $\partial^{i}U$; all remaining inner components are parameterized by the components of $K$. If $J_{0},J_{1}$ represent the boundaries of components $K_{0},K_{1}$ of $K$, then we say that the above $R$ connects $K_{0},K_{1}$.
3.3.2. Universal cover
For compact Riemann surface $V$ as in §2.4.1x, write its non-empty boundary as $\partial V=\bigsqcup_{k}J_{k}$. We set $\pi:\mathbb{D}\rightarrow\operatorname{int}V$ to be the universal covering, and we let ${\Lambda}\subset{\mathbb{T}}=\partial\mathbb{D}$ be the limit set for the group ${\Delta}$ of deck transformations of $\pi$. Then $\pi$ extends continuously to ${\mathbb{T}}\setminus{\Lambda}$, and $\pi:\overline{\mathbb{D}}\setminus{\Lambda}\rightarrow V$ is the universal covering of $V$. Moreover, $\pi$ restricted to any component of ${\mathbb{T}}\setminus{\Lambda}$ gives us a universal covering of some component $J_{k}$ of $\partial V$. Such a component of ${\mathbb{T}}\setminus{\Lambda}$ will be denoted ${\mathbf{J}}_{k}$ (usually we will need only one component for each $k$, so we will not use an extra label for it). The stabilizer of ${\mathbf{J}}_{k}$ in ${\Delta}$ is a cyclic group generated by a hyperbolic Möbius transformation $\tau_{k}$ with fixed points in $\partial{\mathbf{J}}_{k}$.
3.3.3. Covering annuli and local weights
We let
(3.6) ${\mathbb{A}}(V,J_{k})\coloneqq\mathbb{D}/<\tau_{k}>$
be the covering annulus of $V$ corresponding to $J_{k}$. We call its width
${\mathcal{W}}(J_{k})\coloneqq{\mathcal{W}}\left[{\mathbb{A}}(V,J_{k})\right]$
the local weight of $J_{k}$. We let
${\mathcal{W}}(V)=\sum_{k}{\mathcal{W}}(J_{k})$
be the total weight of $V$.
3.3.4. Canonical arc diagram
Recall that an arc $\alpha$ is a proper path in $V$ up to proper homotopy. Any arc $\alpha$ connects some components $J_{k}$ and $J_{i}$ of $\partial V$ and lifts to an arc $\alpha$ in ${\mathbb{D}}$ connecting some intervals ${\mathbf{J}}_{k}$ and ${\mathbf{J}}_{i}$. We let ${\mathcal{W}}({{\mbox{\boldmath$\displaystyle\alpha$}}})={\mathcal{W}}_{\overline{\mathbb{D}}}({\mathbf{J}}_{k},{\mathbf{J}}_{i})$ to be the width of all curves in $\mathbb{D}$ connecting ${\mathbf{J}}_{k}$ and ${\mathbf{J}}_{i}$. Then ${\mathcal{W}}({{\mbox{\boldmath$\displaystyle\alpha$}}})$ is also the width of the rectangle ${\boldsymbol{\Pi}}_{\mbox{\boldmath$\displaystyle\alpha$}}=\overline{\mathbb{D}}$ with horizontal sides ${\mathbf{J}}_{k}$ and ${\mathbf{J}}_{i}$. It is independent of the lift used.
If ${\mathcal{W}}({{\mbox{\boldmath$\displaystyle\alpha$}}})>2$, then by removing from ${\boldsymbol{\Pi}}_{\mbox{\boldmath$\displaystyle\alpha$}}$ the square buffers we obtain a rectangle ${\boldsymbol{\Pi}}^{{\operatorname{can}}}_{\mbox{\boldmath$\displaystyle\alpha$}}$. In this case we let
${\mathcal{W}}^{\operatorname{can}}(\alpha)={\mathcal{W}}^{\operatorname{can}}({{\mbox{\boldmath$\displaystyle\alpha$}}})={\mathcal{W}}({\boldsymbol{\Pi}}^{\operatorname{can}}_{\mbox{\boldmath$\displaystyle\alpha$}})={\mathcal{W}}({{\mbox{\boldmath$\displaystyle\alpha$}}})-2.$
By construction, the $\Delta$-orbit of ${\boldsymbol{\Pi}}^{\operatorname{can}}_{\mbox{\boldmath$\displaystyle\alpha$}}$ consists of pairwise disjoint rectangles. Therefore, ${\boldsymbol{\Pi}}^{\operatorname{can}}_{\mbox{\boldmath$\displaystyle\alpha$}}$ projects injectively onto a rectangle ${\boldsymbol{\Pi}}^{\operatorname{can}}_{\alpha}$ in $V$. Moreover, different ${\boldsymbol{\Pi}}^{\operatorname{can}}_{{\mbox{\boldmath$\displaystyle\alpha$}}_{i}}$ project to pairwise disjoint ${\boldsymbol{\Pi}}^{\operatorname{can}}_{\alpha_{i}}$.
If ${\mathcal{W}}({{\mbox{\boldmath$\displaystyle\alpha$}}})\leq 2$ we let ${\mathcal{W}}^{\operatorname{can}}(\alpha)=0$ and ${\boldsymbol{\Pi}}^{\operatorname{can}}(\alpha)\coloneqq\emptyset$. Arcs with positive weight form the canonical weighted arc diagram (WAD) on $V$:
(3.7) ${\mathcal{A}}_{V}\equiv{\operatorname{WAD}}(V)\coloneqq\sum_{{\mathcal{W}}^{\operatorname{can}}(\alpha)>0}{\mathcal{W}}^{\operatorname{can}}(\alpha)\alpha.$
Every ${\boldsymbol{\Pi}}^{\operatorname{can}}(\alpha)$ supports the canonical vertical foliation ${\mathcal{F}}({\boldsymbol{\Pi}}^{\operatorname{can}}(\alpha)$. Altogether these foliations form the canonical lamination of $V$:
(3.8) ${\mathcal{F}}({\mathcal{A}}_{V})\coloneqq\bigsqcup_{\alpha\in{\mathcal{A}}_{V}}{\mathcal{F}}({\boldsymbol{\Pi}}^{\operatorname{can}}(\alpha)),$
We set
(3.9) ${\mathcal{W}}({\mathcal{A}}_{V})\coloneqq\sum_{\alpha\in{\mathcal{A}}_{V}}{\mathcal{W}}(\alpha)={\mathcal{W}}\big{(}{\mathcal{F}}({\mathcal{A}}_{V})\big{)}.$
The canonical arc diagram of $V$ is the set of arcs in ${\mathcal{A}}_{V}$
$A_{V}\coloneqq\{\alpha\mid{\mathcal{W}}^{\operatorname{can}}(\alpha)>0\}\equiv{\operatorname{AD}}[{\mathcal{A}}_{V}].$
3.3.5. Local WADs and the thick-thin decomposition
For a boundary component $J_{k}$ of $\partial V$, its local WAD ${\mathcal{A}}_{J_{k}}$ consists of weighted arcs of ${\mathcal{A}}_{V}$ adjacent to $J_{k}$ such that the weights of self-arcs adjacent to $J_{k}$ are doubled. We have:
$\sum_{J_{k}}{\mathcal{A}}_{J_{k}}=2{\mathcal{A}}_{V}.$
See [L3, §7.6.3] for a reference of the following fact.
Theorem 3.1 (Thin-Thick Decomposition).
For any compact Riemann surface $V$ with boundary $\partial V=\bigsqcup_{k}J_{k}$, we have:
${\mathcal{W}}(J_{k})-C\leq{\mathcal{W}}({\mathcal{A}}_{J_{k}})\leq{\mathcal{W}}(J_{k}){\ \ }{\ \ }\text{ for every }J_{k}$
and
$\sum_{J_{k}}{\mathcal{W}}(J_{k})-C\leq 2{\mathcal{W}}({\mathcal{A}}_{V})\leq\sum_{J_{k}}{\mathcal{W}}(J_{k})={\mathcal{W}}(V),$
where $C$ depends only on the topological type of $V$.
3.3.6. Sub-diagrams and removing buffers
Let ${\mathcal{A}}\equiv{\mathcal{A}}_{V}$ be the canonical WAD of $V$ as in §3.3.4, and let $B\subset A$ be a subset of arcs from $A\equiv{\operatorname{AD}}({\mathcal{A}})$. Then the WAD induced by $B$ is the formal sum ${\mathcal{B}}\coloneqq\sum_{\alpha\in B}{\mathcal{W}}^{\operatorname{can}}(\alpha)\alpha$, compare to (3.7). The weight $W({\mathcal{B}})$ of ${\mathcal{B}}$ and its canonical lamination ${\mathcal{F}}({\mathcal{B}})$ is defined as for ${\mathcal{A}}$, see (3.8) and (3.9). In §3.5.2, we will specify the vertical and horizontal parts of the WAD associated with $\psi^{\bullet}$-ql maps.
For $B,{\mathcal{B}}$ as above and $c>0$, we define
${\mathcal{B}}-c\coloneqq\sum_{\begin{subarray}{c}\alpha\in B\\ {\mathcal{W}}(\alpha)>c\end{subarray}}({\mathcal{W}}^{\operatorname{can}}(\alpha)-c)\alpha,$
$B-c\coloneqq{\operatorname{AD}}({\mathcal{B}}-c);$
i.e., we remove the $c$-weight from each arc. For a rectangle ${\boldsymbol{\Pi}}^{\operatorname{can}}(\alpha)$ as in §3.3.4, we denote by ${\boldsymbol{\Pi}}^{\operatorname{can}}(\alpha)-c$ the rectangle obtained from ${\boldsymbol{\Pi}}^{\operatorname{can}}(\alpha)$ by removing the $c/2$-buffers on each side. We call
${\mathcal{F}}({\mathcal{B}}-c)\coloneqq\bigsqcup_{\alpha\in{\mathcal{B}}-c}{\mathcal{F}}({\boldsymbol{\Pi}}^{\operatorname{can}}(\alpha)-c)$
the canonical lamination of ${\mathcal{B}}-c$. We also denote by
${\mathcal{F}}^{\operatorname{full}}({\mathcal{B}}-c)\coloneqq\bigsqcup_{\alpha\in{\mathcal{B}}-c}{\mathcal{F}}^{\operatorname{full}}({\boldsymbol{\Pi}}^{\operatorname{can}}(\alpha)-c)$
the full family of vertical curves within respective rectangles, see §3.2.
3.3.7. Transformation rules
Lemma 3.2.
Let $i:S^{\prime}\rightarrow S$ be a holomorphic map between two compact Riemann surfaces with boundary. Then
$\mathrm{(i)}$ If a boundary component $J^{\prime}$ of $S^{\prime}$ is mapped with degree $d$ to a boundary component $J$ of $S$ then ${\mathcal{W}}(J^{\prime})\leq d\cdot{\mathcal{W}}(J)$;
$\mathrm{(ii)}$ If an arc $a^{\prime}$ of $S^{\prime}$ is mapped to an arc $\alpha$ of $S$ then ${\mathcal{W}}(\alpha^{\prime})\leq{\mathcal{W}}(\alpha)$.
Lemma 3.3.
Let ${\mathbf{i}}:{\boldsymbol{\Pi}}^{\prime}\rightarrow{\boldsymbol{\Pi}}$ be a holomorphic map between two rectangles which extends to homeomorphisms between the respective horizontal sides ${\boldsymbol{I}}^{\prime}\rightarrow{\boldsymbol{I}}$ and ${\mathbf{J}}^{\prime}\rightarrow{\mathbf{J}}$. Then
${\mathcal{W}}({\boldsymbol{\Pi}}^{\prime})\leq{\mathcal{W}}({\boldsymbol{\Pi}}).$
3.3.8. WAD under covering
Assume that $f\colon U\to V$ as a finite-degree covering between surfaces with boundaries. Then, see [K, Lemma 3.3]
(3.10) ${\mathcal{A}}_{U}=f^{*}{\mathcal{A}}_{V},{\ \ }\text{ where }{\ \ }f^{*}{\mathcal{A}}_{V}=\sum_{\alpha\in{\mathcal{A}}_{V}}W^{\operatorname{can}}(\alpha)f^{-1}(\alpha)$
and $f^{-1}(\alpha)$ is the set of preimages of $\alpha$.
3.4. $\psi^{\bullet}$-ql renormalization
As in §2.2.3, consider a ql map $f\colon X\to Y$ and its ql prerenormalization
(3.11) $f_{n,0}=f^{p_{n}}\colon X_{n}\to Y_{n}.$
3.4.1. Motivation
A major technical issue is that there is no natural choice of $Y_{n}$ in (3.11) so that ${{\mathcal{W}}(Y_{n}\setminus{\mathfrak{K}}_{n})}$ is optimal and satisfies, in particular, (1.2). To handle this problem, a $\psi$-ql renormalization was introduced in [K]: assuming that (3.11) is primitive and $Y_{n}\cap{\boldsymbol{\mathfrak{K}}}^{[n]}={\mathfrak{K}}^{[n]}_{0}$, extend $f_{n}\colon X_{n}\to Y_{n}$ along all curves in $Y_{n}\setminus{\boldsymbol{\mathfrak{K}}}^{[n]}$; the result is a correspondence $F_{n}=(f_{n},\iota_{n})\colon U_{n}\rightrightarrows V_{n}$, where
$V_{n}\coloneqq{\mathbb{A}}\left(Y\setminus{\boldsymbol{\mathfrak{K}}}^{[n]},\ \partial{\mathfrak{K}}^{[n]}_{0}\right)\bigcup_{\partial{\mathfrak{K}}^{[n]}_{0}}{\mathfrak{K}}^{[n]}_{0}$
is the covering annulus (3.6) of $Y\setminus{\boldsymbol{\mathfrak{K}}}^{[n]}$ rel $\partial{\mathfrak{K}}^{[n]}_{0}$ glued with ${\mathfrak{K}}^{[n]}_{0}$ and
$U_{n}\coloneqq{\mathbb{A}}\left(X\setminus f^{-1}({\boldsymbol{\mathfrak{K}}}^{[n]}),\ \partial{\mathfrak{K}}^{[n]}_{0}\right)\bigcup_{\partial{\mathfrak{K}}^{[n]}_{0}}{\mathfrak{K}}^{[n]}_{0}$
is the covering annulus of $X\setminus f^{-1}\left({\boldsymbol{\mathfrak{K}}}^{[n]}\right)$ rel $\partial{\mathfrak{K}}^{[n]}_{0}$ glued with ${\mathfrak{K}}^{[n]}_{0}$. The correspondence $F_{n}$ is called a $\psi$-ql map and it is independent of the choice of $Y_{n}$ in (3.11). A $\psi$-ql renormalization can be naturally iterated. Moreover, $\psi$-ql bounds can be converted into ql bounds, see Lemma 3.7.
In this section, we will introduce $\psi^{\bullet}$-renormalization by replacing ${\boldsymbol{\mathfrak{K}}}^{[n]}$ with the periodic cycle ${\boldsymbol{\mathfrak{B}}}^{[n]}$ of little bushes so that it is also applicable for satellite combinatorics.
Remark 3.4.
The viewpoint that various classes of self-correspondences ${(g,h)\colon A\rightrightarrows B}$ form interesting dynamical systems was popularized by Sullivan and later by A. Epstein in the context of deformation spaces. In the 2000s, it became apparent that self-correspondences give a natural framework to study some of the classical dynamical systems [IS, K, N]. See also [BD, T] for more recent developments.
3.4.2. Definitions
A ${}^{\bullet}$-quadratic-like map (“$\psi^{\bullet}$-ql map”) is a pair of holomorphic maps
(3.12) $F=(f,\iota)\colon{\ \ }(U,{\mathfrak{B}}^{\prime})\rightrightarrows(V,{\mathfrak{B}})$
between two conformal disks $U$, $V$ with the following properties:
  1. (I)
    $f\colon U\to V$ is a double branched covering (we usually normalize it so that its critical point is located at 00 0);
  2. (II)
    $\iota\colon U\to V$ is an immersion;
  3. (III)
    ${\mathfrak{B}}$ and ${\mathfrak{B}}^{\prime}$ are hulls (compact connected full sets) such that
    $\iota^{-1}({\mathfrak{B}})={\mathfrak{B}}^{\prime}\subset f^{-1}({\mathfrak{B}});$
  4. (IV)
    there exist neighborhoods $X^{\prime}\supset{\mathfrak{B}}^{\prime}$ and $X\supset{\mathfrak{B}}$ with the following property: $\iota:X^{\prime}\rightarrow X$ is a conformal isomorphism such that
    (3.13) $f_{X}\coloneqq f\circ\big{(}\iota\mid X^{\prime}\big{)}^{-1}:X\to f(X^{\prime})\eqqcolon Y$
    is a quadratic-like map with connected filled Julia set
    (3.14) ${\mathfrak{K}}_{F}\coloneqq{\mathfrak{K}}\big{(}f_{X}:X\to Y\big{)},$
    and such that ${\mathfrak{B}}\equiv{\mathfrak{B}}_{F}\equiv{\mathfrak{B}}(f_{X})$ is the bush of $f_{X}$.
Since $\iota$ is a conformal isomorphism in a neighborhood of ${\mathfrak{K}}_{F}$, we will below identify
${\mathfrak{K}}_{F}\simeq\big{(}\iota\mid X\big{)}^{-1}({\mathfrak{K}}_{F}){\ \ }{\ \ }\text{ and hence }{\ \ }{\ \ }{\mathfrak{B}}\simeq{\mathfrak{B}}^{\prime}\equiv{\mathfrak{B}}_{F},$
and write
$F\colon(U,{\mathfrak{B}})\rightrightarrows(V,{\mathfrak{B}}){\ \ }{\ \ }\text{ or }{\ \ }{\ \ }F\colon U\rightrightarrows V.$
Let us say that a subset $\Omega\subset{\mathfrak{K}}_{F}$ is called $\iota$-proper if
$\iota^{-1}(\Omega)=\Omega.$
Item (III) implies that ${\mathfrak{B}}_{F}$ is $\iota$-proper. If ${\mathfrak{K}}_{F}$ is $\iota$-proper, then $F$ is $\psi$-ql map [K].
3.4.3. $\psi^{\bullet}$-ql renormalization
Consider a $\psi^{\bullet}$-ql map $F$ from (3.12), and assume that $f_{X}$ (see (3.13)) is $n+1$ DH renormalizable. Let $f_{n,i}=f^{p_{n}}\colon X_{n,i}\to Y_{n,i}$ be a ql prerenormalization of $f_{X}\colon X\to Y$. The associated $\psi^{\bullet}$-ql renormalization is the extension of $f_{n,i}\colon X_{n,i}\to Y_{n,i}$ along all curves in $V\setminus{\boldsymbol{\mathfrak{B}}}_{n}$, compare with §3.4.1; the result is a $\psi^{\bullet}$-ql map $F_{n,i}=(f_{n,i},\iota_{n,i})\colon U_{n,i}\rightrightarrows V_{n,i}$, where
$V_{n,i}\coloneqq{\mathbb{A}}\left(V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]},\ \partial{\mathfrak{B}}^{[n]}_{i}\right)\bigcup_{\partial{\mathfrak{B}}^{[n]}_{i}}{\mathfrak{B}}^{[n]}_{i}$
is the covering annulus (3.6) of $V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ rel $\partial{\mathfrak{B}}^{[n]}_{i}$ glued with ${\mathfrak{B}}^{[n]}_{i}$ and
$U_{n,i}\coloneqq{\mathbb{A}}\left(U\setminus f^{-1}({\boldsymbol{\mathfrak{B}}}^{[n]}),\ \partial{\mathfrak{B}}^{[n](1)}_{i}\right)\bigcup_{\partial{\mathfrak{B}}^{[n](1)}_{i}}{\mathfrak{B}}^{[n](1)}_{i}$
is the covering annulus of $U\setminus f^{-1}({\boldsymbol{\mathfrak{B}}}^{[n]})$ rel $\partial{\mathfrak{B}}^{[n](1)}_{i}$ glued with ${\mathfrak{K}}^{[n]}_{i}$.
We will write suppress index “00 0” for $F_{n,0}$:
$F_{n,0}\equiv F_{n}=(f_{n},\iota_{n})\colon U_{n}\rightrightarrows V_{n},\hskip 11.38092pt\text{ where }{\ \ }U_{n}\equiv U_{n,0},\ V_{n}\equiv V_{n,0}.$
We write
(3.15) ${\mathcal{R}}^{n\bullet}(F)\coloneqq F_{n}\equiv F_{n,0}.$
Remark 3.5.
The theory works similarly if the bush ${\mathfrak{B}}$ is replaced with any connected forward invariant set $\Upsilon$ satisfying ${\mathfrak{T}}\subset\Upsilon\subset{\mathfrak{K}}$, where ${\mathfrak{T}}$ is the Hubbard continuum. Namely, let us say that a map
(3.16) $F=(f,\iota)\colon{\ \ }(U,\Upsilon^{\prime})\rightrightarrows(V,\Upsilon)$
is $\psi^{\Upsilon}$-ql if it satisfies (I)(IV) from §3.4.2 so that ${\mathfrak{B}}^{\prime}\simeq{\mathfrak{B}}={\mathfrak{B}}(f_{X})$ are replaced with
[…]
For example, one can take $\Upsilon={\mathfrak{T}}$ or $\Upsilon={\mathfrak{T}}\cup{\boldsymbol{\mathfrak{B}}}^{[m]}$. To define “$\psi^{\Upsilon}$-renormalization”, one should extend a ql prerenormalization $f_{n,i}\colon{X_{n,i}\to Y_{n,i}}$ along all curves in $V\setminus\bigcup_{j}\Upsilon_{f_{n,j}}$. Moreover, $\Upsilon_{f_{n,j}}$ can often be enlarged, see §3.4.5.
$\gamma_{1}$$\gamma_{2}$
Figure 10. Illustration to Sup-Chain Rule (3.17), where $n_{1}=n_{2}=1$. Level one bushes ${\boldsymbol{\mathfrak{B}}}^{[1]}$ are depicted black and dashed black (the central bush ${\mathfrak{B}}^{[1]}_{0}$), while level two bushes ${\boldsymbol{\mathfrak{B}}}^{[2]}$ are depicted red and purple (the central bush ${\mathfrak{B}}^{[2]}_{0}$). The map ${\mathcal{R}}^{\bullet}\circ{\mathcal{R}}^{\bullet}(F)$ is obtained from $F$ by extending the ql prerenormalization $f_{2,0}$ (defined near ${\mathfrak{B}}^{[2]}_{0}$, purple) along some paths $\gamma_{i}$ in $V\setminus{\boldsymbol{\Upsilon}}$ (where ${\boldsymbol{\Upsilon}}$ consists of non-dashed sets). The curve $\gamma_{1}$ terminates at ${\boldsymbol{\Upsilon}}$ (i.e., no extension beyond). The curve $\gamma_{2}$ terminates at ${\boldsymbol{\mathfrak{B}}}^{[1]}_{0}$ because it makes a non-trivial loop around ${\boldsymbol{\mathfrak{B}}}^{[1]}\setminus{\mathfrak{B}}^{[1]}_{0}$. Since ${\mathcal{R}}^{2\bullet}F$ is obtained by extending $f_{2,0}$ along all paths in $V\setminus{\boldsymbol{\mathfrak{B}}}^{[2]}$ (where ${\boldsymbol{\mathfrak{B}}}^{[2]}$ is red and purple), we obtain the Sup-Chain Rule ${\mathcal{R}}^{\bullet}\circ{\mathcal{R}}^{\bullet}(F)\subset{\mathcal{R}}^{2\bullet}F$.
3.4.4. Sup-Chain Rule for renormalization domains
Consider a $\psi^{\bullet}$-ql map $F$ as in (3.12). Since
  1. (1)
    ${\mathcal{R}}^{n_{1}\bullet}\circ{\mathcal{R}}^{n_{2}\bullet}(F)$, see (3.15), is the extension of $f_{n_{1}+n_{2},0}$ along curves in $V\setminus{\boldsymbol{\Upsilon}}$ (subject to certain restrictions illustrated on Figure 10), where
    ${\boldsymbol{\Upsilon}}\coloneqq{\boldsymbol{\mathfrak{B}}}^{[n_{1}+n_{2}]}\cup\left({\boldsymbol{\mathfrak{B}}}^{[n_{2}]}\setminus{\mathfrak{B}}^{[n_{2}]}_{0}\right),$
  2. (2)
    ${\mathcal{R}}^{n_{1}+n_{2}\bullet}(F)$ is the extension of $f_{n_{1}+n_{2},0}$ along all curves in $V\setminus{\boldsymbol{\mathfrak{B}}}^{[n_{1}+n_{2}]},$
  3. (3)
    and every connected component of ${\boldsymbol{\mathfrak{B}}}^{[n_{1}+n_{2}]}$ is within a connected component of ${\boldsymbol{\Upsilon}}$ (compare with Figure 10),
we obtain the natural embedding of dynamical systems (respecting all the maps)
(3.17) ${\mathcal{R}}^{n_{1}\bullet}\circ{\mathcal{R}}^{n_{2}\bullet}(F)\ \subset\ {\mathcal{R}}^{n_{1}+n_{2}\bullet}(F).$
3.4.5. Unbalanced renormalization from $\psi^{\bullet}$ to $\psi$-ql maps
Consider a $\psi^{\bullet}$-ql map $F$ from (3.12), and assume that $f_{X}$ (see (3.13)) is twice DH renormalizable. Let $f_{1,0}\colon X_{1,0}\to Y_{1,0}$ be a ql prerenormalization of $f_{X}$.
By Item (III) of the definition of $F$, the little Julia set ${\mathfrak{K}}^{[1]}_{0}\equiv{\mathfrak{K}}(f_{1,0})$ is $\iota$-proper. Therefore, we can extend $f_{1,0}\colon X_{1,0}\to Y_{1,0}$ along all curves in ${\boldsymbol{\Upsilon}}\coloneqq{\mathfrak{K}}^{[1]}_{0}\cup{\boldsymbol{\mathfrak{B}}}^{[1]}$ and obtain $\psi$-ql map
(3.18) $F_{{\boldsymbol{\Upsilon}},1}=(f_{1,0},\iota_{1,0})\colon U_{{\boldsymbol{\Upsilon}},1}\to V_{{\boldsymbol{\Upsilon}},1}.$
We recall from §2.4.5 that the set ${\boldsymbol{\Upsilon}}$ is inhomogeneous: $f_{X}$ does not permute components of ${\boldsymbol{\Upsilon}}$
3.4.6. Restrictions
A $\psi^{\bullet}$-ql map $F$ as in (3.12) has the natural restriction (also known as the pullback, fiber product, graph) denoted by
$F=(f,\iota)\colon U^{2}\rightrightarrows U^{1}=U,{\ \ }{\ \ }\text{ where }{\ \ }U^{2}=\{(x,y)\in U\times U\mid{\ \ }f(x)=\iota(y)\},$
where $f,\iota\colon U^{2}\to U$ are component-wise projections; see [K, §2.2.2]. Note that $F\colon U^{2}\rightrightarrows U^{1}$ is also a $\psi^{\bullet}$-ql map. Repeating the construction, we obtain the sequence
$F\colon U^{k}\rightrightarrows U^{k-1},\hskip 28.45274ptn\geq 1,{\ \ }U^{0}=V,{\ \ }U^{1}=U,$
together with induced iterations denoted by
(3.19) $F^{k}=\big{(}f^{k},\iota^{k}\big{)}\colon U^{k}\to V.$
Since $\iota\mid U^{k+1}$ is the lift of $\iota\mid U$ under the covering $f^{k}$, we have:
Lemma 3.6.
The set $f^{-k}({\mathfrak{B}}_{F})$ is $\iota$-proper for $F\colon U^{k+1}\rightrightarrows U^{k}$.∎
3.4.7. Sup-Chain rule for iterations
Consider a ql map $g\colon A\to B$. Assume that a $\psi$-ql map $F\colon U\rightrightarrows V$ is obtained from $g\colon A\to B$ using finitely many $\psi^{\bullet}$-ql renormalizations followed by the unbalanced renormalization from §3.4.5. Then $F$ has a ql restriction $f_{X}\colon X\to Y$ as in (3.13) that is realized in the dynamical plane of $g$:
$f_{X}\simeq f_{X,g},{\ \ }{\ \ }\text{ where }{\ \ }f_{X,g}=g^{p}\colon X_{g}\to Y_{g}.$
Consider the filled Julia set ${\mathfrak{K}}^{[n]}_{i}\equiv{\mathfrak{K}}(f_{X,g})\subset B$ of $f_{X,g}=g^{p}\colon X_{g}\to Y_{g}$. Let $G\colon N\rightrightarrows M$ be the $\psi$-ql map obtained from $g\colon A\to B$ by extending $f_{X,g}=g^{p}\colon X_{g}\to Y_{g}$ along all paths in ${\boldsymbol{\Upsilon}}\coloneqq{\mathfrak{K}}^{[n]}_{i}\cup{\boldsymbol{\mathfrak{B}}}^{[n]}$; compare with §3.4.5 where the case $(n,i)=(1,0)$ is defined. By construction, we have a covering map
$\rho\colon M\setminus{\mathfrak{K}}_{G}\to B\setminus{\boldsymbol{\Upsilon}}$
so that $\rho$ induces an isomorphism $\rho\colon\partial{\mathfrak{K}}_{G}\overset{\simeq}{\longrightarrow}\partial{\mathfrak{K}}^{[n]}_{i}.$ The Sup-Chain Rule §3.4.4 implies by induction that $F\subset G$. Therefore, we have the induced partial covering:
[…]
so that $\rho^{\prime}\colon\partial{\mathfrak{K}}_{F}\overset{\simeq}{\longrightarrow}\partial{\mathfrak{K}}^{[n]}_{i}.$ Let $\gamma$ be the core hyperbolic geodesic of the annulus $V\setminus{\mathfrak{K}}_{F}$. Then $\gamma_{F}$ can be described as the set points $x$ such that the Brownian motion starting at $x$ has equal probabilities of hitting each of the boundary components of the annulus $V\setminus{\mathfrak{K}}_{F}$. Therefore, the image of the subdisk $V_{\gamma}\Subset V$ bounded by $\gamma$ under $V\setminus{\mathfrak{K}}_{F}\hookrightarrow M\setminus{\mathfrak{K}}_{G}$ is within the subdisk $M_{\beta}\Subset M$ bounded by the core hyperbolic geodesic $\beta$ of the annulus $M\setminus{\mathfrak{K}}_{G}$. Since $\rho$ is a covering with $\rho\colon\partial{\mathfrak{K}}_{G}\overset{\simeq}{\longrightarrow}\partial{\mathfrak{K}}^{[n]}_{i},$ we have $\rho\colon\beta\overset{1:1}{\longrightarrow}\rho(\beta)$ and $\rho(\beta)$ is a hyperbolic geodesic of $B\setminus{\boldsymbol{\Upsilon}}$. Therefore, $V\gamma\setminus{\mathfrak{K}}_{F}$ descents univalently to the dynamical plane of $g\colon A\to B$ via the partial covering (3.20).
3.5. Width and WADs of $\psi^{\bullet}$ maps
For a $\psi^{\bullet}$-ql map $F\colon U\rightrightarrows V$ we write
${\mathcal{W}}_{\bullet}(F)\coloneqq{\mathcal{W}}(V\setminus{\mathfrak{B}}_{F}).$
If $F$ is $\psi$-ql map (i.e., ${\mathfrak{K}}_{F}$ is $\iota$-proper), then we can also measure
${\mathcal{W}}(F)\coloneqq{\mathcal{W}}(V\setminus{\mathfrak{K}}_{F}).$
The Grötzsch inequality easily implies
${\mathcal{W}}\left(U^{k}\setminus{\mathfrak{B}}\right)\leq{\mathcal{W}}\left(U^{k}\setminus f^{-k}({\mathfrak{B}})\right)=2^{k}{\mathcal{W}}_{\bullet}(F).$
The Sup-Chain Rule (3.17) implies that
(3.21) ${\mathcal{W}}\left[{\mathcal{R}}^{n_{1}\bullet}\circ{\mathcal{R}}^{n_{2}\bullet}(F)\right]\geq{\mathcal{W}}\left[{\mathcal{R}}^{n_{1}+n_{2}\bullet}(F)\right].$
3.5.1. Compactness of $\psi$ and $\psi^{\bullet}$-ql maps
If ${\mathcal{W}}(F)$ is bounded for $\psi$-ql map $F$, then $F$ also has regular ql bounds:
Lemma 3.7 ([K, Lemma 2.4]).
There is a positive function $\mu(K)$ with the following property. If $F$ is a $\psi$-ql map with ${\mathcal{W}}(F)\leq K$, then the quadratic-like map $f_{X}\colon X\to Y$ in (3.13) can be selected so that
(3.22) $\operatorname{mod}\left(Y\setminus X\right)\geq\mu(K).$
Remark 3.8.
Assume that $\psi$-ql map $F\colon U\rightrightarrows V$ is obtained from a ql map $g\colon A\to B$ using finitely many $\psi^{\bullet}$-renormalizations followed by the unbalanced renormalization from §3.4.5. As in §3.4.7, let $\gamma$ be the core hyperbolic geodesic of the annulus $V\setminus{\mathfrak{K}}_{F}$. In Lemma 3.7 we can assume (by replacing $X,Y$ with $f_{X}^{-1}(X),X$) that the ql restriction $f_{X}\colon X\to Y$ is within the subdisk $V_{\gamma}\Subset V$ bounded by $\gamma$. It is shown in §3.4.7 that $V_{\gamma}\setminus{\mathfrak{K}}_{F}$ (and hence $V_{\gamma}$) descents univalently to the dynamical plane of $g$ via the partial covering (3.20). Therefore, the ql map $f_{X}\colon X\to Y$ descents univalently to the dynamical plane of $g$.
It is shown in the proof of [K, Lemma 2.4] that $\iota$ is an embedding in a neighborhood of ${\mathfrak{K}}_{F}$; this argument is applicable to $\psi^{\bullet}$ maps:
Lemma 3.9.
There is a positive function $\mu_{\bullet}(K)$ with the following property. If $F$ is a $\psi_{\bullet}$-ql map with ${\mathcal{W}}_{\bullet}(F)\leq K$, then ${\mathfrak{B}}_{F}$ has a neighborhood $\Omega$ such that $\iota$ is injective on $\Omega$ and
$\operatorname{mod}\left(\Omega\setminus{\mathfrak{B}}_{F}\right)\geq\mu_{\bullet}(K).$
3.5.2. WADs of $F$
Assume that ${\boldsymbol{\mathfrak{B}}}^{[n]}$ is well-defined, i.e., $F$ is $n+1$ renormalizable. Let us consider the canonical WAD of $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],m}$:
${\mathcal{A}}^{(m),k}\equiv{\mathcal{A}}^{[n],(m),k}\coloneqq{\operatorname{WAD}}\left(U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],m}\right).$
We say (compare with §1.3.1) that an arc $\alpha\in{\mathcal{A}}^{[n],(m),k}$ is
  • vertical if it connects the outer boundary $\partial^{\mathrm{out}}\left(U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}\right)\coloneqq\partial U^{k}$ of $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ to one of its inner boundary components;
  • horizontal if it connects two inner (i.e., non-outer) boundary components of $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$.
We set (see §3.3.6)
  • ${\mathcal{A}}^{(m),k}_{\mathrm{ver}}$ to be vertical part of ${\mathcal{A}}^{(m),k}$: it consists of arcs connecting $\partial U^{k}$ and one of the components of $\partial{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$;
  • ${\mathcal{A}}^{(m),k}_{\mathrm{hor}}$ to be horizontal part of ${\mathcal{A}}^{(m),k}$: it consists of arcs connecting components of $\partial{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$;
  • ${\mathcal{A}}_{i}^{(m),k}$ to be the local WAD for ${\mathfrak{B}}^{[n],(m),k}_{i}$, see §3.3.5: ${\mathcal{A}}_{i}^{k,(m)}$ consisting of arcs adjacent to ${\mathfrak{B}}^{[n],(m),k}_{i}$ such that the weights of self-arcs adjacent to ${\mathfrak{B}}^{[n],(m),k}_{i}$ are doubled;
  • ${\mathcal{A}}_{i,{\mathrm{ver}}}^{(m),k},{\mathcal{A}}_{i,{\mathrm{hor}}}^{(m),k}$ to be the local parts of ${\mathcal{A}}^{(m),k}_{\mathrm{ver}},\ {\mathcal{A}}^{(m),k}_{\mathrm{hor}}$.
We will usually suppress upper zero induces:
${\mathcal{A}}_{i}={\mathcal{A}}^{(0),0}_{i},\hskip 11.38092pt{\mathcal{A}}_{i}^{(m)}={\mathcal{A}}^{(m),0}_{i},\hskip 11.38092pt{\mathcal{A}}_{i}^{k}={\mathcal{A}}^{(0),k}_{i},\dots$
The following lemma provides a reverse to (3.21) estimate.
Lemma 3.10.
Assume that in the dynamical plane of $F={\mathcal{R}}^{n_{0}\bullet}\colon U\rightrightarrows V$, there is a non-trivial horizontal lamination of curves ${\mathcal{L}}\subset V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ emerging from ${\mathfrak{B}}^{[n]}_{0}$ and landing at ${\boldsymbol{\mathfrak{B}}}^{[n]}$. Then
${\mathcal{W}}\left({\mathcal{R}}^{n_{0}+n\bullet}G\right)\geq{\mathcal{W}}({\mathcal{L}}).$
Proof.
By definition §3.4.2, the ${\mathcal{L}}$ lifts univalently to a horizontal lamination $\widetilde{\mathcal{L}}$ in the dynamical plane of $G$ emerging from ${\mathfrak{B}}^{n_{0}+n}_{G,0}$ and landing at ${\boldsymbol{\mathfrak{B}}}^{[n_{0}+n]}_{G}$. After that $\widetilde{\mathcal{L}}$ lifts univalently to a vertical lamination in the dynamical plane of ${\mathcal{R}}^{n_{0}+n\bullet}G$. We obtain that ${\mathcal{W}}\left({\mathcal{R}}^{n_{0}+n\bullet}G\right)\geq{\mathcal{W}}(\widetilde{\mathcal{L}})\geq{\mathcal{W}}({\mathcal{L}})$. ∎
3.5.3. WADs and ${\mathcal{W}}_{\bullet}(F_{n,i})$
It follows from Theorem 3.1 that
(3.23) ${\mathcal{W}}_{\bullet}(F_{n,i})={\mathcal{W}}({\mathcal{A}}_{i})+O_{p_{n}}(1).$
As with $\psi$-ql renormalization (compare with [McM1, Theorem 9.3]), we have
(3.24) $\frac{1}{2}{\mathcal{W}}(F_{n,0})\leq{\mathcal{W}}(F_{n,i})\leq{\mathcal{W}}(F_{n,0})\hskip 19.91684pt\text{ for all }\ i.$
Indeed, let $\gamma_{i}\subset V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ be the peripheral hyperbolic geodesic around ${\mathfrak{K}}^{[n]}_{i}$. Then the hyperoblic length of $|\gamma_{i}|_{V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}}$ is proportional to ${\mathcal{W}}(F_{n,i})$. Let $\gamma^{1}_{i}$ be the lift of $\gamma_{i}$ under $f\colon U^{(1)}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}\to V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ so that $\gamma^{m}_{i}\subset U^{(m)}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ is the peripheral hyperbolic geodesic around ${\mathfrak{B}}^{[n],(1)}_{i-1}$, where the subscript is $\operatorname{mod}p_{n}$. Counting the degree of $f\colon\gamma^{1}_{i}\to\gamma_{i}$, we obtain
  • $|\gamma^{1}_{i}|_{U^{(1)}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}}=|\gamma_{i}|_{V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}}$ if $i>1$; and
  • $\frac{1}{2}|\gamma^{1}_{1}|_{U^{(1)}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}}=|\gamma_{1}|_{V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}}$.
Finally, $|\gamma^{1}_{i}|_{U^{(1)}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}}\geq|\gamma_{i-1}|_{V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}}$; this implies (3.24).
3.5.4. Covering: from $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{(m)}$ to $U^{k+s}\setminus{\boldsymbol{\mathfrak{B}}}^{(m+s)}$
It follows from §3.3.8 that the WADs change naturally under the covering:
(3.25) ${\mathcal{A}}^{(m+s),k+s}_{\mathrm{hor}}=(f^{s})^{*}{\mathcal{A}}^{(m),k}_{\mathrm{hor}},\hskip 19.91684pt{\mathcal{A}}^{(m+s),k+s}_{i-s}=(f^{s})^{*}{\mathcal{A}}^{(m),k}_{i},$
and similar for other WAD such as ${\mathcal{A}}^{(m),k},{\mathcal{A}}^{(m),k}_{\mathrm{ver}},\dots$.
3.5.5. Restriction by $\iota^{s}$: from $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{(m)}$ to $U^{k+s}\setminus{\boldsymbol{\mathfrak{B}}}^{(m)}$
Assume that $m\leq k$. By Lemma 3.6, the set ${\boldsymbol{\mathfrak{B}}}^{[n],(m)}\subset{\mathfrak{B}}_{F}$ is $\iota$-proper. Let $\gamma$ be a proper curve in $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ emerging from ${\mathfrak{B}}^{[n],(m)}_{a}$. Applying $\iota^{-s}$ along $\gamma$, we construct, see (3.26) below, its restriction $(\iota^{*})^{s}[\gamma]$ which is a proper curve in $U^{k+s}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$. Moreover,
  • if $\gamma$ is vertical, then so is $(\iota^{*})^{s}[\gamma]$,
  • if $\gamma$ is horizontal, then $(\iota^{*})^{s}[\gamma]$ is either horizontal or vertical.
If $(\iota^{*})^{s}[\gamma]$ is horizontal, then we will also call $(\iota^{*})^{s}[\gamma]$ the lift of $\gamma$ under $\iota^{s}$.
More precisely, write $\gamma\colon(0,1)\to U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$. For every $s>0$, there is a $t_{s}\in(0,1]$ such that $\iota^{-s}$ extends along $\gamma\mid(0,t_{s})$ and the resulting curve
(3.26) $(\iota^{*})^{s}[\gamma]\coloneqq\iota^{-s}\circ\gamma\colon(0,t_{s})\to U^{k+s}$
is a proper curve in $U^{k+s}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ (because ${\boldsymbol{\mathfrak{B}}}^{[n],(m)}\subset U^{k}$ is $\iota$-proper). Moreover,
  • if $t_{m}=1$, then $(\iota^{*})^{n}\gamma$ is vertical if and only if $\gamma$ is vertical,
  • if $t_{m}<1$, then $(\iota^{*})^{n}\gamma$ is vertical.
The following lemma implies that curves in a horizontal rectangle that restrict under $(\iota^{s})^{*}$ to vertical curves form buffers of the rectangle.
Lemma 3.11.
Assume that two disjoint horizontal paths $\gamma_{1},\gamma_{2}\subset U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{(m)}$ representing the same arc $\alpha=[\gamma_{1}]=[\gamma_{2}]$ lift under $(\iota^{s})^{*}$ to horizontal paths in $U^{k+s}\setminus{\boldsymbol{\mathfrak{B}}}^{(m)}$. Let $R\subset U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{(m)}$ be the proper rectangle between $\gamma_{1}$ and $\gamma_{2}$ such that all vertical curves in $R$ also represent $\alpha$. Then all curves in ${\mathcal{F}}^{\operatorname{full}}(R)$ lift under $(\iota^{s})^{*}$ to homotopic horizontal curves in $U^{k+s}\setminus{\boldsymbol{\mathfrak{B}}}^{(m)}$.
Proof.
The curve $\gamma\in{\mathcal{F}}^{\operatorname{full}}(R)$ becomes vertical under the restriction by $(\iota^{s})^{*}$ if and only if $\gamma$ hits $\iota^{s}\big{(}\partial U^{k+s}\big{)}$. This is impossible because $\gamma_{1},\gamma_{2}$ do not hit $\iota^{s}\big{(}\partial U^{k+s}\big{)}$. ∎
3.5.6. Monotonicity of the ${\mathcal{A}}^{k}_{\mathrm{hor}}$
Since $\iota\colon U^{k+1}\to U^{k}$ is an embedding in a neighborhood of ${\mathfrak{K}}_{F}$, we identify up to homotopy $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ and $U^{s}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ for all $k,s$. In particular, horizontal arcs $\alpha$ in $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ are naturally viewed as horizontal arcs in $U^{s}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ by realizing the $\alpha$ in a small neighborhood of ${\mathfrak{K}}_{F}$ where $\iota^{s-k}$ is an embedding.
It follows essentially from Lemma 3.3 (by lifting $\iota$ to the universal coverings, see [K, §3.5] for reference) that
(3.27) ${\mathcal{A}}^{[n],(m),k}_{\mathrm{hor}}\geq{\mathcal{A}}^{[n],(m),k+1}_{\mathrm{hor}}\hskip 14.22636pt\text{ and }\hskip 14.22636pt{\mathcal{A}}^{[n],(m),k}_{i,{\mathrm{hor}}}\geq{\mathcal{A}}_{i+1,{\mathrm{hor}}}^{[n],(m),k+1}.$
3.5.7. Domination: from $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{(m)}$ to $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{(m+1)}$ (following [K, §3.6])
Assume that $\gamma\subset U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ is a proper horizontal curve. Then it has the decomposition
(3.28) $\gamma=\ell_{0}\#\gamma_{1}\#\ell_{1}\#\gamma_{2}\#\dots\#\gamma_{s}\#\ell_{s}$
(compare with §2.4.3) such that
  • $\gamma_{j}\subset U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m+1)}$ are proper horizontal curves; and
  • every component of $\ell_{i}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m+1)}$ is trivial in $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m+1)}$ (with respect to a proper homotopy, see §2.4.1).
We call $s$ the expansivity number of $\gamma$ in rel ${\boldsymbol{\mathfrak{B}}}^{(m+1)}$.
It follows from Decomposition (3.28) that there is a constant $C\equiv C_{p_{n},m}$ such that for all $M>0$
[…]
in the following sense, see §3.3.6 for notation. For every arc $\alpha$ in ${\mathcal{A}}^{[n],(m+1),k}_{\mathrm{hor}}-{\operatorname{max}}\{M,1\}C$, there is a vertical sublamination
${\mathcal{L}}_{\alpha}\subset{\mathcal{F}}(e)\equiv{\boldsymbol{\Pi}}^{\operatorname{can}}(\alpha){\ \ }{\ \ }\text{ with }{\ \ }{\ \ }{\mathcal{W}}({\mathcal{L}}_{\alpha})\geq{\mathcal{W}}(\alpha)-{\operatorname{max}}\{M,1\}C$
such that Decomposition 3.28 of every $\gamma\in{\mathcal{L}}_{\alpha}$ has the additional property that
(3.29) $\gamma_{i}\in{\mathcal{F}}^{\operatorname{full}}\left({\mathcal{A}}^{[n],(m+1),k}_{\mathrm{hor}}-C\right).$
The lamination ${\mathcal{L}}_{a}$ is constructed by removing all $\gamma$ from ${\mathcal{F}}(\alpha)$ that do not satisfy (3.29) – the width of removed curves is bounded by ${\operatorname{max}}\{M,1\}C$.
3.5.8. Almost periodic rectangles
Consider a $\psi^{\bullet}$-ql map $F\colon U\rightrightarrows V$ as in (3.12). We say that a proper rectangle $R$ in $V\setminus{\boldsymbol{\mathfrak{B}}}^{[1]}$ is almost periodic if most of the width of $R$ overflows its iterative lift; more precisely:
Definition 3.12.
For a $\psi^{\bullet}$-ql map $F\colon U\rightrightarrows V$, a proper rectangle $R\subset V\setminus{\boldsymbol{\mathfrak{B}}}^{[1]}$ connecting ${\mathfrak{B}}^{[1]}_{a},{\mathfrak{B}}^{[1]}_{a+1}$ is called $\delta$-almost periodic if $R$ represents a genuine periodic arc (of some AD, see §2.4.3) and $R$ has a vertical sublamination ${\mathcal{G}}\subset{\mathcal{F}}(R)$ with ${\mathcal{W}}({\mathcal{G}})\geq(1-\delta){\mathcal{W}}(R)$ such that Conditions (I), (II) stated below hold for all $s\leq 10\overline{p}$, where $\overline{p}$ is the combinatorial bound §2.2.5.
  1. (I)
    Under the immersion, ${\mathcal{G}}$ lifts to the lamination
    ${\mathcal{G}}^{s}\coloneqq(\iota^{sp_{1}})^{*}({\mathcal{G}})\subset U^{sp_{1}}\setminus{\boldsymbol{\mathfrak{B}}}^{[1]}$
    still connecting ${\mathfrak{B}}^{[1]}_{a},{\mathfrak{B}}^{[1]}_{a+1}$.
Let $R^{(s)}\subset U^{(sp_{1})}\setminus{\boldsymbol{\mathfrak{B}}}^{[1],(sp_{1})}$ be the periodic lift of $R$ under $f^{sp_{1}}$; i.e., $R^{(s)}$ connects ${\mathfrak{B}}^{[1],(sp_{1})}_{a},{\mathfrak{B}}^{[1],(sp_{1})}_{a+1}$.
  1. (II)
    The lamination ${\mathcal{G}}^{s}$ overflows $R^{(s)}$ as follows: every curve $\gamma$ in ${\mathcal{G}}^{s}$ is the concatenation $\gamma=\ell_{a}\#\gamma^{\prime}\#\ell_{a+1}$ such that
    • $\gamma^{\prime}\in{\mathcal{F}}^{\operatorname{full}}\big{(}R^{(s)}\big{)}$, see §3.2; and
    • every component of $\ell_{a}\setminus{\mathfrak{B}}^{[1],(sp_{1})}_{a}$ and every component of $\ell_{a+1}\setminus{\mathfrak{B}}^{[1],(sp_{1})}_{a+1}$ is trivial in $U^{sp_{1}}\setminus{\boldsymbol{\mathfrak{B}}}^{[1],(sp_{1})}$ with respect to a proper homotopy, see §2.4.1.
By Lemma 2.3, the first renormalization of $F$ is satellite and ${\mathfrak{B}}^{[1]}_{a},{\mathfrak{B}}^{[1]}_{a+1}$ are neighboring bushes with respect to the cyclic order.
Remark 3.13.
An almost periodic rectangle $R$ between little bushes ${\mathfrak{B}}^{[n]}_{a},{\mathfrak{B}}^{[n]}_{a+1}$ is defined in the same way as in the case $n=1$. Such a rectangle $R$ can be lifted to the dynamical plane of $F_{n-1,c}$ (via the covering map representing the $\psi^{\bullet}$-renormalization, see §3.4.3) and its lift will be an almost periodic rectangle between level 11 11 little bushes as in Definition 3.12 .
4. Pull-off for non-periodic rectangles
In this section we will establish the following theorem that refines a result from [K]:
Theorem 4.1.
For every bound $\bar{p}$ on renormalization periods as in §2.2.5, every small $\delta>0$, and every $n\gg_{\bar{p},\delta}1$, the following holds.
Let $F=(f,\iota)\colon U\rightrightarrows V$ be a $\psi^{\bullet}$-ql map $n+1$ times renormalizable of type $\bar{p}$, and let $F_{n,i}=(f_{n,i},\iota_{n,i})\colon U_{n,i}\rightrightarrows V_{n,i}$ be its $n$th $\psi^{\bullet}$-renormalization, see §3.4.3. If
${\mathcal{W}}_{\bullet}(F_{n,i})=K\gg_{\bar{p},\delta,n}1,$
then
  1. (P)
    either ${\mathcal{W}}_{\bullet}(F)\geq 2K$;
  1. (S)
    or the $n$th renormalization of $F$ is satellite and the $(n-1)$-th $\psi^{\bullet}$-ql renormalization $F_{n-1}$ has a $\delta$-almost periodic rectangle $R$ with ${\mathcal{W}}(R)\geq K/20$ (see Definition 3.12) between two neighboring bushes in its satellite flower.
If the $n$th renormalization of $F$ is primitive, then $F_{n-1}$ has no periodic rectangles; i.e., Case (P) holds.
Proof.
(See also §1.3.1.) Consider the periodic cycle ${\boldsymbol{\mathfrak{B}}}^{[n]}$ of level $n$ little bushes in the dynamical plane of $F\colon U\rightrightarrows V$. Recall from (3.19) that $F^{k}\colon U^{k}\rightrightarrows V$ represents the $k$th iteration of $F$.
The degeneration of $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ is described by the WAD ${\mathcal{A}}^{k}\equiv{\mathcal{A}}^{[n],k}$; let us consider its horizontal and vertical parts ${\mathcal{A}}^{k}_{\mathrm{hor}},{\mathcal{A}}^{k}_{\mathrm{ver}}$, see details in §4.1. Since the $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ decrease, so are the arc diagrams representing ${\mathcal{A}}^{k}_{\mathrm{hor}}$; i.e., the $A^{k}_{\mathrm{hor}}\equiv{\operatorname{AD}}\left({\mathcal{A}}^{k}_{\mathrm{hor}}\right)$ stabilize with $k\leq 3p_{n}$. And since $f^{-1}(U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n]})\subset U^{k+1}\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$, the $A^{k}_{\mathrm{hor}}$ stabilize at an invariant AD; i.e., $A^{k}_{\mathrm{hor}}$ is aligned with the Hubbard tree of the superattracting model.
Consider the local part ${\mathcal{A}}^{k}_{i}$ of ${\mathcal{A}}^{k}$ around ${\mathfrak{B}}^{n}_{i}$. Write $q\coloneqq(10\overline{p}+3)p_{n}$. If for all $i$ we have
(4.1) ${\mathcal{W}}({\mathcal{A}}^{q}_{i})-{\mathcal{W}}({\mathcal{A}}^{0}_{i})\leq\delta_{1}K,{\ \ }{\ \ }{\ \ }\delta_{1}\ll\delta,$
then most of the degeneration in ${\mathcal{A}}^{0}_{i}$ is represented by genuine periodic arcs §2.4.3; this leads to Case (S), see §4.3.
If (4.1) does not hold some $i$, then ${\mathcal{W}}({\mathcal{A}}^{q}_{i,{\mathrm{ver}}})\geq\delta_{1}K$ and applying the Covering Lemma we obtain
${\mathcal{W}}\left(V\setminus{\mathfrak{K}}_{j}^{[n]}\right)\asymp_{\delta_{1},\overline{p}}\ K\hskip 14.22636pt\text{ for all }j.$
Applying the Quasi-Additivity Law and using $p_{n}\gg_{\delta_{1},\overline{p}}1$, we obtain
${\mathcal{W}}(F)\geq\sum_{j}{\mathcal{W}}\left(V\setminus{\mathfrak{K}}_{j}^{[n]}\right)\succeq_{\delta_{1},\overline{p}}\ p_{n}K\gg 2K,$
see §4.2; this is Case (P).
Below we will make a more technical exposition. ∎
4.1. Alignment of WAD with the Hubbard continuum
Following §2.3.1, let
${\boldsymbol{\mathfrak{B}}}\equiv{\boldsymbol{\mathfrak{B}}}^{[n],(0)}{\ \ }{\ \ }\text{ and }{\ \ }{\ \ }{\boldsymbol{\mathfrak{B}}}^{(m)}\equiv{\boldsymbol{\mathfrak{B}}}^{[n],(m)}=f_{X}^{-m}({\boldsymbol{\mathfrak{B}}})$
be the periodic cycle of level-$n$ bushes and its preimage, where $f_{X}=[f\colon X\to Y]$ is a ql restriction of $f$, see (3.13) in §3.4.2. The similar suppression of indices is applied to WAD from §3.5.2.
Recall from §3.5.4 and §3.5.6 that we have
(4.2) ${\mathcal{A}}^{(m+1),k+1}_{{\mathrm{hor}}}=f^{*}\left({\mathcal{A}}^{(m),k}_{{\mathrm{hor}}}\right){\ \ }{\ \ }\text{ and }{\ \ }{\ \ }{\mathcal{A}}^{(m),k+1}_{\mathrm{hor}}\geq{\mathcal{A}}^{(m),k}_{\mathrm{hor}}.$
Let us choose a sufficiently big threshold $C\gg_{p_{n}}1$. By (4.2), the ADs forming ${\mathcal{A}}^{k}_{\mathrm{hor}}-C^{k}$ decrease, hence for a certain
(4.3) ${\mathbf{t}}\leq\ \big{[}\text{maximal number of horizontal arcs on $\displaystyle V\setminus{\boldsymbol{\mathfrak{B}}}_{n}$}\big{]}\ \leq 3p_{n}$
the arc diagram
(4.4) $H={\operatorname{AD}}({\mathcal{H}}),\hskip 14.22636pt\text{ where }\hskip 14.22636pt{\mathcal{H}}\coloneqq{\mathcal{A}}^{\mathbf{t}}_{\mathrm{hor}}-C^{\mathbf{t}}$
coincides with the AD of ${\mathcal{A}}^{{\mathbf{t}}+1}_{\mathrm{hor}}-C^{{\mathbf{t}}+1}$.
Since the WAD ${\mathcal{A}}^{{\mathbf{t}}+1}_{\mathrm{hor}}-C^{{\mathbf{t}}+1}$ on $U^{{\mathbf{t}}+1}\setminus{\boldsymbol{\mathfrak{B}}}$ is dominated by the WAD $f^{*}\left({\mathcal{H}}\right)=f^{*}\left({\mathcal{H}}-{\mathcal{A}}^{\mathbf{t}}_{\mathrm{hor}}-C^{\mathbf{t}}\right)$ on $U^{{\mathbf{t}}+1}\setminus{\boldsymbol{\mathfrak{B}}}^{(1)}$, see §3.5.7, we obtain that $H$ is an invariant AD, and therefore it is aligned with the Hubbard continuum by Lemma 2.2.
4.2. Case (P)
Consider local WAD ${\mathcal{A}}_{i}^{k}$ as in §3.5.2, we have by §3.5.3
(4.5) ${\mathcal{W}}({\mathcal{A}}_{i}^{k})\geq{\mathcal{W}}({\mathcal{A}}_{i}^{0})={\mathcal{W}}_{\bullet}(F_{n,i})-O_{p_{n}}(1)\geq K/2-O_{p_{n}}(1).$
Write $q\coloneqq(10\overline{p}+3)p_{n}$ and fix a sufficiently small $\delta_{1}>0$. Assume that:
[…]
    there is a $\kappa$ such that
    ${\mathcal{W}}({\mathcal{A}}_{\kappa,{\mathrm{ver}}}^{q})\geq\delta_{1}K.$
Since all local weights are comparable §3.5.3, we have the Collar Assumption:
(4.6) ${\mathcal{W}}\left(V\setminus\bigcup_{j\not=i}{\mathfrak{B}}_{j},{\ \ }{\mathfrak{B}}_{i}\right)\asymp K\hskip 22.76228pt\text{ for all $\displaystyle i$.}$
Since ${\mathcal{W}}({\mathcal{A}}_{\kappa,{\mathrm{ver}}}^{q+j})\geq{\mathcal{W}}({\mathcal{A}}_{\kappa,{\mathrm{ver}}}^{q})\geq\delta_{1}K$ and §4.6 holds, we obtain from the [KL1, Covering Lemma] that
${\mathcal{W}}(V,{\mathfrak{B}}_{\kappa+i})\succeq_{\delta_{1},\overline{p}}\ {\mathcal{W}}(U^{q+i},{\mathfrak{B}}_{\kappa})\geq{\mathcal{W}}({\mathcal{A}}_{\kappa,{\mathrm{ver}}}^{q+i})-O_{p_{n}}(1)\succeq_{\delta_{1},\overline{p}}\ K$
for all $i\in\{0,1,\dots,p_{n}-1\}$. Therefore,
$\sum_{j=0}^{p_{n}}{\mathcal{W}}(V,{\mathfrak{B}}_{j})\succeq_{\delta_{1},\overline{p}}\ p_{n}K.$
[…]
${\mathcal{W}}(V\setminus{\boldsymbol{\mathfrak{B}}})\succeq_{\delta_{1},\overline{p}}\ \sum_{j=0}^{p_{n}}{\mathcal{W}}(V,{\mathfrak{B}}_{j})\succeq_{\delta_{1},\overline{p}}\ p_{n}K\geq 2K.$
We conclude that
${\mathcal{W}}^{\bullet}(F)={\mathcal{W}}(V,\ {\mathfrak{B}}_{F})\geq{\mathcal{W}}(V,\ {\boldsymbol{\mathfrak{B}}})\geq 2K.$
4.3. Case (S)
[…]
Let ${\mathcal{H}}$ be the WAD from (4.4) on $(U^{\mathbf{t}},{\boldsymbol{\mathfrak{B}}})$. Consider an arc $e$ in $H$ with maximal ${\mathcal{W}}(e)$. Since the number of arcs in $H$ is bounded by $3p_{n}$ and ${\mathcal{W}}({\mathcal{H}})\geq p_{n}K/4-\delta_{1}p_{n}K$ by (3.24), we obtain that ${\mathcal{W}}(e)\geq K/14$. We claim that $e$ is within a periodic satellite flower ${\mathfrak{B}}^{[n-1]}_{c}$. In particular, the $n$-th renormalization of $f$ is satellite.
Assume converse: $e$ is not in any satellite flower ${\mathfrak{B}}^{[n-1]}_{c}$. By Lemma 2.4, $e$ overflows arcs $e_{1},e_{2}\in(f^{2p_{n}})^{*}(H)$ with $f^{2p_{n}}(e_{1})=f^{2p_{n}}(e_{2})=e_{\operatorname{new}}\in H$. By Lemma 3.11, $\big{(}\iota^{2p_{n}}\big{)}^{*}\left[{\mathcal{F}}(e-\delta_{1}K)\right]$ is a horizontal family of curves in $U^{{\mathbf{t}}+2p_{n}}\setminus{\boldsymbol{\mathfrak{B}}}$; after removing $O_{p_{n}}(1)$ curves, this family is dominated by $(f^{2p_{n}})^{*}(H)$. Therefore, by the Series Law:
${\operatorname{max}}\left({\mathcal{W}}(e_{1}),\ {\mathcal{W}}(e_{2})\right)\geq 2{\mathcal{W}}(e)-2.1\ \delta_{1}K.$
Applying $f^{2p_{1}}$, we obtain the contradiction: ${\mathcal{W}}(e_{\operatorname{new}})>{\mathcal{W}}(e)$.
Consider now a satellite flower ${\mathfrak{B}}^{[n-1]}_{c}$. As in §2.4.4, we denote by ${\mathcal{H}}_{c}$ the WAD consisting of arcs of ${\mathcal{H}}$ that are in ${\mathfrak{B}}^{[n-1]}_{c}$.
Write $q^{\prime}=10\overline{p}p_{n}$. Let ${\mathcal{H}}^{(q^{\prime})}_{c}$ be the WAD consisting of arcs of $\big{(}f^{q^{\prime}}\big{)}^{*}{\mathcal{H}}$ that are in ${\mathfrak{B}}^{[n-1]}_{c}$. By Lemma 3.11, the restriction $(\iota^{q^{\prime}})^{*}\left[{\mathcal{H}}_{c}-\delta_{1}K\right]$ consists of horizontal curves. Since $H={\operatorname{AD}}({\mathcal{H}})$ is aligned with the Hubbard dendrite, $(\iota^{q^{\prime}})^{*}\left[{\mathcal{H}}_{c}-\delta_{1}K\right]$ is, after removing $O_{p_{n}}(1)$ curves, dominated by ${\mathcal{H}}^{(q^{\prime})}_{c}$, see §3.5.7. And since $(f^{q^{\prime}})_{*}\colon{\mathcal{H}}^{q^{\prime}}_{c}\to{\mathcal{H}}_{c}$ is a bijection, as most $\overline{p}\delta_{1}K+2C^{2}<\delta_{2}K$ curves in ${\mathcal{H}}_{c}$ can have expansivity number greater than 22 22, where $\delta_{1}\ll\delta_{2}\ll\delta$. This implies that most curves in ${\mathcal{F}}({\mathcal{H}}_{c})$ are within rectangles representing genuine periodic arcs §2.4.3 that connect neighboring level $n$ bushes of ${\mathfrak{B}}^{[n-1]}_{c}$, see Lemma 2.3. This also demonstrates that all the ${\mathcal{H}}_{c}$ have comparable width – the difference of their weights are bounded by $\delta_{2}K$.
Since one of the ${\mathcal{H}}$ has an edge $e$ with ${\mathcal{W}}(e)\geq K/14$, the map $F\colon U^{{\mathbf{t}}+1}\rightrightarrows U^{\mathbf{t}}$ has a $\delta/2$-almost periodic rectangle $R$ between neighboring level $n$ bushes of ${\mathfrak{B}}^{[n-1]}_{0}$ with ${\mathcal{W}}(R)\geq K/15$ and satisfying Remark 3.13.
4.3.1. Pushing forward $R$ into $V\setminus{\boldsymbol{\mathfrak{B}}}^{n}$ and then into the dynamical plane of $F_{n-1}$
Since ${\mathcal{W}}({\mathcal{A}}_{i,{\mathrm{ver}}}^{q})\leq\delta_{1}K$, curves in ${\mathcal{F}}\big{(}{\mathcal{A}}_{i,{\mathrm{hor}}}-2\delta_{1}K\big{)}$ restrict under $(\iota^{\mathbf{t}})^{*}$ to horizontal curves in $(U^{\mathbf{t}},{\boldsymbol{\mathfrak{B}}})$. Since ${\mathcal{A}}^{{\mathbf{t}}}_{i}\leq{\mathcal{A}}_{i}$, see (3.27), we obtain from Lemma 3.11 that
${\mathcal{F}}\big{(}{\mathcal{A}}^{\mathbf{t}}_{i,{\mathrm{hor}}}-2\delta_{1}K-2\big{)}\subset\ (\iota^{\mathbf{t}})^{*}\left[{\mathcal{F}}^{\operatorname{full}}\big{(}{\mathcal{A}}_{i,{\mathrm{hor}}}-2\delta_{1}K\big{)}\right].$
Therefore, ${\mathcal{F}}({\mathcal{A}}^{\mathbf{t}}_{i,{\mathrm{hor}}}-2\delta_{1}K-2)$ can be univalently pushed forward under $(\iota^{\mathbf{t}})_{*}$. By removing $\delta_{1}K$ buffers from $R$, we push forward $R$ into the dynamical plane of $F\colon U\rightrightarrows V$ and then push forward $R$ into the dynamical plane of $F_{n-1}$, see Remark 3.13. This establishes Case (S) of the theorem.
5. Waves
Given a compact connected filled set $X\subset{\mathbb{C}}$, we denote by $\partial^{c}X$ its Carathéodory boundary; i.e., the set of prime ends of ^$\widehat{\mathbb{C}}\setminus X$. A Riemann map identifies $\partial^{c}\widetilde{X}$ with the unit circle ${\mathbb{T}}$. For a compact connected subset $Y\subset X$, we denote $\partial^{c}_{X}Y\subset\partial^{c}X$ the set of prime ends of $X$ accumulating at $Y$. A side of $Y$ rel $X$ is a connected component of $\partial^{c}_{X}Y$ viewed as a subset of ${\mathbb{T}}\simeq\partial^{c}X$.
Let $F=(f,\iota)\colon U\rightrightarrows V$ be a $\psi^{\bullet}$-ql map. Assume that
  • $({\mathfrak{K}}_{k})_{k}$ is a forward invariant collection of periodic and preperiodic little filled Julia sets of $f$ of the same level;
  • $T\supset\bigcup_{k}{\mathfrak{K}}_{k}$ is forward invariant compact connected filled subset of the Julia set of $f$;
  • $T$ is $\iota$-proper;
  • every ${\mathfrak{K}}_{k}$ has finitely many sides in $T$.
A relevant example for us is $T={\mathfrak{B}}_{F}^{(m)}$. Let us denote by $M$ the total number of sides of all ${\mathfrak{K}}_{k}$.
Consider a side ${\mathfrak{K}}_{k}^{\iota}$ of ${\mathfrak{K}}_{k}$ in $T$. A wave ${\mathcal{S}}$ above ${\mathfrak{K}}_{k}^{\iota}$ is a lamination of proper paths in $U\setminus T$ such that every curve $\gamma\in{\mathcal{S}}$ starts and ends at $\partial^{c}X\setminus{\mathfrak{K}}_{k}^{\iota}$ and goes above ${\mathfrak{K}}_{k}^{\iota}$: the bounded component $O$ of $U\setminus(\gamma\cup T)$ contains ${\mathfrak{K}}_{k}^{\iota}$ on its Carathéodory boundary (i.e., prime ends of ${\mathfrak{K}}_{k}^{\iota}$ are accessible from $O$).
Lemma 5.1 (Wave Lemma).
Let ${\mathcal{S}}$ be a wave as above. Then ${\mathcal{W}}(U\setminus T)\succeq_{M}{\mathcal{W}}({\mathcal{S}}).$
Proof.
We will use the following fact:
Lemma 5.2.
Suppose $f\colon A\to B$ is a degree $m$ covering between closed annuli. Suppose $J\subset\partial A$ is an interval. Let ${\mathcal{S}}$ be a wave in $A$ above $J$. Then ${\mathcal{S}}$ contains a genuine subwave ${\mathcal{S}}^{\operatorname{new}}$ such that $f\mid{\mathcal{S}}^{\operatorname{new}}$ is injective and
${\mathcal{W}}({\mathcal{S}}^{\operatorname{new}})\geq{\mathcal{W}}({\mathcal{S}})-O(\ln m).$
In particular, if $f\mid J$ is not injective, then ${\mathcal{W}}({\mathcal{S}})=O(\ln m)$.
Proof.
Consider universal covering maps $X\to A$ and $X\to B$. Their group of deck transformations are isomorphic to $m{\mathbb{Z}}$ and ${\mathbb{Z}}$ respectively. Let ${\mathcal{S}}_{k}$, $k\in m{\mathbb{Z}}$ be the lifts of ${\mathcal{S}}$ under $X\to A$; all these lifts are disjoint and permuted by $m{\mathbb{Z}}$. Let ${\mathcal{S}}_{k}$, $k\in{\mathbb{Z}}$ be the orbit of ${\mathcal{S}}$ under ${\mathbb{Z}}$. We claim that, by removing $O(\ln m)$ outermost curves from ${\mathcal{S}}$, we obtain ${\mathcal{S}}^{\operatorname{new}}$ and the new ${\mathcal{S}}^{\operatorname{new}}_{k},k\in{\mathbb{Z}}$ so that
${\mathcal{S}}^{\operatorname{new}}_{0},{\ \ }{\mathcal{S}}^{\operatorname{new}}_{1},{\ \ }\dots,{\mathcal{S}}^{\operatorname{new}}_{m-1}$
are pairwise disjoint. Then, by the claim, all ${\mathcal{S}}^{\operatorname{new}}_{k},k\in{\mathbb{Z}}$ are pairwise disjoint, i.e.
$f\colon{\mathcal{S}}^{\operatorname{new}}\longrightarrow f({\mathcal{S}}^{\operatorname{new}})={\mathcal{S}}^{\operatorname{new}}_{0}/{\mathbb{Z}}$
is injective.
Let us verify the claim. Let us denote by $A_{1}$ the component of $\partial A$ containing $J$, by $H\simeq{\mathbb{Z}}/m$ the group of deck transformations of $f\colon A\to B$. Then $H$ acts on $A_{1}$. We can decompose ${\mathcal{S}}={\mathcal{S}}^{0}\sqcup{\mathcal{S}}^{1}\sqcup\dots\mathcal{S}^{T}$ into possibly empty pairwise-disjoint laminations such that
  • curves in ${\mathcal{S}}^{0}$ start and end at an interval $I^{0}\subset A_{1}$ that is a fundamental interval for the action $H\curvearrowright A_{1}$;
  • curves in ${\mathcal{S}}^{t}$, $t>0$ start at an interval $I^{t}_{-}\subset A_{1}$ and end at an interval $I^{t}_{+}\subset A_{1}$ such that one of the intervals $I^{t}_{-},I^{t}_{+}$ is within a union of $1\displaystyle 2^{t}-$ 1 fundamental intervals for the action $H\curvearrowright A_{1}$ and the interval $J^{t}\subset A_{1},$ $J^{t}\cap J\not=\emptyset$ between $I^{t}_{-},I^{t}_{+}$ is the union of exactly $1\displaystyle 2^{t}-$ 1 fundamental intervals of the action $H\curvearrowright A_{1}$;
  • $T=O(\ln m)$.
Then ${\mathcal{W}}({\mathcal{S}}^{t})=O(1)$ for $t>0$ because ${\mathcal{S}}^{t}$ crosses its shift under $H$, see [DL2, Shift Argument §A.3].
Replacing ${\mathcal{S}}$ with ${\mathcal{S}}^{0}$, we obtain that curves in the ${\mathcal{S}}^{0}_{k}$ start and end at pairwise disjoint intervals. Removing an extra $0(1)$ outermost curves, we obtain that the new ${\mathcal{S}}^{\operatorname{new}}_{l}$ are pairwise disjoint. ∎
Let us assume first that ${\mathfrak{K}}_{k}$ is periodic. Fix an iteration $f^{n}\mid T,\ n\leq M^{3}$ such that $f^{n}\mid{\mathfrak{K}}_{k}^{\iota}$ covers $\partial^{c}_{T}{\mathfrak{K}}_{k}$ at least twice. Consider the associated iteration
[…]
Under the immersion $\iota^{n}$, either $\frac{1}{3}{\mathcal{W}}({\mathcal{S}})$ part of ${\mathcal{S}}$ lifts to a vertical family of $U^{n}\setminus T$ or $\frac{2}{3}{\mathcal{W}}({\mathcal{S}})$ part of ${\mathcal{S}}$ lifts univalently into the lamination ${\mathcal{S}}^{\prime}$. In the former case, the lemma is proven:
(5.1) ${\mathcal{W}}(V\setminus T)\geq\frac{1}{2^{n}}{\mathcal{W}}(U^{n}\setminus T)\geq\frac{1}{2^{n}3}{\mathcal{W}}({\mathcal{S}}).$
Assume the latter case. Let $\widetilde{T}$ be the full preimage of $T$ under $f^{n}$. Then ${\mathfrak{K}}_{k}^{\iota}$ splits into finitely many sides $X_{1},X_{2},\dots,X_{t}$ of ${\mathfrak{K}}_{k}$ in $\widetilde{T}$. Every $X_{i}$ maps univalently to the side $f^{n}(X)$ of ${\mathfrak{K}}_{k}$ in $T$ under $f^{n}$. Since $f^{n}\mid{\mathfrak{K}}_{k}^{\iota}$ covers $\partial^{c}_{T}{\mathfrak{K}}_{k}$ at least twice, we can choose two sides $X_{a},X_{b}$ that map onto ${\mathfrak{K}}^{\iota}_{k}$. Every curve $\gamma\in{\mathcal{S}}^{\prime}$ has first shortest subcurves $\gamma_{a},\gamma_{b}$ (which may coincide) in $U^{n}\setminus\widetilde{T}$ above $X_{a},X_{b}$ respectively. By Lemma 5.2, the width of $\gamma$ with $\gamma_{a}=\gamma_{b}$ is $O_{M}(1)$.
Let ${\mathcal{F}}^{\prime}$ be the family of all $\gamma$ in ${\mathcal{S}}$ with $\gamma_{a}\not=\gamma_{b}$. Let ${\mathcal{S}}_{a},{\mathcal{S}}_{b}$ be family of curves consisting of $\gamma_{a}$ and $\gamma_{b}$ with $\gamma$ in ${\mathcal{S}}^{\prime}$. Since ${\mathcal{S}}^{\prime}$ overflows ${\mathcal{S}}_{a}$ and then ${\mathcal{S}}_{b}$, either ${\mathcal{W}}({\mathcal{S}}_{a})$ or ${\mathcal{W}}({\mathcal{S}}_{b})$ is at least $4/3K-O_{M}(1)$. After removing $O_{M}(1)$ buffers and applying Lemma 5.2, the waves ${\mathcal{S}}_{a},{\mathcal{S}}_{b}$ map univalently by onto waves above ${\mathfrak{K}}_{k}^{\iota}$. We obtain a wave with width $\geq\frac{4}{3}K-O_{M}(1)$ and the whole argument can be repeated with a strictly bigger wave leading eventually to the contradiction.
Assume now that ${\mathfrak{K}}_{k}$ is strictly preperiodic. Let $f^{n}$ be the smallest iteration so that $f^{n}({\mathfrak{K}}_{k})$ is periodic. As above, either the $1/3$ part of ${\mathcal{S}}$ lifts to a vertical family under $\iota^{n}$ or the $2/3$ part of ${\mathcal{S}}$ lifts univalently. In the former case, (5.1) concludes the proof. In the latter case, we construct the family ${\mathcal{S}}_{a}$ as above and then we pushforward ${\mathcal{S}}_{a}$ under $f^{n}$. The result will be a wave above a side of a periodic Julia set. This reduces the preperiodic to the periodic case. ∎
6. Pull-off for periodic rectangles
Consider a $\psi$-ql like map $F=(f,\iota)\colon U\rightrightarrows V$. We assume that the first renormalization of $F$ is satellite.
Theorem 6.1.
Fix a combinatorial bounds $\overline{p}$ on the renormalization period §2.2.5. Then for every sufficiently small $\delta>0$ there is a $C_{\delta}=C_{\delta,\overline{p}}>1$ with
$C_{\delta}\to\infty{\ \ }{\ \ }\text{ as }{\ \ }{\ \ }\delta\to 0$
such that the following holds for every $\psi^{\bullet}$-ql map $F$, and its $\psi^{\bullet}$-ql renormalizations $\ F_{1}={\mathcal{R}}^{\bullet}(F),\ F_{2}={\mathcal{R}}^{2\bullet}(F)$, see (3.15).
Suppose that $R$ with ${\mathcal{W}}(R)\gg_{\delta,\overline{p}}1$ is a $\delta$-almost periodic rectangle (see Definition 3.12) in the dynamical plane of $F$ between bushes ${\mathfrak{B}}_{a}^{[1]}$ and ${\mathfrak{B}}_{a+1}^{[1]}$. Then
(6.1) $\text{either }{\ \ }{\ \ }{\ \ }{\mathcal{W}}_{\bullet}(F_{2})\geq C_{\delta}K{\ \ }{\ \ }\text{ or }{\ \ }{\ \ }{\mathcal{W}}_{\bullet}(F)\geq C_{\delta}K.$
Moreover, if $F={\mathcal{R}}^{n\bullet}(G)$, then we also have the alternative
(6.2) $\text{either }{\ \ }{\ \ }{\ \ }{\mathcal{W}}_{\bullet}\left[{\mathcal{R}}^{n+2\bullet}(G)\right]\geq C_{\delta}K{\ \ }{\ \ }\text{ or }{\ \ }{\ \ }{\mathcal{W}}_{\bullet}(F)\geq C_{\delta}K$
(independent of $n$).
6.1. Proof of Theorem 6.1
Let $\Pi\subset{\mathfrak{B}}_{F}$ be the geodesic continuum between ${\mathfrak{B}}^{[n]}_{a}$ and ${\mathfrak{B}}^{[1]}_{a+1}$. We will use notations from Definition 3.12 such as ${\mathcal{G}},{\mathcal{G}}^{s},R^{(s)}$.
6.1.1. Spiraling Numbers
First we will introduce the spiraling parameters rel $\Pi$ for a curve $\gamma\in R$ shortly summarized in the following remark:
${\mathfrak{B}}^{[1]}_{a}$${\mathfrak{B}}^{[1],(q)}_{a}$${\mathfrak{B}}^{\prime\prime}_{1}$$\ell_{1}$${\mathfrak{B}}^{\prime}_{1}$${\mathfrak{B}}_{i}^{[2]}$${\mathfrak{B}}_{i+1}^{[2]}$${\mathfrak{B}}^{\prime\prime}_{2}$${\mathfrak{B}}^{\prime}_{2}$$\ell_{2}$${\mathfrak{B}}^{[1]}_{a+1}$${\mathfrak{B}}^{[1],(q)}_{a+1}$${\mathcal{G}}_{a}$$\alpha$$R^{(q)}$
Figure 11. The non-spiraling case. The rectangle $R$ (orange, violet, red) connects ${\mathfrak{B}}^{[1]}_{a}$ and ${\mathfrak{B}}^{[1]}_{a+1}$. Between ${\mathfrak{B}}^{[1],(q)}_{a}$ and ${\mathfrak{B}}^{[1],(q)}_{a+1}$ most of the rectangle $R$ travels through its lift $R^{(q)}$ (red). Therefore, the part ${\mathcal{G}}_{a}$ (violet) of $R$ between $\ell_{1}$ and $\ell_{2}$ is tremendously wide. The bushes ${\mathfrak{B}}^{\prime}_{1},{\mathfrak{B}}^{\prime\prime}_{1},{\mathfrak{B}}^{\prime}_{2},{\mathfrak{B}}^{\prime\prime}_{2}$ are lifts of ${\mathfrak{B}}^{[1]}_{a+1}$, see Figure 6.
Remark 6.2.
Recall the the pure Mapping Class Group $\operatorname{MCG}$ of a 33 33-punctured sphere $S^{2}\setminus\{a,b,\infty\}$ is trivial. By replacing $a,b$ with small open Jordan disks $D_{a},D_{b}$, we obtain the new group $\operatorname{MCG}(S^{2}\setminus(D_{a}\cup D_{b}\cup\{\infty\}))\simeq{\mathbb{Z}}^{2}$ consisting of Dehn-twists around $D_{a},D_{b}$, see [FM]. This fact has the following implication.
Let $O$ be a small disk-neighborhood of ${\mathfrak{B}}^{[1]}_{a}\cup\Pi\cup{\mathfrak{B}}^{[1]}_{a+1}$. Then $O^{\prime}\coloneqq O\setminus({\mathfrak{B}}^{[1]}_{a}\cup{\mathfrak{B}}^{[1]}_{a+1})$ is topologically a disk with two holes. Therefore, a path (simple curve) $\beta\subset O^{\prime}$ from ${\mathfrak{B}}^{[1]}_{a}$ to ${\mathfrak{B}}^{[1]}_{a+1}$ has a simple description: it first spirals $\tilde{t}_{a}\in{\mathbb{Z}}$ times around ${\mathfrak{B}}^{[1]}_{a}$ and then spirals $\tilde{t}_{a+1}\in{\mathbb{Z}}$ times around ${\mathfrak{B}}^{[1]}_{a+1}$, where $\Pi$ can be used as a reference “path” of zero spiralings. Below we will ignore the spiraling orientation and introduce the absolute quantity $t_{a}=|\tilde{t}_{a}|$.
Recall §3.3 that $\gamma$ lands at the ideal boundary of $V\setminus{\boldsymbol{\mathfrak{B}}}^{[1]}$. Let $\gamma_{\Pi}\subset V\setminus{\boldsymbol{\mathfrak{B}}}^{[1]}$ be a path homotopic rel the endpoints to $\gamma$ in $V\setminus{\boldsymbol{\mathfrak{B}}}^{[1]}$ such that $\gamma_{\Pi}\cap\Pi$ is the minimal possible number. In other words, $\gamma_{\Pi}$ and $\Pi$ are in the minimal position – this is well defined because $\Pi$ has infinitely many cut points.
Consider the components $\gamma_{0},\gamma_{1},\dots,\gamma_{f},\gamma_{f+1}$ of $\gamma_{\Pi}\setminus\Pi.$ Since $\gamma$ is properly homotopic into (a small neighborhood of) $\Pi$, there is a $t_{a}\leq f$ such that
  • $\gamma_{t}\cup\Pi$ surrounds ${\mathfrak{B}}^{[1]}_{a}$ for $t\leq t_{a}$;
  • $\gamma_{t}\cup\Pi$ does not surround ${\mathfrak{B}}^{[1]}_{a}$ for $t>t_{a}$.
(It follows that $\gamma_{t}\cup\Pi$ surrounds ${\mathfrak{B}}^{[1]}_{a+1}$ for $t>t_{a}$, and only $\gamma_{t_{a}}\cup\Pi$ can surround both ${\mathfrak{B}}^{[1]}_{a}$ and ${\mathfrak{B}}^{[1]}_{a}$.)
We say that $t_{a}=t_{a}(\gamma)$ is the spiraling number of $\gamma$ around ${\mathfrak{B}}^{[1]}_{a}$. Similarly, the spiraling number of $\gamma\in R^{(s)}$ around ${\mathfrak{B}}^{[1],(sp_{1})}_{a}$ is introduced. Below are some properties of spiraling numbers:
  1. (A)
    The spiraling numbers of $\gamma_{1},\gamma_{2}\in R$ differ by at most 11 11.
  2. (B)
    If $\gamma_{s}\in R^{(s)}$ is the lift of $\gamma\in R$ into $R^{(s)}$, then $t_{a}(\gamma^{s})\leq\frac{1}{2^{s}}t_{a}(\gamma)$.
Indeed, (A) follows from the fact (the disjoint curves) $\gamma_{1},\gamma_{2}$ can be simultaneously put into the minimal position with $\Pi$. And (B) follows from the fact that $f^{sp_{1}}\colon{\mathfrak{B}}^{[1],(sp_{1})}_{a}\to{\mathfrak{B}}^{[1]}_{a}$ has degree $2^{s}$.
6.1.2. The non-spiraling case:
$t_{a}(\gamma)\leq 4$ for all $\gamma\in R$; see Figure 11 for illustration. Write $q\coloneqq 10\overline{p}$ and note that $qp_{1}>10p_{2}$.
Let us consider the objects introduced in Lemma 2.1. It follows from (B) that curves in $R^{(3)}$ do not spiral around ${\mathfrak{B}}^{[1],(3)}_{a}$. Therefore, $\ell_{2}$ separates the base of $R^{(q)}$ from ${\mathfrak{B}}_{a}^{[1]}$: the base $\partial^{h,0}R^{(q)}$ and ${\mathfrak{B}}_{a}^{[1]}$ are in different components of ${\mathfrak{B}}^{[1],(q)}_{a}\setminus\ell_{2}$.
Let ${\mathcal{G}}_{a}$ be the restriction of ${\mathcal{G}}^{s}$ between $\ell_{1}$ and $\ell_{2}$ – this lamination consists of the first shortest subarcs $\gamma^{\prime}\subset\gamma$ between $\ell_{1}$ and $\ell_{2}$ for all $\gamma\in{\mathcal{G}}^{s}$. Since ${\mathcal{G}}^{s}$ consequently overflows ${\mathcal{G}}_{a}$ and then $R^{(q)}$, we obtain from the Grötzsch inequality that
(6.3) $(1-\delta){\mathcal{W}}(R)\leq{\mathcal{W}}({\mathcal{G}}^{q})\leq{\mathcal{W}}(R^{(q)})\oplus{\mathcal{W}}({\mathcal{G}}_{a})={\mathcal{W}}(R)\oplus{\mathcal{W}}({\mathcal{G}}_{a}),$
where ${\mathcal{W}}(R^{(q)})={\mathcal{W}}(R)$ because $R^{(q)}$ is a lift of $R$. Therefore, ${\mathcal{W}}({\mathcal{G}}_{a})\geq C_{0,\delta}{\mathcal{W}}(R)\geq C_{0,\delta}K,$ where $C_{0,\delta}\to\infty$ as $\delta\to 0$.
There are two possibilities. If a substantial part of ${\mathcal{G}}_{a}$ travels through both ${\mathfrak{B}}^{[2]}_{i},{\mathfrak{B}}^{[2]}_{i+1}$, then this substantial part of ${\mathcal{G}}_{a}$ restricts to a wide lamination ${\mathcal{L}}$ between ${\mathfrak{B}}^{[2]}_{i},{\mathfrak{B}}^{[2]}_{i+1}$. Pushing ${\mathcal{L}}$ with respect to $f_{*}$ and $\iota^{*}$ and using Lemma 3.2 (the number of iterations and the degree is bounded in terms of $\overline{p}$), we obtain a family ${\mathcal{F}}$ of non-trivial proper curves in $V\setminus{\boldsymbol{\mathfrak{B}}}^{[2]}$ starting at ${\mathfrak{B}}^{[2]}_{0}$ such that ${\mathcal{W}}({\mathcal{F}})\succeq_{\overline{p}}C_{0,\delta}K$. If a substantial part of ${\mathcal{F}}$ is vertical, then we obtain the second estimate in (6.1), (6.2). If a substantial part of ${\mathcal{F}}$ is horizontal, then we obtain the first estimate in (6.1); applying Lemma 3.10 to ${\mathcal{F}}$, we obtain the first estimate in (6.2).
If a substantial part of ${\mathcal{G}}_{a}$ omits either ${\mathfrak{B}}^{[2]}_{i}$ or ${\mathfrak{B}}^{[2]}_{i+1}$, then we obtain a wide wave above one of the sides of ${\mathfrak{B}}^{[2]}_{i},{\mathfrak{B}}^{[2]}_{i+1}$; Wave Lemma 5.1 implies that ${\mathcal{W}}_{\bullet}(F)\geq C_{\delta}K$.
[…]
Figure 12. The spiraling case. Since $R$ and its lift $R^{(1)}$ have different spiraling numbers around ${\mathfrak{B}}^{[1]}_{a}$ and ${\mathfrak{B}}^{[1]}_{a+1}$ respectively, (most of) $R$ spirals first around ${\mathfrak{B}}^{[1]}_{a}$ before it continues as $R^{(1)}$ (red). The part ${\mathcal{G}}_{a}$ (violet) of $R$ before $R^{(1)}$ is tremendously wide.
6.1.3. The spiraling case:
$t_{a}(\gamma)\geq 4$ for all $\gamma\in R$; see Figure 12 for illustration. It follows from (B) that the spiraling number of any curve in ${\mathcal{G}}^{1}$ around ${\mathfrak{B}}^{[1]}_{a}$ is strictly less than the spiraling number of any curve in $R^{(1)}$ around ${\mathfrak{B}}^{[1],(1)}_{a}$. Therefore, curves in ${\mathcal{G}}^{1}$ first spirals around ${\mathfrak{B}}_{a}$ before they continue within ${\mathcal{F}}^{\operatorname{full}}(R^{(1)})$. Applying the Grötzsch inequality as in (6.3), we obtain a wide lamination ${\mathcal{G}}_{a}$ with ${\mathcal{W}}({\mathcal{G}}_{a})\geq C_{2,\delta}{\mathcal{W}}(R)$ such that ${\mathcal{G}}_{a}$ creates a wave above one of the sides of ${\mathfrak{B}}_{a}$. Wave Lemma 5.1 implies that ${\mathcal{W}}_{\bullet}(f)\geq C_{\delta}K$.∎
Remark 6.3.
The argument of this section will be repeated in §7.0.1 with ${\mathfrak{B}}^{[1]}_{a+1}$ being replaced by a little Julia set ${\mathfrak{K}}^{[1]}_{0}$. Observe that no specific properties of ${\mathfrak{B}}^{[1]}_{a+1}$ has been used in this section (only that it is a forward invariant set that is disjoint from ${\mathfrak{B}}^{[1]}$).
7. Conclusions
In this section, we will now deduce the main theorems.
Theorem 7.1 (A priori beau $\psi^{\bullet}$-bounds).
For any combinatorial bound $\bar{p}$, there is an $n>1$ and $K_{\bar{p}}>1$ such that the following holds. If $F$ is an infinitely renormalizable $\psi^{\bullet}$-ql map of bounded type $\bar{p}$, then
${\mathcal{W}}_{\bullet}\left[({\mathcal{R}}^{n\bullet})^{m}(F)\right]\leq K_{\overline{p}}{\ \ }{\ \ }{\ \ }\text{ for }{\ \ }m\ \gg_{{\mathcal{W}}(F)}\ 1,$
${\mathcal{R}}^{n\bullet}$ is a $\psi^{\bullet}$-ql renormalization (3.15).
Proof.
Let us choose a sufficiently small $\delta>0$ such that $C_{\delta}\gg\Delta_{\overline{p}}$, where $C_{\delta}$ is from Theorem 6.1 and $\Delta_{\overline{p}}$ is from Proposition 2.7. We next choose a sufficiently big $n\gg_{\overline{p},\delta}1$ such that Theorem 4.1 is applicable as follows: if
${\mathcal{W}}_{\bullet}\left[({\mathcal{R}}^{n\bullet})^{m}(F)\right]=K\gg_{\overline{p},\delta,n}1,$
then
  1. (I)
    either ${\mathcal{W}}_{\bullet}\left[({\mathcal{R}}^{n\bullet})^{(m-1)}(F)\right]\geq 2K$;
  2. (II)
    or ${\mathcal{W}}_{\bullet}\left[{\mathcal{R}}^{n+1\bullet}\circ({\mathcal{R}}^{n\bullet})^{m-1}(F)\right]\geq C_{\delta}K$,
where (II) follows from Case (S) of Theorem 4.1 combined with Theorem 6.1.
Write $G\coloneqq({\mathcal{R}}^{n\bullet})^{(m-1)}(F).$ Reapplying Theorem 4.1 for $G$ and its renormalization ${\mathcal{R}}^{n+1\bullet}G$, we obtain the alternative “(I) vs (II’)”, where
  1. (II’)
    or ${\mathcal{W}}_{\bullet}\left[{\mathcal{R}}^{n+2\bullet}(G)\right]\geq C^{2}_{\delta}K$.
Repeating the argument, we eventually obtain the alternative “(I) vs (II”)”, where
  1. (II”)
    or ${\mathcal{W}}_{\bullet}\left[{\mathcal{R}}^{2n\bullet}(G)\right]\geq C^{n}_{\delta}K$.
Therefore, if there are no a priori beau $\psi^{\bullet}$-bounds, then we obtain
${\mathcal{W}}_{\bullet}\left[({\mathcal{R}}^{n\bullet})^{m}(F)\right]\succeq_{F}C^{nm}_{\delta}K.$
This contradicts the Teichmüller contraction: by Proposition 2.7, see also Remark 2.8, the sequence $({\mathcal{R}}^{n\bullet})^{m}(F)$ restricts to ql maps $f_{nm}:X_{nm}\to Y_{nm}$ such that
${\mathcal{W}}\big{(}Y_{nm}\setminus X_{nm}\big{)}=O_{F}(\Delta_{\overline{p}}^{nm}),$
where $\Delta_{\overline{p}}\ll C_{\delta}.$
Theorem 7.2 (A priori beau ql-bounds).
For any combinatorial bound $\bar{p}$, there is a $K_{\bar{p}}>1$ such that the following holds. If $f\colon U\to V$ is an infinitely renormalizable ql map of bounded type $\bar{p}$, then for $n\gg_{{\mathcal{W}}(f)}1$, the map $f_{n}$ has a ql restriction
$f_{n}\colon X_{n}\to Y_{n}{\ \ }{\ \ }{\ \ }\text{ with }{\ \ }{\ \ }{\mathcal{W}}(Y_{n}\setminus X_{n})\leq K_{\overline{p}}.$
Proof.
Write $F\coloneqq f$, and consider the sequence $G_{m}\coloneqq({\mathcal{R}}^{n\bullet})^{m}(F)$ from Theorem 7.1. We have
(7.1) ${\mathcal{W}}_{\bullet}(G_{m})\leq K_{\overline{p}}\hskip 28.45274pt\text{ for}m\gg_{{\mathcal{W}}(F)}1$
for $K_{\overline{p}}$ from Theorem 7.1.
For $m$ satisfying (7.1), consider the dynamical plane of $G_{m}\colon U\to V$. Let ${\boldsymbol{\mathfrak{B}}}^{[1]}$ be the cycle of little bushes in the dynamical plane of $G_{m}$. We  claim that there is a space between ${\mathfrak{K}}^{[1]}_{0}$ and ${\boldsymbol{\mathfrak{B}}}^{[1]}\setminus{\mathfrak{B}}^{[1]}_{0}$; i.e., that there is a $K_{2}>0$ depending on $K_{\overline{p}},\overline{p}$ such that
(7.2) ${\mathcal{W}}\left(V\setminus\bigcup_{i\not=0}{\mathfrak{B}}^{[1]}_{i},\ {\mathfrak{K}}^{[1]}_{0}\right)\leq K_{2}.$
Then (7.2) together with §3.4.5, Lemma 3.7, and Remark 3.8 will imply that $G_{m}$ has a ql renormalization around ${\mathfrak{K}}_{0}^{[1]}$ with a definite modulus.
7.0.1. Proof of (7.2)
Assuming converse, the associated degeneration is arbitrary big:
$K\coloneqq{\mathcal{W}}\left({\mathcal{G}}\right)\gg_{K_{\overline{p}}}1,\hskip 19.91684pt\text{ where }{\ \ }{\mathcal{G}}\coloneqq{\mathcal{F}}\left(V\setminus\bigcup_{i\not=0}{\mathfrak{B}}^{[1]}_{i},\ {\mathfrak{K}}^{[1]}_{0}\right).$
We will now argue that a substantial part of ${\mathcal{G}}$ travels through level two little bushes; this will lead to a contradiction to the $\psi^{\bullet}$-bounds of Theorem 7.2.
Write ${\boldsymbol{\Upsilon}}\coloneqq{\boldsymbol{\mathfrak{B}}}^{[1]}\cup{\mathfrak{K}}^{[1]}_{0}$, and consider the horizontal and vertical WAD ${\mathcal{A}}^{[1],k}_{\mathrm{hor}},{\mathcal{A}}^{[1],k}_{\mathrm{ver}}$ of $U^{k}\setminus{\boldsymbol{\Upsilon}}$. By (7.1),
(7.3) ${\mathcal{W}}({\mathcal{A}}^{[1],k}_{\mathrm{ver}})=O_{k,\overline{p}}(K_{\overline{p}})\hskip 14.22636pt\text{ and hence }\hskip 11.38092pt{\mathcal{W}}({\mathcal{A}}^{[1],k}_{\mathrm{hor}})=K-O_{k,\overline{p}}(K_{\overline{p}}).$
As in the proof of Theorems 4.1, see §4.1, the arc diagrams ${\mathcal{A}}^{[1],kp_{1}}_{\mathrm{hor}}$ eventually stabilize:
${\operatorname{AD}}\left({\mathcal{A}}^{[1],(t+1)p_{1}}_{\mathrm{hor}}-C^{t+1}\right)={\operatorname{AD}}\left({\mathcal{A}}^{[1],tp_{1}}_{\mathrm{hor}}-C^{t}\right){\ \ }{\ \ }{\ \ }\text{ for }{\ \ }t\leq 3p_{1},$
where $C\gg_{p_{1}}1$ is fixed. By Lemma 2.5, $H={\operatorname{AD}}\left({\mathcal{A}}^{[1],(t+1)p_{1}}_{\mathrm{hor}}-C^{t+1}\right)$ is invariant under $f^{p_{1}}$. Since most of the horizontal curves restrict to horizontal curves (by (7.3)) the same argument as in §4.3 provides a $\delta$-almost periodic rectangle $R\subset{V\setminus{\boldsymbol{\Upsilon}}},\ {\mathcal{W}}(R)\asymp K$ between ${\mathfrak{K}}^{[1]}_{0}$ and some ${\mathfrak{B}}^{[1]}_{a}$, where $\delta$ is sufficiently small and ${\mathfrak{K}}^{[1]}_{0}$ replacing ${\mathfrak{B}}^{[1]}_{a+1}$ in Definition 3.12. By Lemma 2.6, the first renormalization of $G_{m}$ is satellite.
We can now repeat the argument of Theorem 6.1, see Remark 6.3, with ${\mathfrak{K}}^{[1]}_{0}$ replacing ${\mathfrak{B}}^{[1]}_{a+1}$ to obtain
$\text{either }{\ \ }{\ \ }{\ \ }{\mathcal{W}}_{\bullet}\left[{\mathcal{R}}^{2\bullet}(G_{m})\right]\succeq_{\overline{p}}K{\ \ }{\ \ }\text{or }{\ \ }{\ \ }{\mathcal{W}}_{\bullet}(G_{m})\succeq_{\overline{p}}K;$
both estimates leading to a contradiction with $\psi^{\bullet}$-bounds of Theorem 7.1. ∎

### v3

@@
 It followed up with the work [KL2, KL3] handling the definitely primitive case.
-It took more than 10 years to prove MLC at some infinitely renormalizable parameters of bounded satellite type [DL1] (based upon the Pacmen Renormalization Theory developed in [DLS]).
+It took more than 10 years to prove MLC at some infinitely renormalizable parameters of bounded satellite type [DL1] (based upon the Pacman Renormalization Theory developed in [DLS]).
 However, the most prominent parameter, the period-doubling Feigenbaum point corresponding to the cascade of doubling renormalizations (see Figure 1) was not covered by this result.
-In this paper we are filling in this gap by proving MLC at the Feigenbaum point, and in fact, in all infinitely renormalizable maps of bounded type (for which we will still preserve the same name).
+In this paper we are filling in this gap by proving MLC at the Feigenbaum point, and in fact, at all infinitely renormalizable parameters of bounded type (for which we will still preserve the same name).
 Remark 1.1.
-The universality of period-doubling cascade was discovered in the mid 70s by Feigenbaum [F1, F2] (the parameter part, see Figure 1) and independently by Coullet and Tresser [TC, CT] (the dynamical part).
+The universality of the period-doubling cascade was discovered in the mid 1970s by Feigenbaum [F1, F2] (the parameter part, see Figure 1) and independently by Coullet and Tresser [TC, CT] (the dynamical part).
 It is intimately related to the renormalization phenomenon in the Quantum Field Theory and Statistical Mechanics; the discovery opened up a new universality paradigm in Dynamical Systems.
@@
 Let us consider the set ${\mathcal{I}}_{\bar{p}}$ of Feigenbaum parameters $c\in{\mathcal{M}}$ of type $\bar{p}$.
-Theorem A implie that each $c\in{\mathcal{I}}_{\bar{p}}$ is the intersection of the nest $M^{n}(c)$ of little $M$-copies corresponding to the $n$-fold renormalizations of $f_{c}$.
+Theorem B implies that each $c\in{\mathcal{I}}_{\bar{p}}$ is the intersection of the nest ${\mathcal{M}}^{n}(c)$ of little ${\mathcal{M}}$-copies corresponding to the $n$-fold renormalizations of $f_{c}$.
 Let us say that Mandelbrot set is self-similar over ${\mathcal{I}}_{\bar{p}}$ if the straightening map $\chi:{\mathcal{M}}^{n+1}(c)\rightarrow{\mathcal{M}}^{n}(\chi(c))$ (see §2.2.3) is $C^{1+{\delta}}$-conformal at any $c\in{\mathcal{I}}_{\bar{p}}$ (with uniform constants).
@@
 If ${\mathcal{W}}_{\mathrm{loc}}^{\mathrm{ver}}({\mathfrak{K}}_{j})\ll{\mathcal{W}}_{\mathrm{loc}}({\mathfrak{K}}_{j})$, then most curves of the horizontal rectangles ${\mathcal{G}}_{1},{\mathcal{G}}_{2}$ overflow their preimages ${\mathcal{G}}^{\prime}_{1},{\mathcal{G}}^{\prime}_{2},{\mathcal{G}}^{\prime\prime}_{1}$ as illustrated.
-Such a configuration is impossible by the Grötzsch inequality; i.e, (1.4) holds.
+Such a configuration is impossible by the Grötzsch inequality; i.e., (1.4) holds.
 Peripheral preimages are omitted.
@@
 and, for every ${\mathfrak{K}}_{j}$:
-(1.2) ${\mathcal{W}}(f_{n})={\mathcal{W}}_{\mathrm{loc}}({\mathfrak{K}}_{j})+O_{p}(1)\coloneqq\sum_{\begin{subarray}{c}{\mathcal{G}}_{i}\text{ adjacent}\\ \text{to }{\mathfrak{K}}_{j}\end{subarray}}{\mathcal{W}}({\mathcal{G}}_{i})+O_{p}(1),$
-i.e. the sum in (1.2) is taken over all rectangles ${\mathcal{G}}_{i}$ adjacent to ${\mathfrak{K}}_{j}$. (A “$\psi$-ql renormalization” is used to achieve (1.2), see §3.43.5.2, and (3.23).) Set also
+(1.2) ${\mathcal{W}}(f_{n})\asymp{\mathcal{W}}_{\mathrm{loc}}({\mathfrak{K}}_{j})\ \coloneqq\sum_{\begin{subarray}{c}{\mathcal{G}}_{i}\text{ adjacent}\\ \text{to }{\mathfrak{K}}_{j}\end{subarray}}{\mathcal{W}}({\mathcal{G}}_{i})+O_{p}(1),$
+i.e. the sum in (1.2) is taken over all rectangles ${\mathcal{G}}_{i}$ adjacent to ${\mathfrak{K}}_{j}$. (A $\psi$-ql renormalization” is used to achieve (1.2), see §3.43.5.2, and (3.23).) Set also
 (1.3) ${\mathcal{W}}_{\mathrm{loc}}^{\mathrm{ver}}({\mathfrak{K}}_{j})\coloneqq\sum_{\begin{subarray}{c}\text{vertical }{\mathcal{G}}_{i}\\ \text{adjacent to }{\mathfrak{K}}_{j}\end{subarray}}{\mathcal{W}}({\mathcal{G}}_{i}).$
-A fundamental fact is that the horizontal ${\mathcal{G}}_{i}$ are eventually (after several restrictions of the domain) aligned with the Hubbard tree $T_{f}$ of $f$ as it shown on Figure 3.
+A fundamental fact is that the horizontal ${\mathcal{G}}_{i}$ are eventually (after several restrictions of the domain) aligned with the Hubbard tree $T_{f}$ of $f$ as shown on Figure 3.
 Applying the Grötzsch inequality to the horizontal rectangles and their pullbacks (see the caption of Figure 3), we experience a definite loss of the horizontal weight in favor of the vertical one.
 The created definite vertical weight can then be pushed forward by means of the Covering Lemma [KL1].
-Since all the local weights are comparable,  [^1] This observation goes back to McMullen [McM1, Theorem 9.3].
-we obtain:
+Since all the local weights are comparable [^1] This observation goes back to McMullen [McM1, Theorem 9.3]., we obtain:
 (1.4) ${\mathcal{W}}_{\mathrm{loc}}^{\mathrm{ver}}({\mathfrak{K}}_{j})\asymp{\mathcal{W}}_{\mathrm{loc}}({\mathfrak{K}}_{j}){\ \ }{\ \ }{\ \ }\text{ for every }{\mathfrak{K}}_{j},$
@@
 We stress that a key combinatorial ingredient used for (1.4) is the absence of periodic horizontal rectangles; i.e.
-a rectangle that has an iterated lift homotopic to itself rel small Julia sets, see §2.4, §3.5.8.
+a rectangle that has an iterated lift homotopic to itself rel small Julia sets, see §1.3.2, §2.4, §3.5.8.
 Combining  (1.1), (1.4), (1.2), and assuming $p\gg 1$, we obtain:
@@
 Either most of the horizontal degeneration is within periodic rectangles ${\mathcal{R}}_{0},{\mathcal{R}}_{1}$ such that a $(1-\delta)$-portion of each ${\mathcal{R}}_{i}$ overflows its periodic lifts ${\mathcal{R}}^{\prime}_{i}$.
-Or we have ${\mathcal{W}}^{\mathrm{ver}}_{\mathrm{loc}}({\mathfrak{K}}_{j})\succeq_{\delta}{\mathcal{W}}_{\mathrm{loc}}({\mathfrak{K}}_{j}).$
+Or we have ${\mathcal{W}}^{\mathrm{ver}}_{\mathrm{loc}}({\mathfrak{B}}_{j})\succeq_{\delta}{\mathcal{W}}_{\mathrm{loc}}({\mathfrak{B}}_{j}).$
 1.3.2. Satellite case
-Let us now assume that $f$ is an infinitely renormalizable map with the periodic-doubling combinatorics; i.e., $f$ is hybrid equivalent to a map in the period-doubling Feigenbaum combinatorial class.
+Let us now assume that $f$ is an infinitely renormalizable map with the period-doubling combinatorics; i.e., $f$ is hybrid equivalent to a map in the period-doubling Feigenbaum combinatorial class.
 The general bounded case is similar.
-Let ${\mathfrak{B}}\coloneqq{\mathfrak{K}}^{1}_{0}\cup{\mathfrak{K}}^{1}_{1}$ be the bouquet consisting of two level 11 11 little Julia sets of $f$. Instead of ${\mathcal{W}}(f)$, we will measure
+Let ${\mathfrak{B}}\coloneqq{\mathfrak{K}}^{1}_{0}\cup{\mathfrak{K}}^{1}_{1}$ be the bouquet consisting of two level $1$ little Julia sets of $f$. Instead of ${\mathcal{W}}(f)$, we will measure
 ${\mathcal{W}}_{\bullet}(f)\coloneqq{\mathcal{W}}(V\setminus{\mathfrak{B}}).$
@@
     or else we have:
-    ${\mathcal{W}}_{\mathrm{loc}}^{\mathrm{ver}}({\mathfrak{K}}_{j})\asymp_{\delta}{\mathcal{W}}_{\mathrm{loc}}({\mathfrak{K}}_{j}){\ \ }{\ \ }{\ \ }\text{ for every }{\mathfrak{K}}_{j}$
+    ${\mathcal{W}}_{\mathrm{loc}}^{\mathrm{ver}}({\mathfrak{B}}_{j})\asymp_{\delta}{\mathcal{W}}_{\mathrm{loc}}({\mathfrak{B}}_{j}){\ \ }{\ \ }{\ \ }\text{ for every }{\mathfrak{B}}_{j}$
 If (II) holds, then assuming $p\gg_{\delta}1$ and ${\mathcal{W}}_{\bullet}(f_{n})\gg_{p,\delta}1$ and repeating the argument of (1.5), we obtain
-  1. (${}^{\prime}$)
+  1. (II)
     ${\mathcal{W}}_{\bullet}(f)>2{\mathcal{W}}_{\bullet}(f_{n})$;
-the Record Argument leads to a contradiction in case (${}^{\prime}$).
-The dichotomy “Case (I) vs Case (${}^{\prime}$)” is stated as Theorem 4.1. This is a refinement of [K].
+the Record Argument leads to a contradiction in case (II).
+The dichotomy “Case (I) vs Case (II)” is stated as Theorem 4.1. This is a refinement of [K].
 Consider now Case (I).
-As Figure 5 illustrates, bouquets ${\mathfrak{B}}_{0},{\mathfrak{B}}_{1}$ grow under lifting thus the “combinatorial distance” between $\widetilde{\mathfrak{B}}_{0},\widetilde{\mathfrak{B}}_{1}$ shrinks.
+As Figure 5 illustrates, bouquets ${\mathfrak{B}}_{0},{\mathfrak{B}}_{1}$ grow under lifting, thus the “combinatorial distance” between $\widetilde{\mathfrak{B}}_{0},\widetilde{\mathfrak{B}}_{1}$ shrinks.
 This allows us to detect laminations ${\mathcal{S}}_{i}$ (the “differences” between ${\mathcal{R}}={\mathcal{R}}_{k}$ and its lift ${\mathcal{R}}^{\prime}$) that are much wider than ${\mathcal{W}}({\mathcal{R}})\succeq{\mathcal{W}}_{\bullet}(f_{n})$; i.e.
@@
 In the latter case, the Wave Lemma implies that ${\mathcal{W}}_{\bullet}(f_{n-1})\gg{\mathcal{W}}_{\bullet}(f_{n})$ and the Record Argument is applicable.
-In the former case, we obtain ${\mathcal{W}}_{\bullet}(f_{n+1})\succeq{\mathcal{W}}({\mathcal{S}}_{i})\succeq C_{\delta}{\mathcal{W}}(f_{n})$; i.e.:
-${\mathcal{W}}_{\bullet}(f_{n+1})\geq C^{\prime}_{\delta}{\mathcal{W}}(f_{n}){\ \ }{\ \ }\text{ with }{\ \ }C^{\prime}_{\delta}\to\infty\text{{\ \ }}{as}{\ \ }\delta\to 0.$
-For a sufficiently small $\delta$, this eventually contradicts the Teichmüller contraction within hyperbolic classes [S1] (stated as Proposition 2.7):
+In the former case, we obtain ${\mathcal{W}}_{\bullet}(f_{n+1})\succeq{\mathcal{W}}({\mathcal{S}}_{i})\succeq C_{\delta}{\mathcal{W}}_{\bullet}(f_{n})$; i.e.:
+${\mathcal{W}}_{\bullet}(f_{n+1})\geq C^{\prime}_{\delta}{\mathcal{W}}_{\bullet}(f_{n}){\ \ }{\ \ }\text{ with }{\ \ }C^{\prime}_{\delta}\to\infty\text{{\ \ }}{as}{\ \ }\delta\to 0.$
+For a sufficiently small $\delta$, this eventually contradicts the Teichmüller contraction within hybrid classes [S1] (stated as Proposition 2.7):
 ${\mathcal{W}}_{\bullet}(f_{n})=O\left(\Delta^{n}\right){\ \ }{\ \ }{\ \ }\text{ for some }\Delta>1.$
@@
 Acknowledgments
-The first author was partially supported by the NSF grant DMS 20555322055532 20555322055532 and the ERC grant “HOLOGRAM”.
+The first author was partially supported by the NSF grant DMS $2055532$ and the ERC grant “HOLOGRAM”.
 The second author has been partly supported by the NSF, the Hagler and Clay Fellowships, the Institute for Theoretical Studies at ETH (Zurich), and MSRI (Berkeley).
 We would like to thank Jeremy Kahn for many stimulating discussions.
+We are grateful to the referee for a comprehensive reading of the manuscript.
+A number of our pictures are made with W.
+Jung’s program Mandel.
 Results of this paper were first announced in 2021 at the Quasiworld Seminar and at the CIRM conference “Advancing Bridges in Complex Dynamics”, see the talk by the second author [L4].
-A number of our pictures are made with W. Jung’s program Mandel.
-${\mathcal{R}}$${\mathfrak{B}}_{0}$${\mathfrak{B}}_{1}$${\mathcal{R}}^{\prime}$${\mathcal{S}}_{1}$${\mathcal{S}}_{2}$$\widetilde{\mathfrak{B}}_{0}$$\widetilde{\mathfrak{B}}_{1}$
+[…]
 Figure 5.
@@
   • ${\mathfrak{B}}_{f}^{(m)}\coloneqq f^{-m}({\mathfrak{B}}_{f})$ the little bush of height $m$,
-  • ${\mathfrak{B}}^{[n]}_{i}\equiv{\mathfrak{B}}^{[n],(0)}\subset{\mathfrak{B}}^{[n],(m)}\subset{\mathfrak{K}}^{[n]}_{i}$ the associated little bushes,
+  • ${\mathfrak{B}}^{[n]}_{i}\equiv{\mathfrak{B}}^{[n],(0)}_{i}\subset{\mathfrak{B}}^{[n],(m)}_{i}\subset{\mathfrak{K}}^{[n]}_{i}$ the associated little bushes,
   • ${\boldsymbol{\mathfrak{B}}}^{[n]}\coloneqq\bigcup_{i}{\mathfrak{B}}^{[n]}_{i}\subset{\boldsymbol{\mathfrak{K}}}^{[n]}$ the union of periodic level $n$ little bushes,
-  • ${\boldsymbol{\mathfrak{B}}}^{[n],(m)}\coloneqq f^{-1}\left({\boldsymbol{\mathfrak{B}}}^{[n]}\right)$ the union of level $n$ and height $m$ little periodic and preperiodic bushes,
+  • ${\boldsymbol{\mathfrak{B}}}^{[n],(m)}\coloneqq f^{-m}\left({\boldsymbol{\mathfrak{B}}}^{[n]}\right)$ the union of level $n$ and height $m$ little periodic and preperiodic bushes,
   • ${\boldsymbol{\mathfrak{B}}}^{[n],(m)}_{\operatorname{per}}\coloneqq{\boldsymbol{\mathfrak{B}}}^{[n],(m)}\cap{\boldsymbol{\mathfrak{K}}}^{[n]}$ the union of periodic level $n$ and height $m$ little bushes.
 If ${\mathfrak{B}}_{f}={\boldsymbol{\mathfrak{K}}}^{[1]}$, i.e., ${\mathfrak{T}}_{f}\subset{\boldsymbol{\mathfrak{K}}}^{[1]}$, then we will also refer to ${\mathfrak{B}}_{f}$ as the satellite flower of $f$.
-Preperiodic little Julia sets and bushes are defined accordingly, 2.3.1.
+Preperiodic little Julia sets and bushes are defined accordingly, §2.3.1.
 We will usually assume that
@@
     ${\mathfrak{K}}^{[n]}_{0}$ contains the critical point of $f$.
+Periodic little Julia sets will be usually labeled dynamically:
+$f\left({\mathfrak{K}}^{[n]}_{i}\right)={\mathfrak{K}}^{[n]}_{i+1},{\ \ }{\ \ }{\ \ }i,\ i+1\in{\mathbb{Z}}/{p_{n}}\simeq\{0,1,\dots,p_{n}-1\}.$
+However, in the satellite combinatorics, we will often label the ${\mathfrak{K}}^{[n]}_{i}$ with respect to their cyclic order in the satellite flower; see, for instance, Lemma 2.3.
+Preperiodic little Julia sets ${\mathfrak{K}}^{[n]}_{j}$ will be labeled with some subindex.
+The labeling of little bushes is naturally inherited from the labeling of little Julia sets: ${\mathfrak{B}}^{[n],(m)}_{i}\subset{\mathfrak{K}}^{[n]}_{i}$.
+We will always indicate when the cyclic order is used.
+And we will also mention when a preperiodic object appears.
 The Teichmüller contraction of ql renormalization will be established in §2.5.
@@
 A quadratic-like map $f:X\rightarrow Y$, which will also be abbreviated as a ql map, is a holomorphic double branched covering between two Jordan disks $X\Subset Y\subset{\mathbb{C}}$.
-It has a single critical point that we usually put at the origin 00 0.
+It has a single critical point that we usually put at the origin $0$.
 The annulus $A=Y\setminus\overline{X}$ is called the fundamental annulus of $f\colon X\to Y$.
@@
 where the supremum is taken over all $f\colon X^{\prime}\to Y^{\prime}$ representing the ql germ of $f$.
-Two quadratic-like maps $f:X\rightarrow Y$ and $\tilde{f}:\tilde{X}\rightarrow\tilde{Y}$ are called hybrid conjugate if they are conjugate by a quasiconformal map $h:(Y,X)\rightarrow(\tilde{Y},\tilde{X})$ such that ¯$\bar{\partial}h=0$ a.e.
+Two quadratic-like maps $f:X\rightarrow Y$ and $\tilde{f}:\tilde{X}\rightarrow\tilde{Y}$ are called hybrid conjugate if they are conjugate by a quasiconformal map $(Y~,X~)\displaystyle h$$\rightarrow(\tilde{Y},\tilde{X})$ such that $\bar{\partial}h=0$ a.e.
 on ${\mathfrak{K}}$.
@@
 2.2.2. Hubbard continua
-In the superattracting case (when the crittical point is periodic) the Hubbard tree ${\mathfrak{T}}_{f}$ is defined as the smallest tree in ${\mathfrak{K}}$ containing the postcritical set such that ${\mathfrak{T}}_{f}$ intersects the components of $\operatorname{int}{\mathfrak{K}}$ along internal rays.
-It is invariant under $f$ and is marked with the orbit of 00 0.
+In the superattracting case (when the critical point is periodic) the Hubbard tree ${\mathfrak{T}}_{f}$ is defined as the smallest tree in ${\mathfrak{K}}$ containing the postcritical set such that ${\mathfrak{T}}_{f}$ intersects the components of $\operatorname{int}{\mathfrak{K}}$ along internal rays.
+It is invariant under $f$ and is marked with the orbit of $0$.
 If ${\mathfrak{K}}$ is periodically repelling and locally connected, we can define the Hubbard dendrite ${\mathfrak{T}}_{f}$ as the smallest connected closed forward-invariant subset of ${\mathfrak{K}}$ containing the postcritical set.
@@
 For a periodically repelling $f$ and $x,y\in{\mathfrak{J}}_{f}={\mathfrak{K}}_{f}$, we define a geodesic continuum
-(2.2) $[x,y]\coloneqq\pi^{-1}[\pi(x),\pi(y)],$
+(2.2) $x,y]\coloneqq\pi^{-1}[\pi(x),\pi(y)],$
 where $[\pi(x),\pi(y)]$ is a unique arc in ${\mathfrak{J}}^{\mathrm{com}}$ connecting $\pi(x),\pi(y)$.
@@
 We say that a Feigenbaum map is primitive/satellite if all its renormalizations are such.
-A Feigenabaum map has a priori bounds if
+A Feigenbaum map has a priori bounds if
 (2.7) $\operatorname{mod}{\mathcal{R}}^{n}f\geq{\epsilon}>0$
@@
 We write
-${\boldsymbol{\mathfrak{B}}}^{[n]}\coloneqq\bigcup_{i}{\mathfrak{B}}_{i}^{[n]},{\ \ }{\boldsymbol{\mathfrak{B}}}^{[n],(m)}\coloneqq f^{-1}\left({\boldsymbol{\mathfrak{B}}}^{[n]}\right),{\ \ }{\boldsymbol{\mathfrak{B}}}^{[n],(m)}_{\operatorname{per}}\coloneqq{\boldsymbol{\mathfrak{B}}}^{[n],(m)}\cap{\boldsymbol{\mathfrak{K}}}^{[n]}.$
+${\boldsymbol{\mathfrak{B}}}^{[n]}\coloneqq\bigcup_{i}{\mathfrak{B}}_{i}^{[n]},{\ \ }{\boldsymbol{\mathfrak{B}}}^{[n],(m)}\coloneqq f^{-m}\left({\boldsymbol{\mathfrak{B}}}^{[n]}\right),{\ \ }{\boldsymbol{\mathfrak{B}}}^{[n],(m)}_{\operatorname{per}}\coloneqq{\boldsymbol{\mathfrak{B}}}^{[n],(m)}\cap{\boldsymbol{\mathfrak{K}}}^{[n]}.$
 Observe that ${\boldsymbol{\mathfrak{B}}}^{[n],(m)}_{\operatorname{per}}$ is the union of periodic little bushes of height $m$.
@@
 Every periodic or preperiodic little bush ${\mathfrak{B}}_{a}^{[n],(m)}$ is within a unique little periodic or preperiodic filled Julia set ${\mathfrak{K}}^{[n]}_{a}$ associated with $f_{n}$.
-The ${\mathfrak{B}}_{t}^{[n],(m)}$ exhaust ${\mathfrak{K}}^{[n]}_{a}$:
-${\mathfrak{K}}^{[n]}_{a}=\overline{\bigcup_{m}{\mathfrak{B}}^{[n],(m)}_{a}}\ .$
+The ${\mathfrak{B}}_{a}^{[n],(m)}$ exhaust ${\mathfrak{K}}^{[n]}_{a}$:
+${\mathfrak{K}}^{[n]}_{a}=\overline{\bigcup_{m}{\mathfrak{B}}^{[n],(m)}_{a}}\$.
 By construction, the ${\mathfrak{B}}^{[n],(m)}_{a}$ are pairwise disjoint but the ${\mathfrak{K}}^{[n]}_{a}$ may touch each other in the satellite case.
@@
 It is well-known (follows, for instance, from the lamination theory) that the Hubbard continua ${\mathfrak{T}}_{f}$ and ${\mathfrak{T}}_{\circ}\equiv{\mathfrak{T}}_{f_{\circ}}$ are combinatorially equivalent up to ${\boldsymbol{\mathfrak{B}}}^{[n]}$: there is a bijection between
-  • components of ${\mathfrak{T}}_{f}\setminus f^{-m}\left({\boldsymbol{\mathfrak{B}}}^{[n]}_{f}\right)$ and components of ${\mathfrak{T}}_{\circ}\setminus f^{-m}_{\circ}\left({\boldsymbol{\mathfrak{B}}}^{[n]}_{\circ}\right)$ for all $m\geq 0$; and
-  • components of $f^{-m}\left({\boldsymbol{\mathfrak{B}}}^{[n]}_{f}\right)$ and components of $f^{-m}_{\circ}\left({\boldsymbol{\mathfrak{B}}}^{[n]}_{\circ}\right)$ for all $m\geq 0$
+[…]
+[…]
 that respects the adjacency, natural embedding, and dynamical relations between respective components. In other words, the above components define equivalent Markov partitions for $f,f_{\circ}$.
-${\mathfrak{B}}^{[1]}_{a}$${\mathfrak{B}}^{[1],(6p_{2})}_{a}$${\mathfrak{B}}^{\prime\prime}_{1}$$\ell_{1}$${\mathfrak{B}}^{\prime}_{1}$${\mathfrak{B}}_{i}^{[2]}$${\mathfrak{B}}_{i+1}^{[2]}$${\mathfrak{B}}^{\prime\prime}_{2}$${\mathfrak{B}}^{\prime}_{2}$$\ell_{2}$${\mathfrak{B}}^{[1]}_{b}$$\alpha$
+${\mathfrak{B}}^{[1]}_{a}$${\mathfrak{B}}^{[1],(6p_{2})}_{a}$${\mathfrak{B}}^{\prime\prime}$$\ell_{1}$${\mathfrak{B}}^{\prime}$${\mathfrak{B}}_{i_{1}}^{[2]}$${\mathfrak{B}}_{i_{2}}^{[2]}$${\mathfrak{B}}^{\prime\prime\prime}$${\mathfrak{B}}^{\text{iv}}$$\ell_{2}$${\mathfrak{B}}^{[1]}_{b}$$\alpha$
 Figure 6. Illustration to Lemma 2.1.
 2.3.3. Satellite combinatorics
-Suppose that $f$ is at least 33 33 times renormalizable and the first renormalization of $f$ is satellite.
+Suppose that $f$ is at least $3$ times renormalizable and the first renormalization of $f$ is satellite.
 Then all ${\mathfrak{K}}_{i}^{[1]}$ are organized in the satellite flower around the $\alpha$ fixed point of $f$:
-[…]
+${\mathfrak{B}}_{f}={\boldsymbol{\mathfrak{K}}}^{[1]}_{f}\hskip 28.45274pt\text{ because }\hskip 28.45274pt{\boldsymbol{\mathfrak{K}}}^{[1]}_{f}\supset{\mathfrak{T}}_{f},$
 where ${\mathfrak{T}}_{f}$ is the Hubbard continuum of $f$.
@@
 Lemma 2.1 (Satellite flower).
-For a satellite $f$ as above, let ${\mathfrak{B}}^{[1]}_{a},{\mathfrak{B}}^{[1]}_{b}$ be its two level 11 11 bushes.
-Then the continuum ${\mathfrak{T}}^{(6p_{2})}_{f}$ contains geodesic continua $\ell_{1}$ and $\ell_{2}$ (as in §2.2.2) connecting preperiodic lifts ${\mathfrak{B}}^{\prime}_{1},{\mathfrak{B}}^{\prime\prime}_{1}$ and ${\mathfrak{B}}^{\prime}_{2},{\mathfrak{B}}^{\prime\prime}_{2}$ of ${\mathfrak{B}}^{[1]}_{b}$ of preperiods $\leq 6p_{2}$ such that, see Figure 6:
+For a satellite $f$ as above, let ${\mathfrak{B}}^{[1]}_{a},{\mathfrak{B}}^{[1]}_{b}$ be two of its level $1$ periodic bushes.
+Then the continuum ${\mathfrak{T}}^{(6p_{2})}_{f}$ contains geodesic continua $\ell_{1}$ and $\ell_{2}$ (as in §2.2.2) connecting preperiodic lifts ${\mathfrak{B}}^{\prime},{\mathfrak{B}}^{\prime\prime}$ and ${\mathfrak{B}}^{\prime\prime\prime},{\mathfrak{B}}^{\text{iv}}$ of ${\mathfrak{B}}^{[1]}_{b}$ of preperiods $\leq 6p_{2}$ such that, see Figure 6:
   • $\ell_{1}$ and $\ell_{2}$ separate ${\mathfrak{B}}^{[1]}_{a}$ from $\alpha$ (within ${\mathfrak{K}}_{f}$),
-  • the geodesic continuum $\ell\subset{\mathfrak{T}}_{f}$ connecting $\ell_{1},\ell_{2}$ intersects little preperidic bushes ${\mathfrak{B}}_{i}^{[2]},{\mathfrak{B}}^{[2]}_{i+1}$of level 22 22 with preperiods less than $6p_{1}$;
-  • ${\mathfrak{B}}_{i}^{[2]},{\mathfrak{B}}^{[2]}_{i+1}$ are disjoint from $\ell_{1}\cup\ell_{2}$.
-We remark that $\ell,{\mathfrak{B}}_{i}^{[2]},{\mathfrak{B}}^{[2]}_{i+1}$ are within ${\mathfrak{B}}^{[1],(6p_{2})}_{a}$.
+  • the geodesic continuum $\ell\subset{\mathfrak{T}}_{f}$ connecting $\ell_{1},\ell_{2}$ intersects little preperiodic bushes ${\mathfrak{B}}_{i_{1}}^{[2]},{\mathfrak{B}}^{[2]}_{i_{2}}$ of level $2$ with preperiods less than $6p_{1}$;
+  • ${\mathfrak{B}}_{i_{1}}^{[2]},{\mathfrak{B}}^{[2]}_{i_{2}}$ are disjoint from $\ell_{1}\cup\ell_{2}$.
+We remark that $\ell,{\mathfrak{B}}_{i_{1}}^{[2]},{\mathfrak{B}}^{[2]}_{i_{2}}$ are within ${\mathfrak{B}}^{[1],(6p_{2})}_{a}$.
 Proof.
-Let ${\mathfrak{B}}_{\gamma}^{[2]}\subset{\mathfrak{B}}^{[1]}_{a}$ be the level 22 22 periodic bush closest to $\alpha$; i.e.
+Let ${\mathfrak{B}}_{\gamma}^{[2]}\subset{\mathfrak{B}}^{[1]}_{a}$ be the level $2$ periodic bush closest to $\alpha$; i.e.
 $({\mathfrak{T}}_{f}\cap{\mathfrak{B}}^{[1]}_{a})\setminus{\mathfrak{B}}^{[2]}_{\gamma}$ and $\alpha$ are in different components of ${\mathfrak{T}}_{f}\setminus{\mathfrak{B}}_{\gamma}^{[2]}$.
-Let the ${\mathfrak{B}}_{t}^{[2]}\subset{\mathfrak{B}}^{[1],(tp_{1})}_{a}$ be the lifts of ${\mathfrak{B}}_{\gamma}^{[2]}$ under $f^{tp_{1}}$ towards $\alpha$; i.e.
-each ${\mathfrak{B}}_{t}^{[2]}$ separate $\alpha$ from ${\mathfrak{B}}^{[1],(tp_{1}-p_{1})}_{a}$.
-Since $f^{6p_{2}}$ has critical points in ${\mathfrak{B}}^{[2]}_{1}$ and in ${\mathfrak{B}}_{4}^{[2]}$, we can select $\ell_{1}$ and $\ell_{2}$ passing through these critical points such that $\ell_{1},\ell_{2}$ connect preperiodic lifts ${\mathfrak{B}}^{\prime}_{1},{\mathfrak{B}}^{\prime\prime}_{1}$ and ${\mathfrak{B}}^{\prime}_{2},{\mathfrak{B}}^{\prime\prime}_{2}$ of ${\mathfrak{B}}^{[1]}_{b}$.
-Then $\ell_{1},\ell_{2}$ separate ${\mathfrak{B}}^{[1]}_{a}$ from $\alpha$, and $\ell_{1}\cup\ell_{2}$ separate ${\mathfrak{B}}^{[2]}_{2},{\mathfrak{B}}^{[2]}_{3}$ from $\alpha$ and ${\mathfrak{B}}^{[1]}_{a}$; i.e.
-we can take ${\mathfrak{B}}^{[2]}_{i},{\mathfrak{B}}^{[2]}_{i+1}$ to be ${\mathfrak{B}}^{[2]}_{2},{\mathfrak{B}}^{[2]}_{3}$.
+Let the ${\mathfrak{B}}_{i(t)}^{[2]}\subset{\mathfrak{B}}^{[1],(tp_{1})}_{a}$ be the lifts of ${\mathfrak{B}}_{\gamma}^{[2]}$ under $f^{tp_{1}}$ towards $\alpha$; i.e.
+each ${\mathfrak{B}}_{i(t)}^{[2]}$ separates $\alpha$ from ${\mathfrak{B}}^{[1],(tp_{1}-p_{1})}_{a}$.
+Since $f^{6p_{2}}$ has critical points in ${\mathfrak{B}}^{[2]}_{i(1)}$ and in ${\mathfrak{B}}_{i(4)}^{[2]}$, we can select $\ell_{1}$ and $\ell_{2}$ passing through these critical points such that $\ell_{1},\ell_{2}$ connect preperiodic lifts ${\mathfrak{B}}^{\prime},{\mathfrak{B}}^{\prime\prime}$ and ${\mathfrak{B}}^{\prime\prime\prime},{\mathfrak{B}}^{\text{iv}}$ of ${\mathfrak{B}}^{[1]}_{b}$.
+Then $\ell_{1},\ell_{2}$ separate ${\mathfrak{B}}^{[1]}_{a}$ from $\alpha$, and $\ell_{1}\cup\ell_{2}$ separate ${\mathfrak{B}}^{[2]}_{i(2)},{\mathfrak{B}}^{[2]}_{i(3)}$ from $\alpha$ and ${\mathfrak{B}}^{[1]}_{a}$; i.e.
+we can take ${\mathfrak{B}}^{[2]}_{i_{1}},{\mathfrak{B}}^{[2]}_{i_{2}}$ to be ${\mathfrak{B}}^{[2]}_{i(2)},{\mathfrak{B}}^{[2]}_{i(3)}$.
 ∎
@@
   • every component of $\ell_{i}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}$ is trivial in $X\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}$ (with respect to a proper homotopy, see §2.4.1).
-(In other words, the $\ell_{i}$ are contractible into ${\boldsymbol{\mathfrak{B}}}^{[n],(1)}$.) Figure 7 gives an example of an invariant arc diagram.
+(In other words, the $\ell_{i}$ are contractible into ${\boldsymbol{\mathfrak{B}}}^{[n],(1)}$.) Here $f^{*}(A)$ is the pullback of $A$. Figure 7 gives an example of an invariant arc diagram.
 $f$${\mathfrak{B}}^{[1]}_{0}$$\alpha$$\beta$${\mathfrak{B}}^{[1],(1)}_{0}$$\tilde{\alpha}$$\tilde{\beta}$
@@
 Then one of the little bushes ${\mathfrak{B}}^{[n]}_{i}$ is accessible from the outermost component of $\partial Y^{\prime}$.
-Therefore, we can select a proper path $\ell\subset Y^{\prime}$ connecting $\partial Y$ and ${\boldsymbol{\mathfrak{B}}}^{n}$.
-For every such a path $\ell$, its legal pullback $\widetilde{\ell}$ is any of its lift connecting ${\boldsymbol{\mathfrak{B}}}^{[n]}$ and $\partial X$ and concatenated by a path in $Y\setminus X$ so that ~$\widetilde{\ell}\subset Y\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ is a proper path connecting $\partial Y$ and ${\boldsymbol{\mathfrak{B}}}^{n}$.
+Therefore, we can select a proper path $\ell\subset Y^{\prime}$ connecting $\partial Y$ and ${\boldsymbol{\mathfrak{B}}}^{[n]}$.
+For every such a path $\ell$, its legal pullback $\widetilde{\ell}$ is any of its lift connecting ${\boldsymbol{\mathfrak{B}}}^{[n]}$ and $\partial X$ and concatenated by a path in $Y\setminus X$ so that $\widetilde{\ell}\subset Y\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ is a proper path connecting $\partial Y$ and ${\boldsymbol{\mathfrak{B}}}^{n}$.
 Since $A$ is invariant, $\widetilde{\ell}$ can be again represented in $Y^{\prime}$ up to a proper homotopy in $Y\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$.
@@
   • $[\gamma_{j}]\in\big(f^{m}\big)^{*}(A)$ are proper arcs in $f^{-m}(Y)\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}=f^{-m}\left(Y\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}\right)$; and
-  • every component of $\ell_{i}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ is trivial in $X\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ (with respect to a proper homotopy, see §2.4.1) .
+  • every component of $\ell_{i}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ is trivial in $X^{m}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ (with respect to a proper homotopy, see §2.4.1) .
 We call (2.11) the decomposition of $\gamma$ rel ${\boldsymbol{\mathfrak{B}}}^{[n],(m)}$. We define the expansivity number
@@
 $\alpha_{1},\alpha_{2}\in(f^{2p_{n}})^{*}(A)\hskip 17.07182pt\text{ with }{\ \ }f^{2p_{n}}(\alpha_{1})=f^{2p_{n}}(\alpha_{2})=\alpha_{\operatorname{new}}\in A$
-such that $\alpha$ overflows $\alpha_{1},\alpha_{2}$ in the following sense: any decomposition 2.11 of $\alpha=[\gamma]$ rel ${\boldsymbol{\mathfrak{B}}}^{[n],(2p_{n})}$ contains two paths $\gamma_{a},\gamma_{b}$ representing $\alpha_{1}=[\gamma_{a}]$ and $\alpha_{2}=[\gamma_{b}]$.
+such that $\alpha$ overflows $\alpha_{1},\alpha_{2}$ in the following sense: any decomposition (2.11) of $\alpha=[\gamma]$ rel ${\boldsymbol{\mathfrak{B}}}^{[n],(2p_{n})}$ contains two paths $\gamma_{a},\gamma_{b}$ representing $\alpha_{1}=[\gamma_{a}]$ and $\alpha_{2}=[\gamma_{b}]$.
 Proof.
@@
 A horizontal AD $A=\{\alpha_{i}=[\gamma_{i}]\}$ is $f^{p_{1}}$-invariant if every $\alpha_{i}$ can be represented up to a proper homotopy in $X\setminus{\boldsymbol{\Upsilon}}$ in an arbitrary small neighborhood of $f^{-p_{1}}\left({\boldsymbol{\Upsilon}}\cup\Gamma\right)$, where $\Gamma=\bigcup_{i}\gamma_{i}$.
-By replacing $f$ with its superattracting model (§2.3.2, compare also wit Lemma 2.2), we obtain:
+By replacing $f$ with its superattracting model (§2.3.2, compare also with Lemma 2.2), we obtain:
 Lemma 2.5 (Alignment with ${\mathfrak{T}}_{f}$).
@@
 For a primitive copy ${\mathcal{M}}_{i}$, set ${\mathcal{M}}_{i}^{\circ}\coloneqq{\mathcal{M}}_{i}$.
-By compactness, there is a $\Delta>1$ such that every $f_{c}$ with $c\in\bigcup_{i}{\mathcal{M}}_{i}^{\circ}$ has a ql restriction $g_{1}=f_{c}^{p_{1}}\colon X_{1,c}\to Y_{1,c}$ around its maximal little Julia set containing 00 0 such that $g_{1}$ satisfies (II) and such that $g_{1}$ is hybrid conjugate via $h_{c}\colon Y_{1,1}\to Y_{R(c)}$ with dilatation $\leq\Delta$ to a ql restriction $f_{R(c)}\colon X_{R(c)}\to Y_{R(c)}$.
+By compactness, there is a $\Delta>1$ such that every $f_{c}$ with $c\in\bigcup_{i}{\mathcal{M}}_{i}^{\circ}$ has a ql restriction $g_{1}=f_{c}^{p_{1}}\colon X_{1,c}\to Y_{1,c}$ around its maximal little Julia set containing $0$ such that $g_{1}$ satisfies (II) and such that $g_{1}$ is hybrid conjugate via $h_{c}\colon Y_{1,1}\to Y_{R(c)}$ with dilatation $\leq\Delta$ to a ql restriction $f_{R(c)}\colon X_{R(c)}\to Y_{R(c)}$.
 Moreover, we can select the $Y_{1,c}$ and $Y_{c}$ so that $Y_{c}\Supset Y_{1,c}$.
@@
 Moreover, by induction, $Y_{n}$ is disjoint from
-$\Upsilon\coloneqq\bigcup_{m\leq n}\left({\boldsymbol{\mathfrak{B}}}^{[m]}_{f}\setminus{\mathfrak{B}}^{[m]}_{f,0}\right).$
+[…]
 This implies that the quadratic-like prerenormalizations $g_{n}\colon X_{n}\to Y_{n}$ can be univalently lifted to the dynamical plane of any $\psi^{\bullet}$ renormalization
@@
 In this section, we will introduce $\psi^{\bullet}$-ql renormalization and discuss tools to detect its degeneration; see [A, K, KL1, L3, DL2] for details.
-Given a compact subset $K\Subset S$, we denote by ${\mathcal{F}}(S,K)$ the family of non-trivial proper curves in $S\setminus K$ emerging from $\partial S$. We write
+Given a compact subset $K\Subset S$, we denote by ${\mathcal{F}}(S,K)$ the family of non-trivial proper curves in $S\setminus K$ connecting $\partial K$ to $\partial S$. We write
 (3.1) ${\mathcal{W}}(S,K)={\mathcal{W}}\big({\mathcal{F}}(S,K)\big).$
@@
 The correspondence $F_{n,i}$ in (3.3) is called a $\psi^{\bullet}$-ql map.
-It consists of a degree 22 22 branched covering $f_{n,i}$ and an immersion $\iota_{n,i}$, see definitions in §3.4.2.
+It consists of a degree $2$ branched covering $f_{n,i}$ and an immersion $\iota_{n,i}$, see definitions in §3.4.2.
 We define the Julia set and bush of $F_{n,i}$ to be that of the ql map $f_{n,i}$ (3.2):
@@
 A $\psi^{\bullet}$-ql renormalization can be naturally defined for a $\psi^{\bullet}$-ql map; i.e., $\psi^{\bullet}$-ql renormalization can be iterated.
-Sup-Chain Rule for renormalization domains is stated in §3.4.4.
+The Sup-Chain Rule for renormalization domains is stated in §3.4.4.
 Just like a ql map $f\colon X\to Y$ can be restricted to $f\colon X^{k+1}\to X^{k}$, where $X^{k+1}=f^{-k}(X)$, a $\psi^{\bullet}$-ql map $F=(f,\iota)\colon U\rightrightarrows V$ can be restricted to
 $F=(f,\iota)\colon U^{k+1}\rightrightarrows U^{k}$
-by considering the fiber product §3.4.6. We have a natural $1\displaystyle 2^{k}:$ 1 branched covering $f^{k}\colon U^{k}\to V$ and an immersion $\iota^{k}\colon U^{k}\to V$.
+by considering the fiber product §3.4.6. We have a natural $2^{k}:1$ branched covering $f^{k}\colon U^{k}\to V$ and an immersion $\iota^{k}\colon U^{k}\to V$.
 The width of a $\psi^{\bullet}$-ql map $F\colon U\rightrightarrows V$ is
@@
 By monotonicity: ${\mathcal{W}}({\mathcal{R}}_{2})\leq{\mathcal{W}}({\mathcal{R}})$.
-Assume that ${\mathcal{W}}(E_{x})>2$. The left and right 11 11-buffers of $E_{x}$ are defined
+Assume that ${\mathcal{W}}(E_{x})>2$. The left and right $1$-buffers of $E_{x}$ are defined
 $B^{\ell}_{1}\coloneqq[0,1]\times[0,1]{\ \ }\text{ and }{\ \ }B^{\rho}_{1}\coloneqq[x-1,x]\times[0,1]$
@@
 $E^{\operatorname{new}}_{x}\coloneqq[1,x-1]\times[0,1]=E_{x}\setminus\big(B^{\ell}_{1}\cup B^{\rho}_{1}\big)$
-is obtained from $E_{x}$ by removing 11 11-buffers. If ${\mathcal{W}}(E_{x})\leq 2$, then we set $E_{x}^{\operatorname{new}}\coloneqq\emptyset$. Similarly, buffers of any width are defined.
+is obtained from $E_{x}$ by removing $1$-buffers. If ${\mathcal{W}}(E_{x})\leq 2$, then we set $E_{x}^{\operatorname{new}}\coloneqq\emptyset$. Similarly, buffers of any width are defined.
 3.2.1. Monotonicity and Grötzsch inequality
-We say a family of curves $SS\displaystyle\SS$ overflows a family ${\mathcal{G}}$ if every curve $SS\displaystyle\gamma\in\SS$ contains a subcurve $\gamma^{\prime}\in{\mathcal{G}}$. We also say that
+We say a family of curves $S\displaystyle\SS$ overflows a family ${\mathcal{G}}$ if every curve $S\displaystyle\gamma\in\SS$ contains a subcurve $\gamma^{\prime}\in{\mathcal{G}}$. We also say that
   • a family of curves ${\mathcal{F}}$ overflows a rectangle ${\mathcal{R}}$ if ${\mathcal{F}}$ overflows ${\mathcal{F}}^{\operatorname{full}}({\mathcal{R}})$;
@@
 We naturally view vertical curves of $R$ as proper (and open) paths in $S$.
-We say that $R$ connects boundary components $J_{0},J_{1}\subset\partial^{i}R$ if $\partial^{h,0}R\subset J_{0},\ \partial^{h,1}R\subset J_{1}$.
+We say that $R$ connects boundary components $J_{0},J_{1}\subset\partial^{i}S$ if $\partial^{h,0}R\subset J_{0},\ \partial^{h,1}R\subset J_{1}$.
 If $S=U\setminus K$, where $U$ is a topological disk and $K\Subset U$ is a compact subset with finitely many connected components, then $\partial^{i}S$ has one outer component $\partial^{i}U$; all remaining inner components are parameterized by the components of $K$.
@@
 3.3.2. Universal cover
-For compact Riemann surface $V$ as in §2.4.1x, write its non-empty boundary as $\partial V=\bigsqcup_{k}J_{k}$.
+For compact Riemann surface $V$ as in §2.4.1, write its non-empty boundary as $\partial V=\bigsqcup_{k}J_{k}$.
+We assume that $V$ has negative Euler characteristic $\chi(V)<0$.
 We set $\pi:\mathbb{D}\rightarrow\operatorname{int}V$ to be the universal covering, and we let ${\Lambda}\subset{\mathbb{T}}=\partial\mathbb{D}$ be the limit set for the group ${\Delta}$ of deck transformations of $\pi$.
@@
 Any arc $\alpha$ connects some components $J_{k}$ and $J_{i}$ of $\partial V$ and lifts to an arc $\alpha$ in ${\mathbb{D}}$ connecting some intervals ${\mathbf{J}}_{k}$ and ${\mathbf{J}}_{i}$.
-We let ${\mathcal{W}}({{\mbox{\boldmath$\displaystyle\alpha$}}})={\mathcal{W}}_{\overline{\mathbb{D}}}({\mathbf{J}}_{k},{\mathbf{J}}_{i})$ to be the width of all curves in $\mathbb{D}$ connecting ${\mathbf{J}}_{k}$ and ${\mathbf{J}}_{i}$.
-Then ${\mathcal{W}}({{\mbox{\boldmath$\displaystyle\alpha$}}})$ is also the width of the rectangle ${\boldsymbol{\Pi}}_{\mbox{\boldmath$\displaystyle\alpha$}}=\overline{\mathbb{D}}$ with horizontal sides ${\mathbf{J}}_{k}$ and ${\mathbf{J}}_{i}$.
+We let ${\mathcal{W}}({{\mbox{$\displaystyle\alpha$}}})={\mathcal{W}}_{\overline{\mathbb{D}}}({\mathbf{J}}_{k},{\mathbf{J}}_{i})$ to be the width of all curves in $\mathbb{D}$ connecting ${\mathbf{J}}_{k}$ and ${\mathbf{J}}_{i}$.
+Then ${\mathcal{W}}({{\mbox{$\displaystyle\alpha$}}})$ is also the width of the rectangle ${\boldsymbol{\Pi}}_{\mbox{$\displaystyle\alpha$}}=\overline{{\mathbb{D}}}$ with horizontal sides ${\mathbf{J}}_{k}$ and ${\mathbf{J}}_{i}$.
 It is independent of the lift used.
-If ${\mathcal{W}}({{\mbox{\boldmath$\displaystyle\alpha$}}})>2$, then by removing from ${\boldsymbol{\Pi}}_{\mbox{\boldmath$\displaystyle\alpha$}}$ the square buffers we obtain a rectangle ${\boldsymbol{\Pi}}^{{\operatorname{can}}}_{\mbox{\boldmath$\displaystyle\alpha$}}$.
+If ${\mathcal{W}}({{\mbox{$\displaystyle\alpha$}}})>2$, then by removing from ${\boldsymbol{\Pi}}_{\mbox{$\displaystyle\alpha$}}$ the square buffers we obtain a rectangle ${\boldsymbol{\Pi}}^{{\operatorname{can}}}_{\mbox{$\displaystyle\alpha$}}$.
 In this case we let
-${\mathcal{W}}^{\operatorname{can}}(\alpha)={\mathcal{W}}^{\operatorname{can}}({{\mbox{\boldmath$\displaystyle\alpha$}}})={\mathcal{W}}({\boldsymbol{\Pi}}^{\operatorname{can}}_{\mbox{\boldmath$\displaystyle\alpha$}})={\mathcal{W}}({{\mbox{\boldmath$\displaystyle\alpha$}}})-2.$
-By construction, the $\Delta$-orbit of ${\boldsymbol{\Pi}}^{\operatorname{can}}_{\mbox{\boldmath$\displaystyle\alpha$}}$ consists of pairwise disjoint rectangles.
-Therefore, ${\boldsymbol{\Pi}}^{\operatorname{can}}_{\mbox{\boldmath$\displaystyle\alpha$}}$ projects injectively onto a rectangle ${\boldsymbol{\Pi}}^{\operatorname{can}}_{\alpha}$ in $V$.
-Moreover, different ${\boldsymbol{\Pi}}^{\operatorname{can}}_{{\mbox{\boldmath$\displaystyle\alpha$}}_{i}}$ project to pairwise disjoint ${\boldsymbol{\Pi}}^{\operatorname{can}}_{\alpha_{i}}$.
-If ${\mathcal{W}}({{\mbox{\boldmath$\displaystyle\alpha$}}})\leq 2$ we let ${\mathcal{W}}^{\operatorname{can}}(\alpha)=0$ and ${\boldsymbol{\Pi}}^{\operatorname{can}}(\alpha)\coloneqq\emptyset$.
+${\mathcal{W}}^{\operatorname{can}}(\alpha)={\mathcal{W}}^{\operatorname{can}}({{\mbox{$\displaystyle\alpha$}}})={\mathcal{W}}({\boldsymbol{\Pi}}^{\operatorname{can}}_{\mbox{$\displaystyle\alpha$}})={\mathcal{W}}({{\mbox{$\displaystyle\alpha$}}})-2.$
+By construction, the $\Delta$-orbit of ${\boldsymbol{\Pi}}^{\operatorname{can}}_{\mbox{$\displaystyle\alpha$}}$ consists of pairwise disjoint rectangles.
+Therefore, ${\boldsymbol{\Pi}}^{\operatorname{can}}_{\mbox{$\displaystyle\alpha$}}$ projects injectively onto a rectangle ${\boldsymbol{\Pi}}^{\operatorname{can}}_{\alpha}$ in $V$.
+Moreover, different ${\boldsymbol{\Pi}}^{\operatorname{can}}_{{\mbox{$\displaystyle\alpha$}}_{i}}$ project to pairwise disjoint ${\boldsymbol{\Pi}}^{\operatorname{can}}_{\alpha_{i}}$.
+If ${\mathcal{W}}({{\mbox{$\displaystyle\alpha$}}})\leq 2$ we let ${\mathcal{W}}^{\operatorname{can}}(\alpha)=0$ and ${\boldsymbol{\Pi}}^{\operatorname{can}}(\alpha)\coloneqq\emptyset$.
 Arcs with positive weight form the canonical weighted arc diagram (WAD) on $V$:
@@
 $\sum_{J_{k}}{\mathcal{A}}_{J_{k}}=2{\mathcal{A}}_{V}.$
-See [L3, §7.6.3] for a reference of the following fact.
+See [L3, §7.6] for a reference of the following fact.
 Theorem 3.1 (Thin-Thick Decomposition).
@@
 3.3.8. WAD under covering
-Assume that $f\colon U\to V$ as a finite-degree covering between surfaces with boundaries. Then, see [K, Lemma 3.3]
+Assume that $f\colon U\to V$ as a finite-degree covering between surfaces with boundaries, where $\chi(V)<0$. Then, see [K, Lemma 3.3]
 (3.10) ${\mathcal{A}}_{U}=f^{*}{\mathcal{A}}_{V},{\ \ }\text{ where }{\ \ }f^{*}{\mathcal{A}}_{V}=\sum_{\alpha\in{\mathcal{A}}_{V}}W^{\operatorname{can}}(\alpha)f^{-1}(\alpha)$
@@
 3.4.2. Definitions
-A ${}^{\bullet}$-quadratic-like map (“$\psi^{\bullet}$-ql map”) is a pair of holomorphic maps
+A pseudo-quadratic-like map ($\psi^{\bullet}$-ql map”) is a pair of holomorphic maps
 (3.12) $F=(f,\iota)\colon{\ \ }(U,{\mathfrak{B}}^{\prime})\rightrightarrows(V,{\mathfrak{B}})$
@@
   1. (I)
-    $f\colon U\to V$ is a double branched covering (we usually normalize it so that its critical point is located at 00 0);
+    $f\colon U\to V$ is a double branched covering (we usually normalize it so that its critical point is located at $0$);
   2. (II)
@@
 is the covering annulus (3.6) of $V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ rel $\partial{\mathfrak{B}}^{[n]}_{i}$ glued with ${\mathfrak{B}}^{[n]}_{i}$ and
-$U_{n,i}\coloneqq{\mathbb{A}}\left(U\setminus f^{-1}({\boldsymbol{\mathfrak{B}}}^{[n]}),\ \partial{\mathfrak{B}}^{[n](1)}_{i}\right)\bigcup_{\partial{\mathfrak{B}}^{[n](1)}_{i}}{\mathfrak{B}}^{[n](1)}_{i}$
+$U_{n,i}\coloneqq{\mathbb{A}}\left(U\setminus f_{n,i}^{-1}({\boldsymbol{\mathfrak{B}}}^{[n]}),\ \partial{\mathfrak{B}}^{[n](1)}_{i}\right)\bigcup_{\partial{\mathfrak{B}}^{[n](1)}_{i}}{\mathfrak{B}}^{[n](1)}_{i}$
 is the covering annulus of $U\setminus f^{-1}({\boldsymbol{\mathfrak{B}}}^{[n]})$ rel $\partial{\mathfrak{B}}^{[n](1)}_{i}$ glued with ${\mathfrak{K}}^{[n]}_{i}$.
-We will write suppress index “00 0” for $F_{n,0}$:
+We will write suppress index $0$” for $F_{n,0}$:
 $F_{n,0}\equiv F_{n}=(f_{n},\iota_{n})\colon U_{n}\rightrightarrows V_{n},\hskip 11.38092pt\text{ where }{\ \ }U_{n}\equiv U_{n,0},\ V_{n}\equiv V_{n,0}.$
@@
 is $\psi^{\Upsilon}$-ql if it satisfies (I)(IV) from §3.4.2 so that ${\mathfrak{B}}^{\prime}\simeq{\mathfrak{B}}={\mathfrak{B}}(f_{X})$ are replaced with
-[…]
+$\Upsilon^{\prime}=\iota^{-1}(\Upsilon)\simeq\Upsilon=\Upsilon(f_{X}),\hskip 42.67912pt\text{ i.e., }\Upsilon\text{ is $\displaystyle\iota$-proper}.$
 For example, one can take $\Upsilon={\mathfrak{T}}$ or $\Upsilon={\mathfrak{T}}\cup{\boldsymbol{\mathfrak{B}}}^{[m]}$.
-To define “$\psi^{\Upsilon}$-renormalization”, one should extend a ql prerenormalization $f_{n,i}\colon{X_{n,i}\to Y_{n,i}}$ along all curves in $V\setminus\bigcup_{j}\Upsilon_{f_{n,j}}$.
+To define $\psi^{\Upsilon}$-renormalization”, one should extend a ql prerenormalization $f_{n,i}\colon{X_{n,i}\to Y_{n,i}}$ along all curves in $V\setminus\bigcup_{j}\Upsilon_{f_{n,j}}$.
 Moreover, $\Upsilon_{f_{n,j}}$ can often be enlarged, see §3.4.5.
@@
     ${\mathcal{R}}^{n_{1}\bullet}\circ{\mathcal{R}}^{n_{2}\bullet}(F)$, see (3.15), is the extension of $f_{n_{1}+n_{2},0}$ along curves in $V\setminus{\boldsymbol{\Upsilon}}$ (subject to certain restrictions illustrated on Figure 10), where
-    ${\boldsymbol{\Upsilon}}\coloneqq{\boldsymbol{\mathfrak{B}}}^{[n_{1}+n_{2}]}\cup\left({\boldsymbol{\mathfrak{B}}}^{[n_{2}]}\setminus{\mathfrak{B}}^{[n_{2}]}_{0}\right),$
+[…]
   2. (2)
@@
 (3.18) $F_{{\boldsymbol{\Upsilon}},1}=(f_{1,0},\iota_{1,0})\colon U_{{\boldsymbol{\Upsilon}},1}\to V_{{\boldsymbol{\Upsilon}},1}.$
-We recall from §2.4.5 that the set ${\boldsymbol{\Upsilon}}$ is inhomogeneous: $f_{X}$ does not permute components of ${\boldsymbol{\Upsilon}}$
+We recall from §2.4.5 that the set ${\boldsymbol{\Upsilon}}$ is inhomogeneous: $f_{X}$ does not permute components of ${\boldsymbol{\Upsilon}}$ .
 3.4.6. Restrictions
@@
 $\rho\colon M\setminus{\mathfrak{K}}_{G}\to B\setminus{\boldsymbol{\Upsilon}}$
-so that $\rho$ induces an isomorphism $\rho\colon\partial{\mathfrak{K}}_{G}\overset{\simeq}{\longrightarrow}\partial{\mathfrak{K}}^{[n]}_{i}.$ The Sup-Chain Rule §3.4.4 implies by induction that $F\subset G$.
+so that $\rho$ induces an isomorphism $\rho\colon\partial{\mathfrak{K}}_{G}\overset{\simeq}{\longrightarrow}\partial{\mathfrak{K}}^{[n]}_{i}$.
+The Sup-Chain Rule §3.4.4 implies by induction that $F\subset G$.
 Therefore, we have the induced partial covering:
-[…]
-so that $\rho^{\prime}\colon\partial{\mathfrak{K}}_{F}\overset{\simeq}{\longrightarrow}\partial{\mathfrak{K}}^{[n]}_{i}.$ Let $\gamma$ be the core hyperbolic geodesic of the annulus $V\setminus{\mathfrak{K}}_{F}$.
+(3.20) $\rho^{\prime}\colon V\setminus{\mathfrak{K}}_{F}\hookrightarrow M\setminus{\mathfrak{K}}_{G}\overset{\rho}{\longrightarrow}B\setminus{\boldsymbol{\Upsilon}}$
+so that $\rho^{\prime}\colon\partial{\mathfrak{K}}_{F}\overset{\simeq}{\longrightarrow}\partial{\mathfrak{K}}^{[n]}_{i}$.
+Let $\gamma$ be the core hyperbolic geodesic of the annulus $V\setminus{\mathfrak{K}}_{F}$.
 Then $\gamma_{F}$ can be described as the set points $x$ such that the Brownian motion starting at $x$ has equal probabilities of hitting each of the boundary components of the annulus $V\setminus{\mathfrak{K}}_{F}$.
@@
 3.5.2. WADs of $F$
-Assume that ${\boldsymbol{\mathfrak{B}}}^{[n]}$ is well-defined, i.e., $F$ is $n+1$ renormalizable. Let us consider the canonical WAD of $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],m}$:
-${\mathcal{A}}^{(m),k}\equiv{\mathcal{A}}^{[n],(m),k}\coloneqq{\operatorname{WAD}}\left(U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],m}\right).$
+Assume that ${\boldsymbol{\mathfrak{B}}}^{[n]}$ is well-defined, i.e., $F$ is $n+1$ renormalizable.
+Let us consider the canonical WAD of $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$, $k\geq m$:
+${\mathcal{A}}^{(m),k}\equiv{\mathcal{A}}^{[n],(m),k}\coloneqq{\operatorname{WAD}}\left(U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}\right).$
 We say (compare with §1.3.1) that an arc $\alpha\in{\mathcal{A}}^{[n],(m),k}$ is
@@
   • ${\mathcal{A}}^{(m),k}_{\mathrm{hor}}$ to be horizontal part of ${\mathcal{A}}^{(m),k}$: it consists of arcs connecting components of $\partial{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$;
-  • ${\mathcal{A}}_{i}^{(m),k}$ to be the local WAD for ${\mathfrak{B}}^{[n],(m),k}_{i}$, see §3.3.5: ${\mathcal{A}}_{i}^{k,(m)}$ consisting of arcs adjacent to ${\mathfrak{B}}^{[n],(m),k}_{i}$ such that the weights of self-arcs adjacent to ${\mathfrak{B}}^{[n],(m),k}_{i}$ are doubled;
+  • ${\mathcal{A}}_{i}^{(m),k}$ to be the local WAD for ${\mathfrak{B}}^{[n],(m),k}_{i}$, see §3.3.5: ${\mathcal{A}}_{i}^{(m),k}$ consisting of arcs adjacent to ${\mathfrak{B}}^{[n],(m),k}_{i}$ such that the weights of self-arcs adjacent to ${\mathfrak{B}}^{[n],(m),k}_{i}$ are doubled;
   • ${\mathcal{A}}_{i,{\mathrm{ver}}}^{(m),k},{\mathcal{A}}_{i,{\mathrm{hor}}}^{(m),k}$ to be the local parts of ${\mathcal{A}}^{(m),k}_{\mathrm{ver}},\ {\mathcal{A}}^{(m),k}_{\mathrm{hor}}$.
@@
 Lemma 3.10.
-Assume that in the dynamical plane of $F={\mathcal{R}}^{n_{0}\bullet}\colon U\rightrightarrows V$, there is a non-trivial horizontal lamination of curves ${\mathcal{L}}\subset V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ emerging from ${\mathfrak{B}}^{[n]}_{0}$ and landing at ${\boldsymbol{\mathfrak{B}}}^{[n]}$.
+Assume that in the dynamical plane of $F={\mathcal{R}}^{n_{0}\bullet}G\colon U\rightrightarrows V$, there is a non-trivial horizontal lamination of curves ${\mathcal{L}}\subset V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ emerging from ${\mathfrak{B}}^{[n]}_{0}$ and landing at ${\boldsymbol{\mathfrak{B}}}^{[n]}$.
 Then
@@
 As with $\psi$-ql renormalization (compare with [McM1, Theorem 9.3]), we have
-(3.24) $\frac{1}{2}{\mathcal{W}}(F_{n,0})\leq{\mathcal{W}}(F_{n,i})\leq{\mathcal{W}}(F_{n,0})\hskip 19.91684pt\text{ for all }\ i.$
-Indeed, let $\gamma_{i}\subset V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ be the peripheral hyperbolic geodesic around ${\mathfrak{K}}^{[n]}_{i}$.
-Then the hyperoblic length of $|\gamma_{i}|_{V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}}$ is proportional to ${\mathcal{W}}(F_{n,i})$.
-Let $\gamma^{1}_{i}$ be the lift of $\gamma_{i}$ under $f\colon U^{(1)}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}\to V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ so that $\gamma^{m}_{i}\subset U^{(m)}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ is the peripheral hyperbolic geodesic around ${\mathfrak{B}}^{[n],(1)}_{i-1}$, where the subscript is $\operatorname{mod}p_{n}$.
+(3.24) $\frac{1}{2}{\mathcal{W}}_{\bullet}(F_{n,0})\leq{\mathcal{W}}_{\bullet}(F_{n,i})\leq{\mathcal{W}}_{\bullet}(F_{n,0})\hskip 19.91684pt\text{ for all }\ i.$
+Indeed, let $\gamma_{i}\subset V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ be the peripheral hyperbolic geodesic around ${\mathfrak{B}}^{[n]}_{i}$.
+Then the hyperoblic length of $|\gamma_{i}|_{V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}}$ is proportional to ${\mathcal{W}}_{\bullet}(F_{n,i})$.
+Let $\gamma^{1}_{i}$ be the lift of $\gamma_{i}$ under $f\colon U^{1}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}\to V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ so that $\gamma^{1}_{i}$ is the peripheral hyperbolic geodesic around ${\mathfrak{B}}^{[n],(1)}_{i-1}$, where the subscript is $\operatorname{mod}p_{n}$.
 Counting the degree of $f\colon\gamma^{1}_{i}\to\gamma_{i}$, we obtain
-  • $|\gamma^{1}_{i}|_{U^{(1)}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}}=|\gamma_{i}|_{V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}}$ if $i>1$; and
-  • $\frac{1}{2}|\gamma^{1}_{1}|_{U^{(1)}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}}=|\gamma_{1}|_{V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}}$.
-Finally, $|\gamma^{1}_{i}|_{U^{(1)}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}}\geq|\gamma_{i-1}|_{V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}}$; this implies (3.24).
+  • $|\gamma^{1}_{i}|_{U^{1}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}}=|\gamma_{i}|_{V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}}$ if $i\not=1$; and
+  • $\frac{1}{2}|\gamma^{1}_{1}|_{U^{1}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}}=|\gamma_{1}|_{V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}}$.
+Finally, $|\gamma^{1}_{i}|_{U^{1}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(1)}}\geq|\gamma_{i-1}|_{V\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}}$; this implies (3.24).
 3.5.4. Covering: from $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{(m)}$ to $U^{k+s}\setminus{\boldsymbol{\mathfrak{B}}}^{(m+s)}$
@@
 Assume that $m\leq k$.
-By Lemma 3.6, the set ${\boldsymbol{\mathfrak{B}}}^{[n],(m)}\subset{\mathfrak{B}}_{F}$ is $\iota$-proper.
+By Lemma 3.6, the set ${\boldsymbol{\mathfrak{B}}}^{[n],(m)}\subset U^{k}$ is $\iota$-proper.
 Let $\gamma$ be a proper curve in $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ emerging from ${\mathfrak{B}}^{[n],(m)}_{a}$.
@@
 Proof.
-The curve $\gamma\in{\mathcal{F}}^{\operatorname{full}}(R)$ becomes vertical under the restriction by $(\iota^{s})^{*}$ if and only if $\gamma$ hits $\iota^{s}\big{(}\partial U^{k+s}\big{)}$.
-This is impossible because $\gamma_{1},\gamma_{2}$ do not hit $\iota^{s}\big{(}\partial U^{k+s}\big{)}$.
+The curve $\gamma\in{\mathcal{F}}^{\operatorname{full}}(R)$ becomes vertical under the restriction by $(\iota^{s})^{*}$ if and only if $\gamma(t)$ hits $\iota^{s}\big(\partial U^{k+s}\big)$.
+This is impossible because $\gamma_{1}(t),\gamma_{2}(t)$ do not hit $\iota^{s}\big(\partial U^{k+s}\big)$.
 ∎
@@
 In particular, horizontal arcs $\alpha$ in $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ are naturally viewed as horizontal arcs in $U^{s}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ by realizing the $\alpha$ in a small neighborhood of ${\mathfrak{K}}_{F}$ where $\iota^{s-k}$ is an embedding.
-It follows essentially from Lemma 3.3 (by lifting $\iota$ to the universal coverings, see [K, §3.5] for reference) that
-(3.27) ${\mathcal{A}}^{[n],(m),k}_{\mathrm{hor}}\geq{\mathcal{A}}^{[n],(m),k+1}_{\mathrm{hor}}\hskip 14.22636pt\text{ and }\hskip 14.22636pt{\mathcal{A}}^{[n],(m),k}_{i,{\mathrm{hor}}}\geq{\mathcal{A}}_{i+1,{\mathrm{hor}}}^{[n],(m),k+1}.$
+It follows essentially from Lemma 3.3 (by lifting $\iota$ to the universal coverings, see [K, §3.5] for reference) that for $k\geq m$
+(3.27) ${\mathcal{A}}^{[n],(m),k}_{\mathrm{hor}}\geq{\mathcal{A}}^{[n],(m),k+1}_{\mathrm{hor}}\hskip 14.22636pt\text{ and }\hskip 14.22636pt{\mathcal{A}}^{[n],(m),k}_{i,{\mathrm{hor}}}\geq{\mathcal{A}}_{i,{\mathrm{hor}}}^{[n],(m),k+1}.$
 3.5.7. Domination: from $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{(m)}$ to $U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{(m+1)}$ (following [K, §3.6])
@@
 We call $s$ the expansivity number of $\gamma$ in rel ${\boldsymbol{\mathfrak{B}}}^{(m+1)}$.
-It follows from Decomposition (3.28) that there is a constant $C\equiv C_{p_{n},m}$ such that for all $M>0$
-[…]
-in the following sense, see §3.3.6 for notation. For every arc $\alpha$ in ${\mathcal{A}}^{[n],(m+1),k}_{\mathrm{hor}}-{\operatorname{max}}\{M,1\}C$, there is a vertical sublamination
-${\mathcal{L}}_{\alpha}\subset{\mathcal{F}}(e)\equiv{\boldsymbol{\Pi}}^{\operatorname{can}}(\alpha){\ \ }{\ \ }\text{ with }{\ \ }{\ \ }{\mathcal{W}}({\mathcal{L}}_{\alpha})\geq{\mathcal{W}}(\alpha)-{\operatorname{max}}\{M,1\}C$
+It follows from Decomposition (3.28) that there is a combinatorial constant $C\equiv C_{p_{n},m}$ (essentially, the maximal number of possible arcs rel ${\boldsymbol{\mathfrak{B}}}^{[n],(m+1)}$) such that for all $M>1$
+(3.29) ${\mathcal{A}}^{[n],(m),k}_{\mathrm{hor}}-MC{\ \ }{\ \ }\text{ is \emph{dominated} by }{\ \ }{\ \ }{\mathcal{A}}^{[n],(m+1),k}_{\mathrm{hor}}-M$
+in the following sense, see §3.3.6 for notation. For every arc $\alpha$ in ${\mathcal{A}}^{[n],(m),k}_{\mathrm{hor}}-MC$, there is a vertical sublamination
+${\mathcal{L}}_{\alpha}\subset{\mathcal{F}}(e)\equiv{\boldsymbol{\Pi}}^{\operatorname{can}}(\alpha){\ \ }{\ \ }\text{ with }{\ \ }{\ \ }{\mathcal{W}}({\mathcal{L}}_{\alpha})\geq{\mathcal{W}}(\alpha)-MC$
 such that Decomposition 3.28 of every $\gamma\in{\mathcal{L}}_{\alpha}$ has the additional property that
-(3.29) $\gamma_{i}\in{\mathcal{F}}^{\operatorname{full}}\left({\mathcal{A}}^{[n],(m+1),k}_{\mathrm{hor}}-C\right).$
-The lamination ${\mathcal{L}}_{a}$ is constructed by removing all $\gamma$ from ${\mathcal{F}}(\alpha)$ that do not satisfy (3.29) – the width of removed curves is bounded by ${\operatorname{max}}\{M,1\}C$.
+(3.30) $\gamma_{i}\in{\mathcal{F}}^{\operatorname{full}}\left({\mathcal{A}}^{[n],(m+1),k}_{\mathrm{hor}}-M\right).$
+The lamination ${\mathcal{L}}_{a}$ is constructed by removing all $\gamma$ from ${\mathcal{F}}(\alpha)$ that do not satisfy (3.30) – the width of removed curves is bounded by $MC$.
 3.5.8. Almost periodic rectangles
@@
 An almost periodic rectangle $R$ between little bushes ${\mathfrak{B}}^{[n]}_{a},{\mathfrak{B}}^{[n]}_{a+1}$ is defined in the same way as in the case $n=1$.
-Such a rectangle $R$ can be lifted to the dynamical plane of $F_{n-1,c}$ (via the covering map representing the $\psi^{\bullet}$-renormalization, see §3.4.3) and its lift will be an almost periodic rectangle between level 11 11 little bushes as in Definition 3.12 .
+Such a rectangle $R$ can be lifted to the dynamical plane of $F_{n-1,c}$ (via the covering map representing the $\psi^{\bullet}$-renormalization, see §3.4.3) and its lift will be an almost periodic rectangle between level $1$ little bushes as in Definition 3.12 .
 4. Pull-off for non-periodic rectangles
@@
 And since $f^{-1}(U^{k}\setminus{\boldsymbol{\mathfrak{B}}}^{[n]})\subset U^{k+1}\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$, the $A^{k}_{\mathrm{hor}}$ stabilize at an invariant AD; i.e., $A^{k}_{\mathrm{hor}}$ is aligned with the Hubbard tree of the superattracting model.
-Consider the local part ${\mathcal{A}}^{k}_{i}$ of ${\mathcal{A}}^{k}$ around ${\mathfrak{B}}^{n}_{i}$. Write $q\coloneqq(10\overline{p}+3)p_{n}$. If for all $i$ we have
+Consider the local part ${\mathcal{A}}^{k}_{i}$ of ${\mathcal{A}}^{k}$ around ${\mathfrak{B}}^{n}_{i}$.
+Write $q\coloneqq(10\overline{p}+3)p_{n}$.
+We remark that since $q$ is linear in $p_{n}$, the $q$th iterate has a bounded degree on small bushes and the below application of the Covering Lemma will lead to estimates independent of $n$.
+If for all $i$ we have
 (4.1) ${\mathcal{W}}({\mathcal{A}}^{q}_{i})-{\mathcal{W}}({\mathcal{A}}^{0}_{i})\leq\delta_{1}K,{\ \ }{\ \ }{\ \ }\delta_{1}\ll\delta,$
@@
 If (4.1) does not hold some $i$, then ${\mathcal{W}}({\mathcal{A}}^{q}_{i,{\mathrm{ver}}})\geq\delta_{1}K$ and applying the Covering Lemma we obtain
-${\mathcal{W}}\left(V\setminus{\mathfrak{K}}_{j}^{[n]}\right)\asymp_{\delta_{1},\overline{p}}\ K\hskip 14.22636pt\text{ for all }j.$
+${\mathcal{W}}\left(V\setminus{\mathfrak{B}}_{j}^{[n]}\right)\asymp_{\delta_{1},\overline{p}}\ K\hskip 14.22636pt\text{ for all }j.$
 Applying the Quasi-Additivity Law and using $p_{n}\gg_{\delta_{1},\overline{p}}1$, we obtain
-${\mathcal{W}}(F)\geq\sum_{j}{\mathcal{W}}\left(V\setminus{\mathfrak{K}}_{j}^{[n]}\right)\succeq_{\delta_{1},\overline{p}}\ p_{n}K\gg 2K,$
+${\mathcal{W}}(F)\succeq_{\delta_{1},\overline{p}}\ \sum_{j}{\mathcal{W}}\left(V\setminus{\mathfrak{B}}_{j}^{[n]}\right)\succeq_{\delta_{1},\overline{p}}\ p_{n}K\gg 2K,$
 see §4.2; this is Case (P).
@@
 Recall from §3.5.4 and §3.5.6 that we have
-(4.2) ${\mathcal{A}}^{(m+1),k+1}_{{\mathrm{hor}}}=f^{*}\left({\mathcal{A}}^{(m),k}_{{\mathrm{hor}}}\right){\ \ }{\ \ }\text{ and }{\ \ }{\ \ }{\mathcal{A}}^{(m),k+1}_{\mathrm{hor}}\geq{\mathcal{A}}^{(m),k}_{\mathrm{hor}}.$
-Let us choose a sufficiently big threshold $C\gg_{p_{n}}1$. By (4.2), the ADs forming ${\mathcal{A}}^{k}_{\mathrm{hor}}-C^{k}$ decrease, hence for a certain
+(4.2) ${\mathcal{A}}^{(m+1),k+1}_{{\mathrm{hor}}}=f^{*}\left({\mathcal{A}}^{(m),k}_{{\mathrm{hor}}}\right){\ \ }{\ \ }\text{ and }{\ \ }{\ \ }{\mathcal{A}}^{(m),k+1}_{\mathrm{hor}}\leq{\mathcal{A}}^{(m),k}_{\mathrm{hor}}.$
+Let us choose a sufficiently big threshold $C\gg_{p_{n}}1$ so that (3.29) is applicable for $m=0$. By (4.2), the ADs forming ${\mathcal{A}}^{k}_{\mathrm{hor}}-C^{k}$ decrease, hence for a certain
 (4.3) ${\mathbf{t}}\leq\ \big[\text{maximal number of horizontal arcs on $\displaystyle V\setminus{\boldsymbol{\mathfrak{B}}}_{n}$}\big]\ \leq 3p_{n}$
@@
 coincides with the AD of ${\mathcal{A}}^{{\mathbf{t}}+1}_{\mathrm{hor}}-C^{{\mathbf{t}}+1}$.
-Since the WAD ${\mathcal{A}}^{{\mathbf{t}}+1}_{\mathrm{hor}}-C^{{\mathbf{t}}+1}$ on $U^{{\mathbf{t}}+1}\setminus{\boldsymbol{\mathfrak{B}}}$ is dominated by the WAD $f^{*}\left({\mathcal{H}}\right)=f^{*}\left({\mathcal{H}}-{\mathcal{A}}^{\mathbf{t}}_{\mathrm{hor}}-C^{\mathbf{t}}\right)$ on $U^{{\mathbf{t}}+1}\setminus{\boldsymbol{\mathfrak{B}}}^{(1)}$, see §3.5.7, we obtain that $H$ is an invariant AD, and therefore it is aligned with the Hubbard continuum by Lemma 2.2.
+Since the WAD ${\mathcal{A}}^{{\mathbf{t}}+1}_{\mathrm{hor}}-C^{{\mathbf{t}}+1}$ on $U^{{\mathbf{t}}+1}\setminus{\boldsymbol{\mathfrak{B}}}$ is dominated by the WAD $f^{*}\left({\mathcal{H}}\right)=f^{*}\left({\mathcal{A}}^{\mathbf{t}}_{\mathrm{hor}}-C^{\mathbf{t}}\right)$ on $U^{{\mathbf{t}}+1}\setminus{\boldsymbol{\mathfrak{B}}}^{(1)}$, see (3.29) (where $M=C^{\mathbf{t}}$) in §3.5.7, we obtain that $H$ is an invariant AD, and therefore it is aligned with the Hubbard continuum by Lemma $2.$
 4.2. Case (P)
@@
 Write $q\coloneqq(10\overline{p}+3)p_{n}$ and fix a sufficiently small $\delta_{1}>0$. Assume that:
-[…]
+  1. (Ploc)
     there is a $\kappa$ such that
@@
 $\sum_{j=0}^{p_{n}}{\mathcal{W}}(V,{\mathfrak{B}}_{j})\succeq_{\delta_{1},\overline{p}}\ p_{n}K.$
-[…]
+Applying [KL1, Quasi Additivity Law with separation] together with (4.6) and using $p_{n}\gg_{\delta_{1},\overline{p}}\ 1$, we obtain
 ${\mathcal{W}}(V\setminus{\boldsymbol{\mathfrak{B}}})\succeq_{\delta_{1},\overline{p}}\ \sum_{j=0}^{p_{n}}{\mathcal{W}}(V,{\mathfrak{B}}_{j})\succeq_{\delta_{1},\overline{p}}\ p_{n}K\geq 2K.$
 We conclude that
-${\mathcal{W}}^{\bullet}(F)={\mathcal{W}}(V,\ {\mathfrak{B}}_{F})\geq{\mathcal{W}}(V,\ {\boldsymbol{\mathfrak{B}}})\geq 2K.$
+${\mathcal{W}}_{\bullet}(F)={\mathcal{W}}(V,\ {\mathfrak{B}}_{F})\geq{\mathcal{W}}(V,\ {\boldsymbol{\mathfrak{B}}})\geq 2K.$
+This establishes Case (P) from (Ploc).
 4.3. Case (S)
-[…]
+Let us now assume that (Ploc) does not hold: for all $i$ we have ${\mathcal{W}}({\mathcal{A}}_{i,{\mathrm{ver}}}^{q})\leq\delta_{1}K$.
+This implies that most paths in the canonical lamination ${\mathcal{F}}({\mathcal{A}}_{i})$ of ${\mathcal{A}}_{i}$ are horizontal and they restrict to horizontal paths under the immersions $\iota^{s}\colon(U^{s},{\boldsymbol{\mathfrak{B}}})\to(V,{\boldsymbol{\mathfrak{B}}})$ for $s\leq q$.
+We will use notations of §2.4.4.
 Let ${\mathcal{H}}$ be the WAD from (4.4) on $(U^{\mathbf{t}},{\boldsymbol{\mathfrak{B}}})$.
@@
 Since $H={\operatorname{AD}}({\mathcal{H}})$ is aligned with the Hubbard dendrite, $(\iota^{q^{\prime}})^{*}\left[{\mathcal{H}}_{c}-\delta_{1}K\right]$ is, after removing $O_{p_{n}}(1)$ curves, dominated by ${\mathcal{H}}^{(q^{\prime})}_{c}$, see §3.5.7.
-And since $(f^{q^{\prime}})_{*}\colon{\mathcal{H}}^{q^{\prime}}_{c}\to{\mathcal{H}}_{c}$ is a bijection, as most $\overline{p}\delta_{1}K+2C^{2}<\delta_{2}K$ curves in ${\mathcal{H}}_{c}$ can have expansivity number greater than 22 22, where $\delta_{1}\ll\delta_{2}\ll\delta$.
+And since $(f^{q^{\prime}})_{*}\colon{\mathcal{H}}^{q^{\prime}}_{c}\to{\mathcal{H}}_{c}$ is a bijection, as most $\overline{p}\delta_{1}K+2C^{2}<\delta_{2}K$ curves in ${\mathcal{H}}_{c}$ can have expansivity number greater than $2$, where $\delta_{1}\ll\delta_{2}\ll\delta$.
 This implies that most curves in ${\mathcal{F}}({\mathcal{H}}_{c})$ are within rectangles representing genuine periodic arcs §2.4.3 that connect neighboring level $n$ bushes of ${\mathfrak{B}}^{[n-1]}_{c}$, see Lemma 2.3.
@@
 5. Waves
-Given a compact connected filled set $X\subset{\mathbb{C}}$, we denote by $\partial^{c}X$ its Carathéodory boundary; i.e., the set of prime ends of ^$\widehat{\mathbb{C}}\setminus X$.
+Given a compact connected filled set $X\subset{\mathbb{C}}$, we denote by $\partial^{c}X$ its Carathéodory boundary; i.e., the set of prime ends of $\widehat{\mathbb{C}}\setminus X$.
 A Riemann map identifies $\partial^{c}\widetilde{X}$ with the unit circle ${\mathbb{T}}$.
@@
   • curves in ${\mathcal{S}}^{0}$ start and end at an interval $I^{0}\subset A_{1}$ that is a fundamental interval for the action $H\curvearrowright A_{1}$;
-  • curves in ${\mathcal{S}}^{t}$, $t>0$ start at an interval $I^{t}_{-}\subset A_{1}$ and end at an interval $I^{t}_{+}\subset A_{1}$ such that one of the intervals $I^{t}_{-},I^{t}_{+}$ is within a union of $1\displaystyle 2^{t}-$ 1 fundamental intervals for the action $H\curvearrowright A_{1}$ and the interval $J^{t}\subset A_{1},$ $J^{t}\cap J\not=\emptyset$ between $I^{t}_{-},I^{t}_{+}$ is the union of exactly $1\displaystyle 2^{t}-$ 1 fundamental intervals of the action $H\curvearrowright A_{1}$;
+  • curves in ${\mathcal{S}}^{t}$, $t>0$ start at an interval $I^{t}_{-}\subset A_{1}$ and end at an interval $I^{t}_{+}\subset A_{1}$ such that one of the intervals $I^{t}_{-},I^{t}_{+}$ is within a union of $2^{t}-1$ fundamental intervals for the action $H\curvearrowright A_{1}$ and the interval $J^{t}\subset A_{1},$ $J^{t}\cap J\not$$\emptyset$ between $I^{t}_{-},I^{t}_{+}$ is the union of exactly $2^{t}-1$ fundamental intervals of the action $H\curvearrowright A_{1}$;
   • $T=O(\ln m)$.
-Then ${\mathcal{W}}({\mathcal{S}}^{t})=O(1)$ for $t>0$ because ${\mathcal{S}}^{t}$ crosses its shift under $H$, see [DL2, Shift Argument §A.3].
+Then ${\mathcal{W}}({\mathcal{S}}^{t})=O(1)$ for $t>0$ because, after removing $1$ buffers, ${\mathcal{S}}^{t}$ crosses its shift under $H$, see [DL2, Shift Argument §A.3].
 Replacing ${\mathcal{S}}$ with ${\mathcal{S}}^{0}$, we obtain that curves in the ${\mathcal{S}}^{0}_{k}$ start and end at pairwise disjoint intervals.
-Removing an extra $0(1)$ outermost curves, we obtain that the new ${\mathcal{S}}^{\operatorname{new}}_{l}$ are pairwise disjoint.
+Removing an extra $O(1)$ outermost curves, we obtain that the new ${\mathcal{S}}^{\operatorname{new}}_{l}$ are pairwise disjoint.
 ∎
 Let us assume first that ${\mathfrak{K}}_{k}$ is periodic.
-Fix an iteration $f^{n}\mid T,\ n\leq M^{3}$ such that $f^{n}\mid{\mathfrak{K}}_{k}^{\iota}$ covers $\partial^{c}_{T}{\mathfrak{K}}_{k}$ at least twice.
+Since the number sides is $M$ and since the period of small Julia sets is bounded by $M$, we can fix an iteration $f^{n}\mid T$, where $n$ is bounded in terms of $M$ (for instance, $\ n\leq M^{3}$) such that $f^{n}\mid{\mathfrak{K}}_{k}^{\iota}$ covers $\partial^{c}_{T}{\mathfrak{K}}_{k}$ at least twice.
 Consider the associated iteration
@@
 By Lemma 5.2, the width of $\gamma$ with $\gamma_{a}=\gamma_{b}$ is $O_{M}(1)$.
+Write $K\coloneqq{\mathcal{W}}({\mathcal{S}})$.
 Let ${\mathcal{F}}^{\prime}$ be the family of all $\gamma$ in ${\mathcal{S}}$ with $\gamma_{a}\not=\gamma_{b}$.
@@
 Since ${\mathcal{S}}^{\prime}$ overflows ${\mathcal{S}}_{a}$ and then ${\mathcal{S}}_{b}$, either ${\mathcal{W}}({\mathcal{S}}_{a})$ or ${\mathcal{W}}({\mathcal{S}}_{b})$ is at least $4/3K-O_{M}(1)$.
-After removing $O_{M}(1)$ buffers and applying Lemma 5.2, the waves ${\mathcal{S}}_{a},{\mathcal{S}}_{b}$ map univalently by onto waves above ${\mathfrak{K}}_{k}^{\iota}$.
-We obtain a wave with width $\geq\frac{4}{3}K-O_{M}(1)$ and the whole argument can be repeated with a strictly bigger wave leading eventually to the contradiction.
+After removing $O_{M}(1)$ buffers and applying Lemma 5.2, the waves ${\mathcal{S}}_{a},{\mathcal{S}}_{b}$ map univalently by $f^{n}$ onto waves above ${\mathfrak{K}}_{k}^{\iota}$.
+We obtain a wave with width $\geq\frac{4}{3}K-O_{M}(1)$ and the whole argument can be repeated with a bigger wave leading eventually to the contradiction.
 Assume now that ${\mathfrak{K}}_{k}$ is strictly preperiodic.
@@
 First we will introduce the spiraling parameters rel $\Pi$ for a curve $\gamma\in R$ shortly summarized in the following remark:
-${\mathfrak{B}}^{[1]}_{a}$${\mathfrak{B}}^{[1],(q)}_{a}$${\mathfrak{B}}^{\prime\prime}_{1}$$\ell_{1}$${\mathfrak{B}}^{\prime}_{1}$${\mathfrak{B}}_{i}^{[2]}$${\mathfrak{B}}_{i+1}^{[2]}$${\mathfrak{B}}^{\prime\prime}_{2}$${\mathfrak{B}}^{\prime}_{2}$$\ell_{2}$${\mathfrak{B}}^{[1]}_{a+1}$${\mathfrak{B}}^{[1],(q)}_{a+1}$${\mathcal{G}}_{a}$$\alpha$$R^{(q)}$
+[…]
 Figure 11.
@@
 Therefore, the part ${\mathcal{G}}_{a}$ (violet) of $R$ between $\ell_{1}$ and $\ell_{2}$ is tremendously wide.
-The bushes ${\mathfrak{B}}^{\prime}_{1},{\mathfrak{B}}^{\prime\prime}_{1},{\mathfrak{B}}^{\prime}_{2},{\mathfrak{B}}^{\prime\prime}_{2}$ are lifts of ${\mathfrak{B}}^{[1]}_{a+1}$, see Figure 6.
+The bushes ${\mathfrak{B}}^{\prime},{\mathfrak{B}}^{\prime},{\mathfrak{B}}^{\prime\prime\prime},{\mathfrak{B}}^{\text{iv}}$ are lifts of ${\mathfrak{B}}^{[1]}_{a+1}$, see Figure 6.
 Remark 6.2.
-Recall the the pure Mapping Class Group $\operatorname{MCG}$ of a 33 33-punctured sphere $S^{2}\setminus\{a,b,\infty\}$ is trivial.
+Recall the the pure Mapping Class Group $\operatorname{MCG}$ of a $3$-punctured sphere $S^{2}\setminus\{a,b,\infty\}$ is trivial.
 By replacing $a,b$ with small open Jordan disks $D_{a},D_{b}$, we obtain the new group $\operatorname{MCG}(S^{2}\setminus(D_{a}\cup D_{b}\cup\{\infty\}))\simeq{\mathbb{Z}}^{2}$ consisting of Dehn-twists around $D_{a},D_{b}$, see [FM].
 This fact has the following implication.
-Let $O$ be a small disk-neighborhood of ${\mathfrak{B}}^{[1]}_{a}\cup\Pi\cup{\mathfrak{B}}^{[1]}_{a+1}$.
-Then $O^{\prime}\coloneqq O\setminus({\mathfrak{B}}^{[1]}_{a}\cup{\mathfrak{B}}^{[1]}_{a+1})$ is topologically a disk with two holes.
-Therefore, a path (simple curve) $\beta\subset O^{\prime}$ from ${\mathfrak{B}}^{[1]}_{a}$ to ${\mathfrak{B}}^{[1]}_{a+1}$ has a simple description: it first spirals $\tilde{t}_{a}\in{\mathbb{Z}}$ times around ${\mathfrak{B}}^{[1]}_{a}$ and then spirals $\tilde{t}_{a+1}\in{\mathbb{Z}}$ times around ${\mathfrak{B}}^{[1]}_{a+1}$, where $\Pi$ can be used as a reference “path” of zero spiralings.
-Below we will ignore the spiraling orientation and introduce the absolute quantity $t_{a}=|\tilde{t}_{a}|$.
+[…]
 Recall §3.3 that $\gamma$ lands at the ideal boundary of $V\setminus{\boldsymbol{\mathfrak{B}}}^{[1]}$.
@@
 Consider the components $\gamma_{0},\gamma_{1},\dots,\gamma_{f},\gamma_{f+1}$ of $\gamma_{\Pi}\setminus\Pi.$ Since $\gamma$ is properly homotopic into (a small neighborhood of) $\Pi$, there is a $t_{a}\leq f$ such that
-  • $\gamma_{t}\cup\Pi$ surrounds ${\mathfrak{B}}^{[1]}_{a}$ for $t\leq t_{a}$;
-  • $\gamma_{t}\cup\Pi$ does not surround ${\mathfrak{B}}^{[1]}_{a}$ for $t>t_{a}$.
+  • $\gamma_{t}\cup\Pi$ surrounds ${\mathfrak{B}}^{[1]}_{a}$ for $t\in\{1,2,\dots,t_{a}\}$;
+  • $\gamma_{t}\cup\Pi$ does not surround ${\mathfrak{B}}^{[1]}_{a}$ for $t\in\{t_{a}+1,\dots,f\}$.
 (It follows that $\gamma_{t}\cup\Pi$ surrounds ${\mathfrak{B}}^{[1]}_{a+1}$ for $t>t_{a}$, and only $\gamma_{t_{a}}\cup\Pi$ can surround both ${\mathfrak{B}}^{[1]}_{a}$ and ${\mathfrak{B}}^{[1]}_{a}$.)
@@
   1. (A)
-    The spiraling numbers of $\gamma_{1},\gamma_{2}\in R$ differ by at most 11 11.
+    The spiraling numbers of $\gamma_{1},\gamma_{2}\in R$ differ by at most $1$.
   2. (B)
-    If $\gamma_{s}\in R^{(s)}$ is the lift of $\gamma\in R$ into $R^{(s)}$, then $t_{a}(\gamma^{s})\leq\frac{1}{2^{s}}t_{a}(\gamma)$.
+    If $\gamma_{s}\in R^{(s)}$ is the lift of $\gamma\in R$ into $R^{(s)}$, then $t_{a}(\gamma_{s})\leq\frac{1}{2^{s}}t_{a}(\gamma)$.
 Indeed, (A) follows from the fact (the disjoint curves) $\gamma_{1},\gamma_{2}$ can be simultaneously put into the minimal position with $\Pi$.
@@
 There are two possibilities.
-If a substantial part of ${\mathcal{G}}_{a}$ travels through both ${\mathfrak{B}}^{[2]}_{i},{\mathfrak{B}}^{[2]}_{i+1}$, then this substantial part of ${\mathcal{G}}_{a}$ restricts to a wide lamination ${\mathcal{L}}$ between ${\mathfrak{B}}^{[2]}_{i},{\mathfrak{B}}^{[2]}_{i+1}$.
+If a substantial part of ${\mathcal{G}}_{a}$ travels through both ${\mathfrak{B}}^{[2]}_{i_{1}},{\mathfrak{B}}^{[2]}_{i_{2}}$, then this substantial part of ${\mathcal{G}}_{a}$ restricts to a wide lamination ${\mathcal{L}}$ between ${\mathfrak{B}}^{[2]}_{i_{1}},{\mathfrak{B}}^{[2]}_{i_{2}}$.
 Pushing ${\mathcal{L}}$ with respect to $f_{*}$ and $\iota^{*}$ and using Lemma 3.2 (the number of iterations and the degree is bounded in terms of $\overline{p}$), we obtain a family ${\mathcal{F}}$ of non-trivial proper curves in $V\setminus{\boldsymbol{\mathfrak{B}}}^{[2]}$ starting at ${\mathfrak{B}}^{[2]}_{0}$ such that ${\mathcal{W}}({\mathcal{F}})\succeq_{\overline{p}}C_{0,\delta}K$.
@@
 If a substantial part of ${\mathcal{F}}$ is horizontal, then we obtain the first estimate in (6.1); applying Lemma 3.10 to ${\mathcal{F}}$, we obtain the first estimate in (6.2).
-If a substantial part of ${\mathcal{G}}_{a}$ omits either ${\mathfrak{B}}^{[2]}_{i}$ or ${\mathfrak{B}}^{[2]}_{i+1}$, then we obtain a wide wave above one of the sides of ${\mathfrak{B}}^{[2]}_{i},{\mathfrak{B}}^{[2]}_{i+1}$; Wave Lemma 5.1 implies that ${\mathcal{W}}_{\bullet}(F)\geq C_{\delta}K$.
-[…]
+If a substantial part of ${\mathcal{G}}_{a}$ omits either ${\mathfrak{B}}^{[2]}_{i_{1}}$ or ${\mathfrak{B}}^{[2]}_{i_{2}}$, then we obtain a wide wave above one of the sides of ${\mathfrak{B}}^{[2]}_{i_{1}},{\mathfrak{B}}^{[2]}_{i_{2}}$; Wave Lemma 5.1 implies that ${\mathcal{W}}_{\bullet}(F)\geq C_{\delta}K$.
+${\mathfrak{B}}^{[1]}_{a}$${\mathfrak{B}}_{a}^{[1],(1)}$${\mathfrak{B}}^{[1]}_{a+1}$${\mathfrak{B}}_{a+1}^{[1],(1)}$${\mathcal{G}}_{a}$$R^{(1)}$$\alpha$
 Figure 12.
@@
 The argument of this section will be repeated in §7.0.1 with ${\mathfrak{B}}^{[1]}_{a+1}$ being replaced by a little Julia set ${\mathfrak{K}}^{[1]}_{0}$.
-Observe that no specific properties of ${\mathfrak{B}}^{[1]}_{a+1}$ has been used in this section (only that it is a forward invariant set that is disjoint from ${\mathfrak{B}}^{[1]}$).
+Observe that no specific properties of ${\mathfrak{B}}^{[1]}_{a+1}$ has been used in this section (only that it is a forward invariant set that is disjoint from ${\mathfrak{B}}^{[1]}_{a}$).
 7. Conclusions
@@
 If $F$ is an infinitely renormalizable $\psi^{\bullet}$-ql map of bounded type $\bar{p}$, then
-${\mathcal{W}}_{\bullet}\left[({\mathcal{R}}^{n\bullet})^{m}(F)\right]\leq K_{\overline{p}}{\ \ }{\ \ }{\ \ }\text{ for }{\ \ }m\ \gg_{{\mathcal{W}}(F)}\ 1,$
+${\mathcal{W}}_{\bullet}\left[({\mathcal{R}}^{n\bullet})^{m}(F)\right]\leq K_{\overline{p}}{\ \ }{\ \ }{\ \ }\text{ for }{\ \ }m\ \gg_{{\mathcal{W}}_{\bullet}(F)}\ 1,$
 ${\mathcal{R}}^{n\bullet}$ is a $\psi^{\bullet}$-ql renormalization (3.15).
@@
 Therefore, if there are no a priori beau $\psi^{\bullet}$-bounds, then we obtain
-${\mathcal{W}}_{\bullet}\left[({\mathcal{R}}^{n\bullet})^{m}(F)\right]\succeq_{F}C^{nm}_{\delta}K.$
+${\mathcal{W}}_{\bullet}\left[({\mathcal{R}}^{n\bullet})^{m}(F)\right]\succeq_{F}\ C^{n(m-1)}_{\delta}K.$
 This contradicts the Teichmüller contraction: by Proposition 2.7, see also Remark 2.8, the sequence $({\mathcal{R}}^{n\bullet})^{m}(F)$ restricts to ql maps $f_{nm}:X_{nm}\to Y_{nm}$ such that
@@
 Write $F\coloneqq f$, and consider the sequence $G_{m}\coloneqq({\mathcal{R}}^{n\bullet})^{m}(F)$ from Theorem 7.1. We have
-(7.1) ${\mathcal{W}}_{\bullet}(G_{m})\leq K_{\overline{p}}\hskip 28.45274pt\text{ for}m\gg_{{\mathcal{W}}(F)}1$
+(7.1) ${\mathcal{W}}_{\bullet}(G_{m})\leq K_{\overline{p}}\hskip 28.45274pt\text{ for }m\gg_{{\mathcal{W}}_{\bullet}(F)}1$
 for $K_{\overline{p}}$ from Theorem 7.1.
@@
 Assuming converse, the associated degeneration is arbitrary big:
-$K\coloneqq{\mathcal{W}}\left({\mathcal{G}}\right)\gg_{K_{\overline{p}}}1,\hskip 19.91684pt\text{ where }{\ \ }{\mathcal{G}}\coloneqq{\mathcal{F}}\left(V\setminus\bigcup_{i\not=0}{\mathfrak{B}}^{[1]}_{i},\ {\mathfrak{K}}^{[1]}_{0}\right).$
+[…]
 We will now argue that a substantial part of ${\mathcal{G}}$ travels through level two little bushes; this will lead to a contradiction to the $\psi^{\bullet}$-bounds of Theorem 7.2.